A Bibliography of Publications in Scientific Programming

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

02 June 2023 Version 1.15

Title word cross-reference

+ [BDV03]. 1 [TPKP13]. 2 [PPJ^{+}15]. 3 [JR10, MRLF12]. ^{3} [FP00a]. ^{H} [Che93a, Che93b]. ^{SM} [Ano07a]. _{2} [WCZ^{+}16]. d [TOF^{+}14]. K [RLC04]. N [HJ96, War14, WHG93]. O(N) [HJ96].

-body [HJ96, War14]. -D [TPKP13]. -dimensional [WHG93]. -means [RLC04]. -spline-based [TOF^{+}14].

1 [HR12a, KBRS95]. 1999 [Met99a, Met99b].

2 [GBJ94, HR12b]. 2.0 [HMSM08]. 20 [Met99a]. 200 [GBJ94, MJ95]. 2001 [Ger02]. 2003 [BCCP05, Mor15, RMX05].

2003/2008 [Mor15]. 2008 [Mor15]. 2012 [VR13]. 2048-processor [HMCH07]. 2HOT [War14]. 2nd [Nag05b].

3 [GBH14]. 3Es [SZC16].

4M [DTV00].

590 [Was95].

77 [OPE^{+}95, Zim07].

860 [CBCM93]. 8i [KGBB09].

90 [McC96]. 95 [RMX05, Sch03a]. 95/2003 [RMX05]. 9th [SO11].

ABC [JFPL16]. Abstract [Vol97, Rou08b]. abstractions [Hav00b]. Acceleration [Jes10, ZSS^{+}10]. Accelerator [TLC15, CIN^{+}96]. accelerators [KK11]. access [CGL08, GL04, GBH14, LOHA01, TKS02]. Accessing [TC96]. Accuracy [GMCT15, PDROFRM16, BBDN11]. Achieving [BAN02, HLPY16, KFFZ05]. Acquisition [ZGS16]. across [GR93]. active [RLC04]. Activity [ZG15]. Actor [CSM15, KPS16]. Actor-Fact [CSM15]. ad [BDM^{+}04]. Adaptation [SRO^{+}15, IRSD99]. Adaptive [PVL^{+}04, Shu94, BCC^{+}93, BAN02, CW93, EAS^{+}97, GRN99, OHS00, VCT05]. address [HUN08]. ADIFOR [BCC^{+}92]. Adjacency [ZSS^{+}10]. Adjacency-based [ZSS^{+}10]. adjoints [FC01, HUN08]. adoption [NDSG07]. Advanced [ATZ^{+}15]. advances [Spo12, HPS^{+}13]. Advantages [DF14]. agent [CJS^{+}02, CGK^{+}05, PVL^{+}04]. agent-based [CJS^{+}02]. aggregation [IMB^{+}13]. Agile [AKA16]. aided [IJL^{+}01]. AIPS [WHG93]. Alan [Ste97]. algebra [ACIK97, HKL^{+}12]. algebraic [DHH00, Hav00a]. Algorithm [ACLS16, CLTX16, FMA16, HJ15, JKR92, KS15, LWW16, MWQ16, SLBZ16, SZC16, TT16, WS15, WWLG16, ZYLZ15, ZHL16, KHSJ95, MJLM07, MSLS13, War14, ZSS^{+}10]. Algorithms [CS94, CBRRC^{+}16, CW93, DRL13, GIKP95, GA96, HJ96, KTP05, KK09, Lin04, LKDB10, NKV^{+}02, OHS00, RLC04, SZ09]. Alice [Nag04]. aligned [AGIS94]. All-Subgraph [SJ15]. All-to-all [MJ95]. Allocation [BS01, HLPY16, JFPL16, MPA16, Mor94, RMX05]. Amazon [JMR^{+}11, TFN11]. American [ZG15]. Amesos2 [BHRT12]. among [LZF16]. Analysis [CTCG15, HPZ16, HBCM94, KAS15, KPS16, KOM94, SC16, YG16, AE03, BCK07, BB08, BBF^{+}04, BCHL05, CO93, FO96, HMSM08, JMR^{+}11, LDV07, Lin04, PPS12a, PPS^{+}12b, PDGQ05, Rou08b, SCSJ09, SS00, SGM^{+}08, Sne95, TKS02, TF04, WGW08, GM14]. analysis-driven [Rou08b]. Analyzing [LSB12, ZCXQ16]. ANCS [HLPY16]. angular [KMB09]. Answers [SHM97]. Ant [FMA16]. Antispam [PDROFRM16]. AntStar [FMA16]. aperture [PTS^{+}13]. AppLeS [COB^{+}00]. Application [ATZ^{+}15, BCS01, KJG^{+}08, Mal01, PPS^{+}12b, YTD15, BBF^{+}04, BFGH14, BBK^{+}11, DFP96, HLM^{+}05, IJL^{+}01, KeKR^{+}11, Bor04]. Applications [Ang93, CHM^{+}97, LWYS16, RRM^{+}15, Sne95, BBBM12, BDV03, CHS^{+}99, CKS^{+}09, HZ99, HHML05, LTB02, LPHD04, Nor07, PVL^{+}04, RKR^{+}99, SJR14, SKU^{+}09, Slo12, SB01, SBJV11, WGW08]. applied [WGF93]. Applying [AlS16, PDROFRM16, SCB02]. Approach [AlS16, AKA16, CLA^{+}15, DDMS15, DEŁ01, FOT92, LGC^{+}16, LWYS16, SPPH15, SSN15, YZ16, ZWZ16, FP95, GG95, KJG^{+}08, KLS^{+}96, LSB12, LB02, PSU08]. approaches [OB96]. approximate [GG05]. Approximation [YTD15]. arbitrary [CGL08]. Architecting [IR02]. Architecture [AKA16, KGMN15, BHK^{+}13, PMTL12, RX04, Rou08b, WB95, KCO^{+}05]. Architectures [CCTY15, CKKK15, BFGH14, BBK^{+}11, HPD09, LKDB10, SH94, SJR14, WSP94]. area [JEM07]. arising [ARvW03]. Aristotle [FSV14]. arithmetic [Sch03a]. ARMS [CJS^{+}02]. Array [HBCM94, McC96, ESP^{+}12, Ott93, WHG93, BCS01]. arrays [FO96, RN07, TC96]. Art [Bry96, BDH^{+}10]. Artificial [Gil09]. ASKALON [Pro07]. Aspects [KAS15]. Aspen [SV15]. Assembly [JBLB15]. Assess [ÁRCPS15]. Assessing [LLW^{+}14]. Assessment [YG16, MCvM10]. Assigning [HY15]. Assigning-Weight [HY15]. Association [ZYLZ15]. Atmospheric [Ros00, SS00]. atomistic [ADLN08, NKV^{+}02]. atomistic-to-continuum [ADLN08]. attachment [MPP^{+}04]. Author [Ano97a, Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14]. auto [DF14, FML^{+}14]. auto-tuning [DF14, FML^{+}14]. Automated [SV15]. Automatic [FB99, HZ99, JJY^{+}03, SLD^{+}16, AGG^{+}97, BFGH14, IJL^{+}01, Kes96, LP99, OPE^{+}95]. Automatically [CBRRC^{+}16]. Automating [CQF05, PPS12a, PPS^{+}12b]. automation [HMSM08]. Autonomic [KeKR^{+}11]. Available [RRV09]. Aware [BHP^{+}03, CGC16, CLYL16, MJB15, OCC^{+}16, TT16, HJYC10, IMB^{+}13, Zim07]. Awareness [ZHL16]. AWS [JMR^{+}11]. Azure [HLM^{+}11].

B.E. [Ano08b, VvAC^{+}09]. Babb [Per08]. backprojection [PTS^{+}13]. backprojection-based [PTS^{+}13]. Backup [CTCG15]. Balance [CKKK15]. Balancing [CGC16, VMRM16, LTB02, PVL^{+}04]. banded [MSSG11]. bartering [Özt04]. Based [ATZ^{+}15, BD16, CLA^{+}15, CLTX16, CLYL16, DPMB16, HJ15, HY15, HHD^{+}16, JFPL16, KS15, KPS16, MFB^{+}15, MPA16, MM16, SPPH15, SSN15, SLBZ16, SZXL16, THS^{+}15, WS15, WB16, XTCC15, XML^{+}16, YZZ15, YZ16, YG16, ZCXQ16, ZWW16, ZWCY16, Abd02, ABB^{+}02, AMM05, BMY13, BB08, BGLR93, CJS^{+}02, CL04, GRN99, GGJ04, GG05, KK11, Lan01, OPP11, PTS^{+}13, PPS12a, Sch03b, TOF^{+}14, TKS02, Tým99, WOC99, ZSS^{+}10]. Basic [BBG^{+}93]. Bayesian [LB02]. BE [RRV09]. behavior [TKS02]. behaviour [GGJ04]. being [BKP99]. Belos [BHRT12]. Benchmark [Was95, DEvdV01, KGBB09, Mül03, PHH95, RST02]. benchmarking [RST02]. Benchmarks [PAR94, AE03, HB94, SBM^{+}10]. benefits [NJ08]. Best [Old00]. better [BPC^{+}09, GA96]. between [SH94]. Beyond [SF03a, Lin04]. biclusters [HB12]. Biermann [Ste97]. Big [CTC^{+}15]. Binary [YDC^{+}16, LSB12]. binding [VRW^{+}03]. Biobjective [CMTAC^{+}16]. Biodiesel [ZJZ^{+}16]. Biological [AHCZ12, ZRP08]. biomolecular [BPK00]. Birds [SCAP16]. BitTorrent [DRL13]. Black [YDC^{+}16]. BLAS [ARvW03]. Blitz [AJJF14]. Blogosphere [ZG15]. Bocca [ANEA08]. bodies [AGIS94]. body [HJ96, War14]. Book [Ano07a, Ano07b, Ano07c, Ano11b, Bai08, Bor04, Bro11a, Bro11b, Bro12, Bro09, Bus09, Fil14, GB10, Les10, Man08, Nag04, Nag05a, Nag05b, Nag09a, Nag09b, Nag11a, Nag11b, Nag12]. books [Met99a, Met99b]. Boosting [PDROFRM16]. Boot [HHD^{+}16]. Both [ZHL16]. breadth [BAP13]. breadth-first [BAP13]. bringing [Pla04]. Broadband [GGP09, IB09, KD09, SKU^{+}09, VGC09, WSZK09]. broadcasts [DRL13]. Broker [ACB^{+}02]. BSML [VRW^{+}03]. BSP [SHM97]. Building [VvAC^{+}09, RLL^{+}02]. Bulk [ZWCY16]. Burgers [MRLF12]. burst [BCHL05]. Bus [MM16]. butterfly [Sch03b].

C [Ang93, CO93, Che93a, Che93b, CP03, DNS97, DDMS15, LH93]. Cache [AlS16, NsP16, TLC15, Bai95, WC96]. caching [Tým99]. CAF [Mar05]. CaKernel [BBK^{+}11]. calculus [RMX05, Rou08b]. Call [Ano97b]. Camera [ACLS16]. can [Ang93, BFH96]. cancer [AMN^{+}12]. CanonicalProducer [KCO^{+}05]. capabilities [MRLF12, PPS12a, PPS^{+}12b]. Case [DS97, Hav00a, McC96, RRM^{+}15, AJJF14, GW11, GB94, GL04, MS00, Rou08b, ZS99]. Cassandra [CTCG15]. Causal [LDV07]. CCA [ANEA08]. Cell [SCAP16, Ano08b, GGP09, IB09, KD09, RRV09, SKU^{+}09, VvAC^{+}09, VGC09, WSZK09]. Cell/B.E. [VvAC^{+}09]. CellSs [BPC^{+}09]. Center [CTCG15, CGC16]. Centered [DDMS15].

Centers [JFPL16, LZF16, ZHL16]. Centric [TLC15]. CFD [Jes10, JR10]. Chain [DPMB16, ZJZ^{+}16]. chains [HS03]. challenges [NJ08]. Chance [SZC16]. Chance-Constrained [SZC16]. Characteristics [VMRM16]. Characterization [PCGLO14, MJLM07]. Characterizing [OdSSP13]. charge [CBCM93]. Cheaper [HUN08]. Checkpoint [CSC16]. checkpointing [IMB^{+}13]. Chemistry [KJG^{+}08, SE95]. chip [JK10]. CHIPPER [MM16]. Chips [TLC15]. choice [AJJF14, FB99]. Cholesky [CDO^{+}96]. circulation [Hem00, SS00]. Class [NCMF15, SO93, Ang93]. Class-specific [SO93, Ang93]. classes [Ott93, VHBR93, WHG93]. Classical [PDROFRM16]. Classification [CLYL16, KG08]. Classifier [MWQ16]. Classifiers [PDROFRM16]. Cloud [ALL^{+}16, AlS16, BR11, CTCG15, CTC^{+}15, CMTAC^{+}16, HPZ16, KPS16, OCC^{+}16, Wu16, XST^{+}16, ZHL16, JMR^{+}11, SBJV11, TSCT11]. Clouds [CLYL16, HLPY16, MFB^{+}15, ZWZ16, OPP11]. Cluster [ACCTA^{+}15, SHHI01, SCD^{+}15, HMCH07, JEM07, GWZ16]. Cluster-enabled [SHHI01]. clustering [DRL13, GGJ04]. Clusters [LWW16, BCHL05, BDV03, JR10, KK11, Nag05a]. CM [GBJ94, MJ95]. CM-2 [GBJ94]. CM-200 [GBJ94, MJ95]. Co [RN07, ZGW08, WCZ^{+}16, BCS01]. Co-Array [BCS01]. Co-arrays [RN07]. co-regulations [ZGW08]. Coaching [Ano07a]. Coarrays [RRM^{+}15]. coarse [KB96]. coarse-grain [KB96]. Code [BCC^{+}92, BKLS01, CCBPGA15, ACIK97, BHK^{+}13, CP03, Jes10, PKE^{+}10]. Codes [O’K00, BGLR93, IJL^{+}01, Kok07, Mal01, NAA^{+}03]. Codesign [MOS16]. Coding [ATZ^{+}15, FC01]. coherence [WC96]. Collaborating [CL04]. collaborative [FML^{+}14, KJG^{+}08, XNQF04]. collection [Hav00b]. Collections [BBC^{+}10].

Collective [FML^{+}14]. coloring [BÇCD12]. combinatorial [BÇCD12]. Combined [DRR12, HY15, ZLL^{+}15]. Combining [LGC^{+}16, MSH99, MS00]. combustion [SE95]. Common [BD16]. Communication [GBK^{+}96, BDM^{+}04, Bry96, ESSL99, KB01, MJ95, PKE^{+}10, TPKP13]. Communications [GBJ94, GG95]. Community [KCS15, SSN15]. Community-Based [SSN15]. Comparative [ZS99, OB96]. Comparing [FC01, PSM^{+}15]. Comparison [PBK01, ZG15]. Compilation [Ano99b, CP03, ESSL99]. compile [MSH99]. compile-time [MSH99]. Compiler [BMN^{+}97, DS97, DZKS13, IRSD99, OSS94, SKS01, Sin99, WC96, Ang93, BGH99, EAS^{+}97, HCJ08, Mül03, OB96, SHHI01, SO93, Ben99]. Compiler-directed [DZKS13]. Compiler-enforced [WC96]. Compilers [CCTY15, Met99a, Met99b]. compiling [DDS99, GG95]. Complete [Nag05b]. complex [Che93a, DSS^{+}05, ESSL99]. Complexity [Rou08a, ANEA08, PPS12a, PPS^{+}12b]. component [KCO^{+}05, KJG^{+}08, Sch03b]. component-based [Sch03b]. Components [ADLN08, ZCXQ16, CL04]. comprehensive [RST02]. compressible [WB95]. compression [IMB^{+}13, LLW^{+}14]. Computation [Nag05b, SRO^{+}15, DSZ96, HST^{+}93, IB09, KB01, Lar93, OSS94, PMTL12, PTS^{+}13, PKE^{+}10, WGF93]. Computational [KAS15, SRTCL16, GW11, GHW00, HLM^{+}05, MCvM10, WD07, SLJ^{+}00]. computations [CLM05, CFR14, HUN08, HJYC10, PCS99, ZSS^{+}10]. compute [JIC^{+}14]. compute-local [JIC^{+}14]. computer [BTS^{+}14, IJL^{+}01, JEM07, KMB09, LDV07, Ste97, WCKD07, WSP94]. Computers [PAR94, HB94, LH93, OS99, PBK01, PMCF94, Shu94]. Computing [ALL^{+}16, BR11, CTC^{+}15, DK02, DDMS15, GM14, GGP09, HPZ16, Nag04, PSM^{+}15, PT09a, PT09b, VR13, WWLG16, YTD15, ACB^{+}02, Ano08b, AB96, BKK^{+}11, BH12, BDG^{+}94, BBK^{+}11, BC99, BÇCD12, BDH^{+}10, CJS^{+}02, CM02, Che93b, DDMS14, DRR12, FB99, FP00b, GW11, GGMS99, GHH^{+}02, Hav00a, JIC^{+}14, KeKR^{+}11, KLS^{+}96, LPHD04, LG03, MMG^{+}02, NJ08, Nor07, NDSG07, Rou08a, SKU^{+}09, VCT05, VF95, WGW08, Zim07, HHML05, JEM07, SO11, Mic97]. concept [BMY13]. concept-based [BMY13]. Concepts [DNS97, NDSG07]. Concurrent [BBC^{+}10]. Conditional [BCK07]. Conference [GM14]. Confidence [ÁRCPS15]. Config [MS00]. configuration [CQF05]. conjugate [GG05]. CONLAB [JKR92]. Connection [GBJ94, MJ95]. connectivity [PSU08]. conquer [MGK^{+}13]. considerations [GHST12]. Consistency [ZLL^{+}15]. Consolidation [CLYL16]. constant [CGL08]. constant-time [CGL08]. Constrained [MFB^{+}15, SZC16, YZ16, LBvBW12]. Constraint [SW16]. Constraints [BD16, LR05]. Constructing [BBBM12]. construction [MOT97]. constructs [CP95, RS94]. continuum [ADLN08]. contour [RLC04]. contraction [PMCF94]. Control [HXX16, FP95, GR93]. convergence [HMCH07]. Converter [VMRM16]. cooperative [CR05]. Coordinate [Bjø00, GHW00]. Coordinated [PRM^{+}14]. Coordination [CHM^{+}97]. coping [Hil97]. Coprocessor [SRO^{+}15]. Coprocessors [TSP^{+}15]. CORBA [Lan01]. Core [CKKK15, TC96, TLC15, JR10, JIC^{+}14, PTS^{+}13, PT09a, PT09b, Zhe10, DGH^{+}15, Nag05b]. coregulation [HB12]. Corral [JDVM10]. correlation [KMB09]. cosmological [War14]. cosmology [BAN02]. Cost [MFB^{+}15, Wu16, BKP99, JMR^{+}11, PMCF94, SZ04]. Cost-Effective [Wu16]. coupling [GRN99]. course [ZGW08]. courses [Met99a, Met99b]. CPU [CFR14]. CPU/GPU [CFR14]. CPUs [LNK^{+}15, YTD15]. CRAFT [PMM94]. cratering [HLM^{+}05]. CRAUL [IRSD99]. Creating [BKLS01]. Critical [NM16, ZA10]. cross [BHK^{+}13, GA96]. cross-architecture [BHK^{+}13]. crunching [GB94]. crystals [GIKP95]. CSE [CHC15]. Cuckoo [WS15]. custom [NAA^{+}03]. customizable [WOC99]. Customized [XST^{+}16]. CX [CM02]. Cycle [ZCXQ16].

D

[Nag05a, JR10, MRLF12, PPJ^{+}15, TPKP13]. Data [AlS16, CTCG15, CTC^{+}15, CSC16, CGC16, HJ15, HMC97, JFPL16, KPS16, LR05, LZF16, MJB15, SZXL16, SSM^{+}02, THS^{+}15, WCG95, Wu16, ZHL16, AMN^{+}12, AHCZ12, AGG^{+}97, CDD^{+}05, CvHK97, CP95, DRT^{+}14, Gil09, Hav00b, HJ96, HS03, IB09, IMB^{+}13, KG08, KB96, KGV97, KVW^{+}07, LOHA01, LLW^{+}14, Lin04, NPP^{+}00, Özt04, PSU08, PDGQ05, Pla04, PMCF94, RR07, RKR^{+}99, Rou08b, SVR^{+}07, Sin99, Slo12, SG96, TFN11, VRW^{+}03, VRM02, WHRH07, War96, ZRP08, ZS99, ZGW08, ZSS^{+}10, vDKH01]. data-aware [IMB^{+}13]. data-intensive [DRT^{+}14, Slo12]. Data-parallel [HMC97, WCG95, CvHK97, Hav00b, HJ96, ZS99]. data-structure-neutral [SG96]. Database [CTC^{+}15, HMSW92, BDLL94, MSLS13, WSZK09]. databases [TSCT11]. Datasets [SZAG15, BCHL05, HB12]. David [Ano07c, Nag05b]. DDT [AGG^{+}97]. de-allocation [RMX05]. Deadline [MFB^{+}15]. Deadline-Constrained [MFB^{+}15]. deadlock [HPS^{+}13]. debugging [FP00b]. Decision [MWQ16]. Decisions [ZGS16]. decomposition [DBVF01, GRN99, JHNP14]. Deduplication [CSC16]. definition [HS03]. definition-use [HS03]. Degree [ZCXQ16, ZWW16]. Delays [HXX16]. Delegated [ITF^{+}08]. Demand [ZJZ^{+}16, Ang93]. Demands [SLBZ16]. Dense [YTD15]. dependence [BH02, Lin04]. Derivative [BCC^{+}92]. derivatives [Met99a, Met99b]. derived [RMX05]. descriptions [MPP^{+}04]. Design [CFR14, CDO^{+}96, CIN^{+}96, GHST12, GSM03, HMR^{+}15, O’K00, PPD05, RX04, SS00, SG96, SV15, ZWW16, DRT^{+}14, DSZ96, FP95, PDA^{+}08]. Designing [HPZ16, Lin04]. designs [OHS00]. Detection [BD16, CLA^{+}15, HJ15, KCS15, XML^{+}16, HPS^{+}13]. Determine [NM16]. Determining [ADS95, BH02]. Development [AKA16, Ano07a, FS01, HLM^{+}05, PT09a, PT09b, SB01, VHBR93, ARvW03, Ano93, JKR92, KJG^{+}08, SDS00]. Diagrams [NCMF15]. Difference [Vol97, VHBR93]. Different [GMCT15]. differential [Gus93, PPS^{+}12b]. digit [BSF96]. digit-index [BSF96]. dilemma [BFH96]. dimensional [Kok07, KBRS95, WHG93]. Dimensionality [SZAG15]. Direct [BHRT12]. directed [DZKS13]. Direction [BAP13]. Direction-optimizing [BAP13]. Directive [XTCC15]. Directive-Based [XTCC15]. Directives [CCBPGA15, LOHA01]. Discover [ZGW08, HB12]. discoveries [Gil09]. Discovering [DPMB16]. Discovery [GB15, AHCZ12, MPP^{+}04, WD07]. Distributed [BBG^{+}93, BMN^{+}97, KPS16, SO11, SO15, BKK^{+}11, DSS^{+}05, FP00a, FS01, FB99, HMCH07, JKR92, KMR^{+}97, KGV97, KABW11, LTB02, MSLS13, OB96, RLL^{+}02, RS95, SHHI01, Shu94, XNQF04, ZA10]. distributed-object [FS01]. Distributing [ASCH^{+}07]. Distribution [MJB15, ZWW16, ACIK97, AGG^{+}97, CP95, KK09, NPP^{+}00]. distributions [KTP05]. divide [MGK^{+}13]. DMZ [DRT^{+}14]. DNA [KTP05]. do [Ang93, GA96]. Domain [JBLB15, GRN99, JHNP14]. Dongarra [Nag05b]. Driven [BR11, Kes96, KVW^{+}07, Rou08b]. DSMPI [SSC97]. DSP [YZZ15]. Dual [ACCTA^{+}15, ZGS16]. Dual-Source [ZGS16]. during [AKA16]. DWD [SDS00]. Dynamic [CMM^{+}02, CP03, GWZ16, HPZ16, HLPY16, KLN^{+}04, LTB02, RMX05, WD07, WHRH07, HS03, MSCS14, RKR^{+}99, Shu94, VCT05]. dynamical [CO93]. dynamics [BCS01, BPK00, GIKP95, GHW00].

Early [HLM^{+}11]. EC2 [TFN11]. ECMWF [DTV00]. Ecology [WCZ^{+}16]. ed [Nag05b]. edited [Bry96]. Editor [Sne95]. Editorial [Ger02, O’K00]. editors [Pre99]. education [DDMS14]. EEG [ADS95]. Effect [AKA16]. Effective [TSP^{+}15, Wu16]. effects [CBCM93, GIKP95, LLW^{+}14]. Efficiency [CGK^{+}05, JFPL16, Mat94, STD^{+}14]. Efficient [CTCG15, CSM15, DRL13, HJ15, IB09, MOS16, PTS^{+}13, RLC04, SCAP16, TT16, CGL08, GL04, KABW11, SZ04, TFN11]. Efficiently [RLL^{+}02, HJ96]. Eigensolver [BH96]. Ejector [CCBPGA15]. ELASTIC [MSCS14]. Elasticity [ALL^{+}16, Kok07]. Electric [WCKD07]. electronic [CL04]. Element [Ano07b, JBLB15, SO15, VHBR93, WGF93, ZSS^{+}10]. elements [Hem00]. Elimination [PPJ^{+}15]. elliptic [AGIS94, JHNP14]. Ellis [Bry96]. Embedded [JBLB15, YZZ15, PPS12a, PPS^{+}12b]. Empirical [CTCG15, HBCM94]. Emulating [Mor15]. enabled [CDD^{+}05, KVW^{+}07, SHHI01]. Enabling [BTS^{+}14, GBH14, HJYC10, Mic97, vDKH01, Nor07]. encoding [LSB12]. Energy [CLYL16, JFPL16, TT16, ZHL16, CKS^{+}09, PRM^{+}14, STD^{+}14]. Energy-Aware [CLYL16]. Energy-Awareness [ZHL16]. Energy-Efficiency [JFPL16]. Energy-Efficient [TT16]. enforced [WC96]. engine [KVW^{+}07, NPP^{+}00, GGP09, IB09, KD09, SKU^{+}09, VGC09, WSZK09]. engineer [Hil97]. Engineering [Ano07a, CHC15, Nag05b, HP02, MS00]. Enhancing [FMA16]. ensemble [AMN^{+}12]. Entire [YZZ15]. Enumeration [SJ15]. Environment [CTC^{+}15, GBK^{+}96, BBF^{+}04, BGLR93, BDG^{+}94, BBLR03, CB99, FS01, HMC97, Lan01, Lar93, MSCS14, PVKE01, Pro07, VF95]. Environments [OCC^{+}16, EAS^{+}97, GGMS99, GRC03, RLL^{+}02, VCT05, VRW^{+}03, ZA10, Bry96]. epidermal [GGJ04]. Epoch [HJ15]. Epoch-Based [HJ15]. ePRO [CKS^{+}09]. ePRO-MP [CKS^{+}09]. Equalization [VMRM16]. equation [MRLF12]. Equations [YTD15, Gus93, PPS^{+}12b]. error [HPS^{+}13]. estimation [TOF^{+}14]. estimator [FP00a]. estuary [KBRS95]. Evaluating [HY15, SZ09]. Evaluation [LNK^{+}15, SCD^{+}15, BC99, ZS99]. Event [HXX16, Kuf15, THS^{+}15, OPP11]. event-based [OPP11]. Event-Triggered [HXX16]. events [ZA10]. Evolution [Bor04]. Example [ZG15, Den96]. Execution [Mar05, FS01, HLM^{+}05, Slo12, ZS99]. executions [RR07]. Expands [WCZ^{+}16]. Expected [YZ16]. Experiences [CvHK97, JR10, JDVM10, Lar93]. experiment [BCCP05]. experiments [BKK^{+}11, PDA^{+}08]. Expert [OB96]. Explaining [SH94]. Explicit [SF03b, SF03a]. exploit [BPC^{+}09]. Exploiting [HPD09]. Exploration [SV15]. Exploring [JK10, JIC^{+}14, MRLF12]. Express [DNS97]. expressing [JK10]. expression [PCS99, ZGW08]. Extended [TC96]. Extending [BCC^{+}00]. extension [PPD05]. Extracting [NCMF15]. extraction [Abd02]. extreme [BAN02].

Fact [CSM15]. factor [GGJ04]. factorial [PDA^{+}08]. Factorization [DEŁ01, CDO^{+}96, KD09, VGC09]. factory [JMR^{+}11]. fair [BTS^{+}14]. Fast [BSF96, JFPL16, SJ15, HST^{+}93]. Fastscat [HST^{+}93]. fault [ZA10]. fault-tolerance [ZA10]. Faults [HXX16]. Feature [CLTX16, KS15]. Feedback [MPA16]. Feedback-Based [MPA16]. FFT [Sch03b, TPKP13]. field [VHBR93]. Fields [ACCTA^{+}15]. file [DZKS13, Met99a, Met99b]. files [CGL08]. Filtering [FGROL^{+}16]. Final [ZGS16]. financial [LPHD04]. fine [HPD09]. fine-grain [HPD09]. Finite [Ano07b, JBLB15, Vol97, VHBR93, WGF93, ZSS^{+}10]. Fink [Hil97]. first [BAP13]. Fitting [VRM02]. five [GA96]. Flexible [ESSL99, RS94, GHST12, MSLS13]. flood [MCvM10]. Flow [CCBPGA15, ADLN08, FP95, GR93, WB95]. fluid [GHW00]. Fly [HJ15]. footprint [SVR^{+}07]. force [AJ94]. forecast [DTV00, Ros00]. forecasting [WCG95]. Foreword [Ano05, CK08]. format [CGL08]. Formation [SCAP16, BAN02]. Formula [AJJF14]. formulation [KHSJ95]. Fortran [Ben99, Ano93, DDS99, Zim07, ACIK97, AJJF14, AGG^{+}97, BCS01, BCC^{+}92, BMN^{+}97, CMZ92, DS97, Hig93, HMR^{+}15, HZ99, KMR^{+}97, KOM94, KGV97, LP99, McC96, MH95, Met99a, Met99b, Mor15, NCMF15, NDSG07, OPE^{+}95, PMM94, PCS99, RRM^{+}15, RN07, RMX05, Sch03a, Sny07, Szy07]. Fortran-like [KGV97]. Fortran-P [OPE^{+}95]. FORTRAN/journal [Ano93]. Fortran90 [DNS97]. Fortran95 [DN04]. ForTrilinos [MRLF12]. ForUML [NCMF15]. Forward [ACCTA^{+}15]. FP [JFPL16, ZYLZ15]. FP-ABC [JFPL16]. FP-Growth [ZYLZ15].

FPGA [OCC^{+}16]. FPGA-Aware [OCC^{+}16]. Fractional [PDA^{+}08]. Fragmentation [GWZ16]. Framework [BGLR93, CKKK15, FGROL^{+}16, Pla04, SZAG15, SW16, SZXL16, ACB^{+}02, ACIK97, BDM^{+}04, BBK^{+}11, BHK^{+}13, CR05, DSS^{+}05, MCvM10, OPP11, TPKP13]. Framework-based [BGLR93]. Free [ZWW16, Bjø00, GHW00]. friendly [Hem00]. Fujitsu [DTV00]. function [KMB09]. Functional [BH96, DFP96, Den96, WC96]. Fusion [MWQ16, PKE^{+}10]. future [JIC^{+}14]. Fuzzy [NM16, SZC16].

G [ABB^{+}02, KVW^{+}07, TF04]. gamma [BCHL05]. gamma-ray [BCHL05]. gap [SH94]. GCR [CKKK15]. GDMP [SSM^{+}02]. Geant4 [SCD^{+}15]. gene [ZGW08]. general [Hem00, Mat94, RR07, SS00, SF03b, SF03a]. Generalized [CBRRC^{+}16]. Generating [SO15, BCC^{+}92]. generation [BDM^{+}04, BHK^{+}13]. generator [SF03b, SF03a]. generic [BH12, PPS12a, VHBR93]. Genetic [CLTX16, KS15]. gentle [Ste97]. GFDL [Hem00]. ghosting [MSLS13]. Giga [GKL^{+}96]. Global [AKA16, KGMN15, Ros00, WB16, FC01, HMCH07, Wan02]. GMA [KCO^{+}05]. GMRES [PBK01]. Gockenbach [Ano07b]. Good [Ano07c, Mat03]. Gordon [Per08]. governed [BH02]. GPU [CFR14, CCBPGA15, GW11, Jes10, XTCC15, YTD15]. GPUs [BBDN11]. gradient [GG05]. grain [HPD09, KB96]. Granularity [WS15]. Graph [Lin04, PMCF94]. Graphical [PHH95]. graphics [BB09]. Gravitational [ACCTA^{+}15]. Great [Ste97]. Grid [DK02, KCO^{+}05, ACB^{+}02, CJS^{+}02, GRC03, MPP^{+}04, Pla04, HP02, ABB^{+}02, ACB^{+}02, AMM05, BBF^{+}04, BBŁ08, COB^{+}00, CQF05, HHML05, HLM^{+}05, KFFZ05, KVW^{+}07, KLN^{+}04, MAG^{+}07, PPD05, Pro07, Rom02, SZ04, SSM^{+}02, TF04, VCT05]. Grid-based [ABB^{+}02]. Grid-enabled [KVW^{+}07]. GRIDCC [MAG^{+}07]. GridLab [KLN^{+}04]. Grids [PPJ^{+}15, SN02, ITF^{+}08, Özt04, SLJ^{+}00]. GridWay [HHML05]. Grinding [WS15]. GROMOS96 [BCS01]. Groovy [PSM^{+}15]. GroovyLab [PSM^{+}15]. Group [DPMB16]. Growth [ZYLZ15, GGJ04, Szy07]. Grunwald [Bry96]. GSSIM [BKK^{+}11]. GTS [PKE^{+}10]. Guest [Ger02, O’K00, Sne95]. Guest-Editorial [Ger02, O’K00]. Guide [Ano07c, Hil97]. guidelines [GB94].

H [Hil97, Mic97]. Hadoop [SZXL16]. Handling [Che93a]. Hardware [DGH^{+}15, MOS16, vDKH01]. Hardware/Software [MOS16]. hashed [War14]. HBase [CTCG15]. help [BFH96]. HeNCE [BDG^{+}94]. heterogeneity [LR05]. Heterogeneous [BB09, CCTY15, HMSW92, LWW16, SRO^{+}15, TT16, WWLG16, BCHL05, BDG^{+}94, BBK^{+}11, BDH^{+}10, DRL13, JHNP14, KK11, PVL^{+}04, PRM^{+}14, RLL^{+}02]. Heuristic [LWW16]. Heuristics [KGMN15, SLBZ16, FB99]. Heuristics-Based [SLBZ16]. Hidden [XML^{+}16, BH02, LB02, VRM02]. hierarchical [BDV03, DZKS13]. Hierarchically [YTD15]. hierarchy [BPC^{+}09]. High [Ano99b, Ano08b, CTCG15, GM14, GGP09, GBK^{+}96, HMR^{+}15, HKL^{+}12, LBvBW12, Nag05a, Nor07, WSZK09, AB96, BTS^{+}14, CB99, FP00b, GGMS99, IR02, KMB09, Lan03, LG03, MMG^{+}02, NJ08, VvAC^{+}09, Zim07, ACIK97, Ano93, AGG^{+}97, BMN^{+}97, DS97, KMR^{+}97, KOM94, KGV97, MH95]. High-level [LBvBW12]. High-Performance [GBK^{+}96, HMR^{+}15, HKL^{+}12, AB96, FP00b, KMB09, LG03]. high-resolution [VvAC^{+}09]. highly [GBH14]. highly-scalable [GBH14]. HMM [Abd02]. HMM-based [Abd02]. hoc [BDM^{+}04]. Hole [YDC^{+}16]. homogeneity [KTP05]. host [Din99]. HPC [BFGH14, DGH^{+}15, STD^{+}14]. HPF [BDV03]. HSIP [WWLG16]. Humphrey [Ano07a]. Huss [Nag05b]. Hybrid [ÁRCPS15, ATZ^{+}15, DFP96, KS15, SW16, SZC16, WS15, CFR14, JR10, KeKR^{+}11]. Hypervisor [OCC^{+}16]. Hysteresis [WS15].

I/O [OSLK12, NsP16]. IBM [KGBB09, Was95]. Ideas [BBG^{+}93, Ste97]. Identifying [ZCXQ16]. IDL [ESSL99]. IFS [DTV00]. II [PPS^{+}12b, Per08]. image [GKL^{+}96, RLC04]. images [VvAC^{+}09]. impact [FSV14, HLM^{+}05]. imperative [DFP96]. Implementation [BHP^{+}03, CTC^{+}15, KMB09, NsP16, SKU^{+}09, SSC97, TOF^{+}14, BS01, BDV03, CBCM93, CDO^{+}96, CIN^{+}96, CW93, FS01, Gus93, PPD05, PCS99]. implementations [ADS95, BH96, PBK01]. implemented [Hav00b]. Implementing [Ano07b, CS94, HJ96, Lan03, VGC09, YZZ15, DN04]. implications [DF14]. implicit [SCSJ09]. Improve [AlS16, GB15, Sin99]. Improved [ZYLZ15, War14, vDKH01]. Improving [BBDN11, RR04]. Impulse [CHS^{+}99]. Incorporating [BGH99]. incremental [TOF^{+}14]. independent [GL04]. Index [Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14, CTC^{+}15, XML^{+}16, Ano97a, BSF96]. indicator [FSV14]. Individual [GGJ04]. Individual-based [GGJ04]. Indoor [LWYS16]. Industrial [Nag04, ZCXQ16]. Inferring [PSU08]. inflation [OdSSP13]. influence [SCSJ09]. influences [RX04]. influenza [LSB12]. informatics [BFH96]. Information [AlS16, Met99a, Met99b, PHH95]. Infrastructure [CMTAC^{+}16, SPPH15, KeKR^{+}11, Rom02, SGM^{+}08].

Infrastructure-Based [SPPH15]. Inheritance [Mor15]. input [GB94]. Insights [WCZ^{+}16]. Instruction [GR93]. instrument [KCO^{+}05]. Integrated [DGH^{+}15, RKR^{+}99]. Integrating [FMA16]. integration [IRSD99]. Integrity [HHD^{+}16]. Intel [CBCM93, DGH^{+}15, SCD^{+}15, SRO^{+}15, TSP^{+}15]. Intelligence [NM16, SRTCL16, WCZ^{+}16, Gil09]. Intelligent [GBK^{+}96, SZC16]. Intel(R) [BGH99]. intensive [DRT^{+}14, Slo12]. Inter [ITF^{+}08]. Inter-operating [ITF^{+}08]. interactive [BBLR03, KVW^{+}07]. interchange [VRW^{+}03]. Interest [DPMB16, MGK^{+}13]. Interface [SLD^{+}16, Spo12, WOC99]. Interfaces [BKLS01]. interference [SCSJ09]. intermediate [Tým99]. International [GM14, PAR94, SO11, HB94]. Internet [WCZ^{+}16]. Interpreting [PDGQ05]. Interprocedural [HBCM94, HS03]. intervals [CGL08]. Intrepid [BEK^{+}12]. Introduction [Ano01, Ano02, Ano04a, MH95, PT09a, Sne95, VC93, WSP94, Ste97]. Inventory [ZGS16]. inverse [Wan02]. inverses [GG05]. Investigating [HMCH07, OPP11]. iPSC [CBCM93]. iPSC/860 [CBCM93]. irradiation [GIKP95]. Irregular [KB96, PCS99, LOHA01, LP99]. Isorropia [BÇCD12]. Issue [JK10, PT09a, PT09b, SO11, VR13, BFGH14, BR11, GM14, HR12a, HR12b]. Issues [Ano99b, McC96, ARvW03]. Itanium(R) [GHH^{+}02]. iterative [BHRT12, DEvdV01, SG96].

J [Hil97]. Jack [Nag05b]. Jacobi [BH96]. JADE [CGK^{+}05]. Janet [BKLS01]. Jason [PCGLO14]. Java [Ano99b, BC99, BKLS01, DDS99, GRN99, GGMS99, MR02, MMG^{+}02, PMTL14, Tým99, WOC99]. Java-based [GRN99, WOC99]. Java^{T M} [BGH99]. JIST [ARFS05]. JIT [BGH99]. JLAPACK [DDS99, PMTL12]. job [FS01, KLN^{+}04]. job-execution [FS01]. Joseph [Nag05a]. journal [Ano93]. Jr. [Hil97]. Just [ARFS05]. Just-In-Time [ARFS05].

Kaiser [Bry96]. Kemari [KMR^{+}97]. kernels [FSV14]. Key [Hem00]. Knowledge [HMSM08, YG16, AHCZ12, Gil09]. Kokkos [ESP^{+}12]. Komzsik [Bor04]. Koniges [Nag04]. KSR [KBRS95]. KSR-1 [KBRS95].

laboratory [ABB^{+}02]. laminar [FP95]. Lanczos [Bor04]. Landau [Hil97]. Language [CHM^{+}97, CP95, Hig93, JBLB15, AJ94, AMM05, AJJF14, BBG^{+}93, CLM05, Che93a, Che93b, DSZ96, GGMS99, LG03, Met99a, Met99b, RS94, Szy07, VRW^{+}03, VF95, WC96]. Languages [CCTY15, HJ96, JK10, KGV97, Mar05, Zim07]. LAPACK [ARvW03, DDS99]. Large [SZAG15, SZXL16, TSCT11, BHRT12, BPK00, CGL08, JEM07, MJLM07, MSCS14, RS94, WGW08]. large-area [JEM07]. Large-Scale [SZAG15, MJLM07, WGW08]. Latency [Kuf15, RS95]. Lattice [IB09]. layout [DZKS13]. leadership [WGW08]. Learning [PDROFRM16, KG08, LB02, PVL^{+}04]. Lederman [Nag05b]. Legacy [RRM^{+}15]. LEO [CCBPGA15]. Letter [Pre99]. Level [OCC^{+}16, Vol97, COB^{+}00, LBvBW12, RRV09]. libraries [ARvW03, PMTL12, PMTL14, SG96]. library [BŻvA^{+}01, ESP^{+}12, GHST12, GL04, GSM03, Lan03, VGC09]. life [KFFZ05]. Lifting [Sch03b]. lightweight [Ott93]. like [KGV97, PMTL14]. Linda [Lar93, Mat94]. line [BBF^{+}04, RLL^{+}02, SCB02]. Linear [ACIK97, WB16, YTD15, BHRT12, HKL^{+}12, Kok07, MSSG11, RR07, SG96]. Linearization [WB16]. Link [CSM15, GMCT15, SSN15]. linked [HS03].

Linpack [KGBB09]. Linux [Nag05a, HHD^{+}16]. Live [JFPL16]. Load [CKKK15, Din99, IRSD99, LTB02, PVL^{+}04]. Local [LGC^{+}16, FSV14, JIC^{+}14, ZGW08]. Locality [MJB15, TLC15, HJYC10, RR04, Sin99, Zim07]. Locality-Aware [MJB15, HJYC10, Zim07]. Locality-Centric [TLC15]. Locations [Kuf15]. Logging [ZLL^{+}15]. long [KTP05]. lookup [WSB11]. loop [NAA^{+}03]. loops [LOHA01]. Louis [Bor04]. Low [RS95, YTD15, HB12, SZ04, TPKP13]. low-communication [TPKP13]. low-cost [SZ04]. Low-Rank [YTD15]. low-variance [HB12]. LPARX [KB96]. LU [CDO^{+}96, DEŁ01]. Lung [AMN^{+}12].

Machine [AGIS94, CGC16, CLTX16, Hav00b, MJ95, PDROFRM16, SCAP16, ZHL16, GBJ94]. Machine-Part [SCAP16]. Machines [BMN^{+}97, SJ15, SO15, BCC^{+}93, BCC^{+}00, HMCH07]. MAGMA [DGH^{+}15]. magnetic [PKE^{+}10]. make [Sch94]. Management [CMTAC^{+}16, THS^{+}15, YG16, BCK07, CJS^{+}02, IB09, KeKR^{+}11, KABW11, KLN^{+}04, PRM^{+}14, vDKH01]. Managing [ANEA08, AKA16, CDD^{+}05, PPS12a, PPS^{+}12b]. Many [CKKK15, TLC15, PTS^{+}13, DGH^{+}15]. Many-Core [CKKK15, TLC15, PTS^{+}13]. Many-Integrated-Core [DGH^{+}15]. Manycore [CCTY15, ESP^{+}12, MJB15]. Maple [PPD05]. mapping [DSS^{+}05, Den96, DRR12, PMCF94, SZ04]. MapReduce [MPA16]. MapReduce-Based [MPA16]. maps [KTP05]. Marc [Nag05b]. Mark [Ano07b]. Markov [BH02, LB02, VRM02]. markup [VRW^{+}03]. Mary [Bry96]. massive [JK10]. Massively [GA96, Nag04, LH93, OPE^{+}95, SS00, Sch94, WMR^{+}94, WB95, WSP94]. Matching [TLC15]. MatchMaking [ITF^{+}08]. Materials [HMSW92, NKV^{+}02].

Mathematical [LGC^{+}16]. Matheuristic [LGC^{+}16]. MATLAB [PMTL14, CB99, Kok07]. MATLAB-like [PMTL14]. Matrices [CSM15]. Matrix [DEŁ01, VMRM16, YTD15, BBDN11, CFR14, KHSJ95, Lin04, VGC09]. Maximizing [HY15]. May [Met99a, Met99b]. MC2 [DTV00]. McrEngine [IMB^{+}13]. means [RLC04]. Measurement [ALL^{+}16, WGW08]. Measuring [DPMB16, STD^{+}14]. mechanics [DFP96, GW11]. Mechanism [THS^{+}15]. mechanisms [Ott93]. medium [JK10]. Memorial [Per08]. Memory [BMN^{+}97, CHS^{+}99, SO15, SZXL16, TKS02, Bai95, BPC^{+}09, DBVF01, FSV14, GL04, GBH14, HMCH07, JKR92, JIC^{+}14, KMR^{+}97, KHSJ95, LR05, NAA^{+}03, OB96, Per08, PBK01, RS95, RMX05, SHHI01, SZ09, Shu94, ZS99]. Memory-Based [SZXL16]. Menhir [CB99]. mental [ADS95]. Mesh [KK09, BAN02, CL04, CW93, MSLS13, KK09]. mesh-based [CL04]. Meshes [SO15]. message [PBK01]. message-passing [PBK01]. messages [RS95]. MESSIAHS [CS94]. Messina [Mic97]. metadata [MPP^{+}04]. metaprogramming [BMY13]. Method [ÁRCPS15, Bor04, HY15, HHD^{+}16, MM16, NM16, TC96, WB16, YDC^{+}16, DN04, GG05, PBK01, TOF^{+}14, Ano07b]. methodologies [Hav00a]. methodology [HCJ08]. Methods [CKKK15, GMCT15, Mor94, Vol97, BCC^{+}93, RR07, STD^{+}14]. Metric [LWYS16, XML^{+}16]. Metrics [GMCT15, LLW^{+}14, RX04]. MHDFS [SZXL16]. MIC [CCBPGA15]. micro [ADLN08]. micro- [ADLN08]. MicroGrid [SLJ^{+}00]. Microimmune [YZ16]. middleware [COB^{+}00, VCT05]. migratable [BŻvA^{+}01]. Migrating [SCAP16]. Migration [CCBPGA15, GWZ16, LZF16, HZ99, KLN^{+}04]. millennium [Met99a, Met99b]. MIMD [PBK01]. mind [FML^{+}14]. Minimum [CBRRC^{+}16]. Mining [AlS16, HB12, KS15, SLD^{+}16, ZYLZ15, AMN^{+}12, AHCZ12, PSU08, ZRP08]. Mirroring [SSM^{+}02]. Missing [WCZ^{+}16]. mission [Nor07]. mission-enabling [Nor07]. mitigating [OdSSP13]. Mixed [LG03, RR07, SB01]. Mixed-language [LG03]. MM5 [Mic00]. MMX^{T M} [BGH99]. Mobile [ATZ^{+}15, MR02, CKS^{+}09]. mobility [BDM^{+}04]. mode [SB01]. Model [BD16, DDMS15, HPZ16, KS15, KPS16, SZC16, XTCC15, BHK^{+}13, BH02, BDV03, CIN^{+}96, DTV00, Hem00, HMCH07, KBRS95, Mar05, PMM94, Ros00, SZ09, SS00, SDS00, SBM^{+}10, Wan02, War96]. Model-Based [BD16]. Model-Centered [DDMS15]. Modeling [DDMS14, SW16, ADLN08, BDM^{+}04, GRR^{+}03, LDV07, SLJ^{+}00]. Modelling [ACCTA^{+}15, LPHD04]. Models [CCTY15, WS15, LB02, MSSG11, VRM02]. Modern [HMR^{+}15, LNK^{+}15, AJJF14, NDSG07]. Modulation [VMRM16]. module [BCS01, Sch03a]. modules [HB12]. molecular [BCS01, BPK00, DFP96, GIKP95]. Monitoring [BBŁ08, Kuf15, BBF^{+}04, KCO^{+}05, TF04]. Monocular [LWYS16]. MOPEX [JEM07]. mosaics [JEM07]. motivated [LLW^{+}14]. MP [CKS^{+}09]. MPDATA [SRO^{+}15]. MPI [Nag05b, GBH14, HZ99, HPS^{+}13, JR10, KB01, RST02, SB01]. MPI-3 [GBH14]. MPI/OpenMP [JR10, SB01]. Multi [CKKK15, GGMS99, PT09a, PT09b, XTCC15, CIN^{+}96, DF14, JR10, PVL^{+}04, Zhe10]. Multi- [CKKK15]. multi-agent [PVL^{+}04]. Multi-core [PT09a, PT09b, JR10, Zhe10]. Multi-GPU [XTCC15]. Multi-language [GGMS99]. multi-model [CIN^{+}96]. multi-objective [DF14]. multiagent [PCGLO14].

Multibillion [SO15]. MultiCache [NsP16]. multicomputers [JKR92]. Multicore [BD16, KGMN15, LNK^{+}15, NJ08, SJ15, YTD15, BB09, HPD09, LKDB10, PCGLO14, SZ09, SJR14]. multidimensional [ESP^{+}12, FO96]. Multidisciplinary [CHM^{+}97]. multidisk [GKL^{+}96]. multigrid [AGIS94, GHST12]. Multilayered [NsP16]. Multilevel [MFB^{+}15, JJY^{+}03]. multiphysics [PPS12a, PPS^{+}12b]. Multiple [CLA^{+}15, Kuf15, MWQ16, Mor15, WS15, MGK^{+}13]. multiplication [KHSJ95]. multiplications [BBDN11, GA96]. Multiplicative [WB16]. multiprocessor [CKS^{+}09, GKL^{+}96]. MultiProcessors [BDV03, LR05, RS95, AGIS94]. Multiprogramming [BHP^{+}03]. multiscale [ADLN08]. Multisupercomputer [LZF16]. Multitasking [ATZ^{+}15]. Multithreaded [SCD^{+}15, HMC97]. multithreading [Sin99]. MUST [HPS^{+}13]. mutual [Szy07].

NaaS [XST^{+}16]. naming [Sch94]. nanofluidic [ADLN08]. NAS [SBM^{+}10]. Native [BKLS01]. navigation [MJLM07]. Navy [Ros00]. nearly [CGL08]. NEC [DTV00]. Nested [BS01]. Net [HPZ16]. Net/SLE [WOC99]. Netherlands [MCvM10]. Network [DPMB16, GMCT15, HLPY16, KAS15, Kuf15, XST^{+}16, ZCXQ16, ZWCY16, ASCH^{+}07, BDG^{+}94, CM02, CDD^{+}05, CIN^{+}96, DRT^{+}14, DRL13, PSU08, WOC99]. Networking [GM14]. Networks [ÁRCPS15, CGC16, CSM15, GMCT15, SSN15, ZWW16, ADS95, BDM^{+}04, DRL13]. neural [ADS95, CIN^{+}96]. neural-network [CIN^{+}96]. neuronal [PSU08]. neutral [SG96]. next [RN07]. Nimrod [PDA^{+}08]. Nimrod/E [PDA^{+}08]. NINJA [MMG^{+}02]. NoC [MM16]. Node [XTCC15, BTS^{+}14, JR10]. non [JIC^{+}14, NAA^{+}03]. non-regular [NAA^{+}03]. non-volatile [JIC^{+}14]. nonaligned [AGIS94]. nonlinear [BCHL05]. normalized [GG05]. NoSQL [CTC^{+}15]. Notation [McC96]. Novel [FGROL^{+}16, HPZ16, LWYS16, WWLG16]. NUMA [BCC^{+}00, MJB15, TKS02]. NUMA-based [TKS02]. Number [GB94, NM16]. Number-crunching [GB94]. numbers [Che93a]. numerical [BC99, BAN02, DSZ96, DHH00, MMG^{+}02, VF95, WCG95]. numerics [Bjø00]. NumPy [AJJF14]. NVRAM [ZLL^{+}15].

O [OSLK12, NsP16]. Object

[BCC^{+}93, CW93, Gus93, HMSW92, MOT97, NCMF15, BGLR93, BBG^{+}93, BKP99, FS01, HST^{+}93, KZRR94, KK11, VF95, WGF93]. object-based [KK11]. Object-Oriented [HMSW92, NCMF15, BCC^{+}93, CW93, Gus93, MOT97, BGLR93, BKP99, HST^{+}93, KZRR94, VF95, WGF93]. objective [DF14]. ObjectMath [VF95]. objects [MR02]. observations [HLM^{+}11]. observed [BH02]. ocean [HMCH07, Mal01]. oct [War14]. oct-tree [War14]. Octagonal [NM16]. ODE [RR04]. Oligonucleotide [KTP05]. Oliveira [Ano07c]. OMPM2001 [AE03]. on-chip [JK10]. on-line [BBF^{+}04, RLL^{+}02, SCB02]. On-the-Fly [HJ15]. One [GBH14]. Online [LWYS16]. onto [DSS^{+}05]. OON [VC93]. OON-SKI [VC93]. open [SGM^{+}08, SGM^{+}08]. OpenACC [CCBPGA15]. OpenCL [FSV14, LNK^{+}15, THS^{+}15]. OpenMP [ARvW03, BHP^{+}03, BCC^{+}00, BS01, BDV03, CCBPGA15, DBVF01, HJYC10, JJY^{+}03, JR10, KB01, LOHA01, Mal01, Mar05, Mat03, MJB15, Mül03, NPP^{+}00, PVKE01, PKE^{+}10, SHHI01, SKS01, SB01, Wan02]. Operating [WSP94, ITF^{+}08, KZRR94, SCSJ09, WMR^{+}94]. Operational [Ros00]. operators [GRC03]. Opinion [KS15]. Optimal [ACLS16, LZF16, PPJ^{+}15, ZWW16]. optimisation [WD07]. Optimization [CLTX16, FMA16, MFB^{+}15, McC96, SW16, SZC16, SCAP16, WB16, YZ16, ZJZ^{+}16, BBLR03, DHH00, DZKS13, GSM03, LBvBW12, SKS01, Sin99, WSB11, XNQF04]. optimizations [Ang93]. Optimized [CGC16, THS^{+}15, ZLL^{+}15]. Optimizing [BMN^{+}97, CSC16, SVR^{+}07, YZZ15, Zhe10, BAP13, BHK^{+}13, CKS^{+}09, SO93]. Opus [CHM^{+}97]. order [BH02]. ordering [BÇCD12]. ordinary [Gus93]. Orientation [SC16]. Oriented [DPMB16, HMSW92, NCMF15, BGLR93, BCC^{+}93, BKP99, CW93, Gus93, HST^{+}93, KZRR94, MOT97, VF95, WGF93]. Otto [Nag05b]. Out-of-Core [TC96, JIC^{+}14]. Outlier [XML^{+}16]. overcome [BFH96]. Overlapping [KB01, KCS15, PKE^{+}10]. Overview [AJ94, HW12].

P [FP00a, OPE^{+}95]. P2P [ASCH^{+}07]. P4 [Lar93]. Package [SSM^{+}02]. packaging [CL04]. PADDA [Ger02]. Pair [YG16]. Papers [Old00, VR13, Ano97b]. Parallel [CBCM93, CLA^{+}15, CLTX16, CR05, DDMS15, FP95, FOT92, GIKP95, GG05, GBK^{+}96, JFPL16, LPHD04, Nag04, O’K00, Ott93, PAR94, PVKE01, SZAG15, SO11, SJ15, SE95, SO15, SCAP16, TLC15, WGF93, YDC^{+}16, ARvW03, ARFS05, AJ94, ADS95, AGIS94, BCC^{+}93, BBK^{+}11, BBG^{+}93, BB09, BÇCD12, BFH96, CLM05, CM02, CvHK97, DBVF01, DRR12, EAS^{+}97, FP00a, FB99, GG95, GL04, GRN99, GRR^{+}03, GA96, HMC97, Hav00b, HZ99, HB94, HJ96, KMR^{+}97, KZRR94, KLS^{+}96, Lar93, LH93, Lin04, Mar05, Mic00, MSLS13, OPE^{+}95, OdSSP13, OSS94, PVL^{+}04, PMCF94, RR07, RKR^{+}99, RS94, SCSJ09, SS00, SDS00, Sch94, SJR14, SGM^{+}08, SBM^{+}10, Shu94, SCB02, SBJV11, VGC09, WCKD07, War96, War14, WMR^{+}94, WB95, WCG95, WSP94, WGW08, ZS99, JEM07, PDGQ05]. parallel/distributed [FB99]. parallelisation [IJL^{+}01]. parallelism [BS01, HPD09, JK10, KB96, RRV09]. Parallelization [JR10, KBRS95, Vol97, BCS01, JJY^{+}03, Kes96, LP99, MSH99]. Parallelize [RRM^{+}15]. Parallelizing [KCS15, OB96]. Parameter [COB^{+}00, CLTX16, PDA^{+}08, TOF^{+}14]. Parkbench [HB94]. PARMACS [HZ99]. Part [SCAP16, ZGS16, HR12a, HR12b, PPS12a, PPS^{+}12b]. Parthenogenetic [SLBZ16]. partial [PPS^{+}12b]. particle [CBCM93]. Partitioning [BÇCD12, MOS16, CP95, LR05]. Partnership [KJG^{+}08]. passing [PBK01]. Path [LZF16, NM16]. Pattern [GRC03, Kes96, DRT^{+}14]. Pattern-driven [Kes96]. Patterns [HMR^{+}15, CFR14, ESSL99, GSM03, ZS99]. Paul [Hil97, Mic97]. pC [BBG^{+}93]. PCN [FOT92]. PDDP [War96]. PDE [KK09, LBvBW12, SF03b, SF03a]. PDE-constrained [LBvBW12]. PDEs [BEK^{+}12]. PDS [BDLL94]. peak [SH94]. peer [XNQF04]. peer-to-peer [XNQF04]. Pegasus [DSS^{+}05]. penultimate [Met99a, Met99b]. PerfExplorer [HMSM08]. Performance [ACIK97, Ano93, Ano99b, ACCTA^{+}15, AGG^{+}97, BMN^{+}97, CCBPGA15, DS97, DTV00, GB15, GM14, GGP09, GBK^{+}96, HMR^{+}15, HP02, JMR^{+}11, KMR^{+}97, KOM94, KZRR94, KGV97, LNK^{+}15, MJLM07, MSSG11, MH95, Nag05a, SCD^{+}15, SSC97, WGW08, Ano08b, AB96, AE03, BBF^{+}04, BDLL94, BTS^{+}14, BDV03, CLM05, CMM^{+}02, CB99, CKS^{+}09, ESP^{+}12, FP00a, FSV14, FP00b, GGMS99, GRR^{+}03, HCJ08, HLM^{+}11, HKL^{+}12, HMSM08, IR02, KMB09, Lan03, LDV07, LG03, MMG^{+}02, NJ08, Nor07, SS00, SH94, SGM^{+}08, SBM^{+}10, SE95, TOF^{+}14, TF04, WHRH07, WSZK09]. performance-portability [ESP^{+}12]. performance-prediction [BDV03]. performing [PMTL12]. permutations [BSF96]. persistence [CDD^{+}05]. Perspective [DPMB16]. peta [Zim07]. peta-scale [Zim07]. petaflops [Mic97]. PetaShare [KABW11]. Petri [HPZ16]. Pfortran [BCS01]. PGHPF [BMN^{+}97]. Phase [ACLS16, TC96, VMRM16]. Phi [ACCTA^{+}15, DGH^{+}15, SCD^{+}15, SRO^{+}15, TSP^{+}15]. physically [LLW^{+}14]. physics [BHK^{+}13, RX04]. PIC [BDV03]. Pictorial [Bry96]. Pipeline [DDMS15]. Pipelining [ATZ^{+}15]. place [FO96]. Placement [ACLS16, CGC16, ZHL16]. Placements [Wu16]. plasma [LG03]. Platform [GWZ16, XST^{+}16, CGK^{+}05, PCGLO14]. platforms [CFR14]. Playa [HKL^{+}12]. plume [WCKD07]. point [KMB09]. pointer [HS03]. pointer-linked [HS03]. Policy [JFPL16, SZ04]. Polish [ZG15]. polymorphism [DN04]. Polyshift [GBJ94]. Port [DGH^{+}15]. portability [ESP^{+}12]. Portable [BŻvA^{+}01, GL04, KMR^{+}97]. Portal [KBG^{+}02]. Porting [HHML05]. Potential [SLBZ16]. Power [KGBB09]. ppOpen [TOF^{+}14]. ppOpen-AT [TOF^{+}14]. Practical [KOM94, FML^{+}14]. Precise [HBCM94, Mor94]. precision [Sch03a]. preconditioned [GG05]. preconditioners [JHNP14, MOT97]. preconditioning [GHST12]. Prediction [GMCT15, Ros00, SSN15, AMN^{+}12, BDV03, FP95]. predictive [GRN99]. predictive-adaptive [GRN99]. Preface [Ano04b]. Prefiltering [GB15]. preliminary [BKP99]. Pressure [CKKK15]. Pricing [ZGS16, BTS^{+}14]. primordial [BAN02]. Problem [CBRRC^{+}16, GWZ16, GB94, Lan01, VRW^{+}03, WGF93]. Problems [BD16, FMA16, SW16, SCAP16, GHW00, JHNP14, SDS00, VHBR93]. Process [MOS16, BH02, Shu94]. processes [CR05]. processing [BB09, Den96, FP95, TFN11, VF95]. processor [GHH^{+}02, HMCH07, KGBB09, Sin99]. Processors [ATZ^{+}15, MJB15, YZZ15, ARFS05, Ano08b, KMR^{+}97, OPE^{+}95, PTS^{+}13, PRM^{+}14, PCGLO14, Nag04]. Produced [CBRRC^{+}16]. product [GA96, KHSJ95]. Production [ZGS16]. Productive [FOT92]. productivity [Zim07]. Professional [ÁRCPS15]. Profile [SC16]. profiling [BB08, CKS^{+}09]. program [CMM^{+}02, HST^{+}93, RR04]. programmable [HKL^{+}12]. programmer [OB96]. Programming [CCTY15, CMZ92, DS97, DDMS15, DGH^{+}15, FOT92, GBK^{+}96, KGBB09, KK11, LGC^{+}16, Nag04, PT09a, PT09b, SW16, SZC16, SRTCL16, WB16, XTCC15, YZ16, BH12, BBK^{+}11, BB09, BFH96, CLM05, Che93a, Che93b, CvHK97, DHH00, EAS^{+}97, GL04, GBH14, GGMS99, GHW00, HMC97, Hav00b, Lar93, Mat94, NDSG07, PVKE01, PMM94, PPS12a, SBM^{+}10, Slo12, Sny07, Szy07, VCT05, Wan02, War96, Hil97]. Programs [BCC^{+}92, MJB15, Den96, DF14, FP00a, GG95, GRR^{+}03, LDV07, LP99, OPE^{+}95, OdSSP13, RS94, SKS01, TKS02, Zhe10]. project [HW12, HR12a, HR12b]. properties [Din99]. propulsion [WCKD07]. Protein [CLA^{+}15, KG08, WSZK09]. protocol [MPP^{+}04]. prototyping [DSZ96]. provisioning [JDVM10, OPP11, Slo12]. PSEs [HP02]. psychological [VRM02]. Public [PAR94, HB94, RST02]. PUMA [WMR^{+}94]. purpose [Mat94]. Python [CLM05, DDMS14, Spo12]. PyTrilinos [Spo12].

QCD [IB09]. QoS

[HLPY16, MAG^{+}07, Slo12]. QR [CDO^{+}96, KD09]. Quality [ÁRCPS15, KFFZ05, PMCF94]. quality-cost [PMCF94]. quantification [GW11]. Quantitative [AE03]. quantum [RX04, KJG^{+}08]. Quasi [PPJ^{+}15].

Quasi-Optimal [PPJ^{+}15]. Questions [SHM97, SN02]. queueing [DDMS14].

R [KCO^{+}05]. R-GMA [KCO^{+}05]. Race [HJ15]. Racing [YZ16]. radar [PTS^{+}13]. Randomization [ZWZ16]. Rank [YTD15]. Ranking [NM16]. Ratio [VMRM16]. ray [BCHL05]. ready [TSCT11]. Real [BD16, PDROFRM16, Bry96, JR10, MAG^{+}07, SH94]. Real-Time [BD16, MAG^{+}07]. receptors [GGJ04]. Recommendation [CSM15, SLD^{+}16]. recommendations [AB96]. reconfigurable [KMB09]. Recursive [DEŁ01]. Reduction [KS15, SZAG15, ZHL16, KHSJ95]. Reference [Nag05b]. refinement [BAN02, CW93]. Regional [ZJZ^{+}16]. regions [MGK^{+}13]. Regular [Ano08c, NAA^{+}03]. regulations [ZGW08]. reinforcement [PVL^{+}04]. Relational [KCO^{+}05]. Reliability [TT16]. Reliability-Aware [TT16]. reliable [DRL13, KABW11]. Remanufacturing [ZGS16]. Remote [CTCG15, GL04, GBH14]. reordering [ZSS^{+}10]. Replicas [Wu16]. Report [PAR94, HB94]. Repositories [SLD^{+}16]. representation [Tým99]. Reproducibility [SPPH15]. Requirements [AKA16, SDS00]. rescheduling [KLN^{+}04, SZ04]. Research [HHD^{+}16, LZF16, ZYLZ15, AGG^{+}97, NKV^{+}02]. Reshaping [TFN11]. resolution [BAN02, VvAC^{+}09]. resolving [BAN02]. Resource [ACB^{+}02, GWZ16, HPZ16, MPA16, Özt04, CJS^{+}02, JDVM10, KLN^{+}04, OPP11]. Resources [HLPY16, KVW^{+}07]. Restart [CSC16]. Restoration [ZWCY16]. results [BHK^{+}13]. reusing [NAA^{+}03]. Reverse [HHD^{+}16]. reversing [HUN08]. Review [Ano07a, Ano07b, Ano07c, Ano11b, Bai08, Bor04, Bro11a, Bro11b, Bro12, Bro09, Bus09, Fil14, GB10, Hil97, Les10, Man08, Mic97, Nag04, Nag05a, Nag05b, Nag09a, Nag09b, Nag11a, Nag11b, Nag12, Ste97, Wes08]. Reviewers [Ano03, PT09b]. RISC [Hil97, Was95]. Risk [YG16, MCvM10]. Robert [Per08]. Robust [ZJZ^{+}16, CM02]. Robustness [ZWW16]. Rough [PDROFRM16]. Router [GWZ16]. routines [CDO^{+}96]. rover [MJLM07]. Rubin [Hil97]. Rules [ZYLZ15]. Run [EAS^{+}97, DN04, IRSD99, MSH99]. Run-time [EAS^{+}97, DN04, IRSD99, MSH99]. Runtime [Ano99b, HPS^{+}13, NPP^{+}00, Pro07].

S [Ano07a]. same [Mic00]. same-source [Mic00]. Sampling [YZ16]. SAMR [LTB02]. Satisfaction [SW16, AMM05]. satisfaction-based [AMM05]. Sawzall [PDGQ05]. SC13 [GM14]. SC2000 [Old00]. Scala [PSM^{+}15]. Scalability [SBJV11]. Scalable [BPK00, HPZ16, JHNP14, KCS15, NKV^{+}02, OHS00, CM02, DRR12, GBH14, IMB^{+}13, OS99, Ros00, Rou08a]. ScalaLab [PMTL12, PMTL14, PSM^{+}15]. ScaLAPACK [CDO^{+}96]. Scale [SZAG15, ZWW16, MJLM07, MSCS14, WGW08, Zim07]. Scale-Free [ZWW16]. SCALEA [TF04]. SCALEA-G [TF04]. Scaling [NAA^{+}03, MGK^{+}13, SCSJ09]. scanning [WSZK09]. SCASH [SHHI01]. scattering [HST^{+}93]. Scenarios [PDROFRM16]. schedules [NAA^{+}03]. Scheduling [CS94, HPZ16, KGMN15, LWW16, LKDB10, MFB^{+}15, MJB15, OCC^{+}16, TT16, WWLG16, ZWZ16, DRR12, FB99, GR93, KLN^{+}04, SJR14, Shu94, SCB02, WD07, ARFS05, BPC^{+}09]. schema [VRW^{+}03]. Scheme [ATZ^{+}15, DPMB16, SJR14]. schemes [SF03b, SF03a]. Science [BR11, HMSW92, KJG^{+}08, KBG^{+}02, DRT^{+}14, Ste97, TSCT11, DRT^{+}14]. Science-Driven [BR11]. Scientific [Ano07c, Che93b, DS97, DDMS15, GHH^{+}02, MFB^{+}15, Nag05b, PSM^{+}15, PT09a, PT09b, SPPH15, Sny07, SRTCL16, AJ94, ANEA08, BH12, BBŁ08, BGLR93, BÇCD12, CLM05, CHS^{+}99, DSS^{+}05, DDMS14, GG95, Gil09, Hav00a, HHML05, IJL^{+}01, JDVM10, KJG^{+}08, Mat94, NJ08, NDSG07, OSS94, PMTL12, PMTL14, Rou08b, SKU^{+}09, SLJ^{+}00, SBJV11, Szy07, VF95]. Scientist [Hil97]. script [BBBM12]. scripting [PMTL14]. SDN [ZWCY16]. SDN-Based [ZWCY16]. Search [LGC^{+}16, BAP13]. Searching [WS15, GA96]. Secondary [CTC^{+}15]. Section [Ano08c]. Sections [TC96]. Security [AlS16, CQF05]. Seed [CLA^{+}15]. Seed-Based [CLA^{+}15]. SEER [AMN^{+}12]. segmentation [RLC04]. segments [Abd02]. SEI [Ano07a]. Selected [VR13]. Selection [CLTX16, AMM05, WHRH07]. semantic [MPP^{+}04, BB08, GB15]. Semi [KG08, IJL^{+}01]. semi-automatic [IJL^{+}01]. Semi-supervised [KG08]. Semiseparable [YTD15]. Sensitivity [CO93]. Sensor [WS15, YZZ15]. Sensor-Based [YZZ15]. sequence [KG08, WSZK09]. sequences [KTP05, LSB12]. sequencing [HB12]. serial [CLM05]. Series [Ano07a, Nag05b, BCHL05]. server [BDLL94, GKL^{+}96]. servers [CDD^{+}05]. Service [CMTAC^{+}16, GB15, KFFZ05, MPP^{+}04, PHH95, WHRH07]. services [AMM05, TSCT11]. Set [PDROFRM16, YG16]. Sets [Wu16]. shallow [KBRS95]. shallow-water [KBRS95]. shared [DBVF01, HMCH07, NAA^{+}03, OB96, PBK01, SHHI01, TKS02, ZS99]. shared-memory [NAA^{+}03, ZS99]. sharing [BTS^{+}14, Ott93]. sided [LKDB10, GBH14]. Sieve [KK09]. signal [Den96]. signals [ADS95]. Signed [SSN15]. SIMD [TSP^{+}15]. SIMDization [IB09]. Similarities [CLA^{+}15]. simplified [DN04]. simulating [BDM^{+}04]. Simulation [CCBPGA15, O’K00, WB95, BBLR03, GIKP95, GGJ04, MJLM07, NKV^{+}02, OPP11, PPS12a, PPS^{+}12b, SE95, War14]. Simulations [SCD^{+}15, YDC^{+}16, KLS^{+}96, LLW^{+}14, LG03, MGK^{+}13, RX04, WCKD07]. Single [AlS16, VMRM16, XTCC15, AGIS94, DF14]. single- [DF14]. Single-Phase-to-Three-Phase [VMRM16]. single/multigrid [AGIS94]. Singularities [PPJ^{+}15]. Sink [WCZ^{+}16]. sites [IR02]. SKaMPI [RST02]. SKI [VC93]. Skillrank [ÁRCPS15]. Skills [ÁRCPS15]. sky [VvAC^{+}09]. SKYHI [Hem00]. SLA [ZHL16]. SLAM [LWYS16]. SLE [WOC99]. Sloan [Nag05a]. Slogger [BB08]. Slope [SC16]. SLPA [KCS15]. Smith [Mic97]. SMP [CCBPGA15]. Snir [Nag05b]. Social [ÁRCPS15, CSM15, DPMB16, GMCT15, KAS15, SSN15]. sockets [BŻvA^{+}01]. Soft [WS15]. Soft-Sensor [WS15]. Software [AKA16, Ano07a, Ano07c, AB96, CHC15, GBJ94, MOS16, OS99, O’K00, PT09a, PT09b, SLD^{+}16, Wes08, ANEA08, BFH96, DHH00, GB94, Gus93, Hav00a, KJG^{+}08, LBvBW12, MS00, MOT97, PMTL12, PPS12a, PPS^{+}12b, RX04, Rou08b, SHHI01, WSB11]. solution [HMCH07, SG96]. solutions [AGIS94]. Solver [CKKK15, DEvdV01, JR10, MSSG11, SCSJ09, SF03b, SF03a]. solvers [BHRT12, RR04]. Solving [BEK^{+}12, SW16, SCAP16, WB16, YTD15, YZ16, Gus93, Lan01, MRLF12, VRW^{+}03]. Some [McC96, Lar93]. Sorting [SCAP16]. Source [ZGS16, Mic00, SGM^{+}08]. Space [DPMB16, SV15, XML^{+}16, CBCM93, Nor07]. Spam [FGROL^{+}16]. Spanning [CBRRC^{+}16]. Spare [ZGS16]. Spare-Part [ZGS16]. Sparse [DEŁ01, BHRT12, CFR14, LP99, Lin04, SG96]. sparse-matrix [CFR14]. SPEC [AE03]. Special [BFGH14, BR11, GM14, HR12a, HR12b, JK10, PT09a, PT09b, SO11, VR13]. specialization [BMY13]. Specific [JBLB15, Ang93, SO93, ZGW08]. Specification [Hig93, Pro07, CMM^{+}02]. spectral [Ros00]. speech [Abd02]. SpeedShop [SGM^{+}08]. spike [PSU08, MSSG11]. spine [KLS^{+}96]. SPINET [KLS^{+}96]. Spitzer [JEM07]. spline [TOF^{+}14]. square [GG05]. Stability [SC16]. stack [Tým99]. stack-based [Tým99]. standard [FP00b, Met99a, Met99b, RN07]. star [BAN02]. State [BDH^{+}10, ADS95]. State-of-the-art [BDH^{+}10]. Static [Den96, GG95, ACIK97]. Statistical [KTP05, Din99, OHS00]. Stencil [CKKK15, SRO^{+}15]. Sterling [Mic97]. Steve [Nag05b, Bry96]. Steven [Nag05b]. Stewart [Ano07c]. Stochastic [HPZ16, HXX16, ZWZ16]. Storage [GM14, Slo12, SZXL16, DZKS13, KABW11]. Strassen [KHSJ95]. Strategies [OCC^{+}16, DBVF01, FC01, PCS99, Sch94]. Strategy [GB15, RRM^{+}15, WS15, Wu16, MGK^{+}13]. streams [Pla04]. Strength [Nag04]. strides [Bai95]. String [TLC15]. Strong [SCSJ09, ZCXQ16]. structural [LB02]. Structure [CLA^{+}15, SG96]. Structured [BFH96]. structures [HS03, KGV97]. studies [ZS99]. Study [ACCTA^{+}15, DS97, HBCM94, McC96, RRM^{+}15, AJJF14, BKP99, GW11, GB94, Hav00a, MS00, OB96, SBM^{+}10]. style [DHH00, Ano07c]. Subgraph [SJ15]. submesh [Mor94]. Suely [Ano07c]. Sundance [LBvBW12]. Super [VR13]. Supercomputer [GBK^{+}96, SH94]. supercomputers [Hil97]. Supercomputing [BCCP05]. Supernova [JMR^{+}11]. Supersonic [CCBPGA15]. supervised [KG08]. Supply [ZJZ^{+}16]. Support [CS94, CLTX16, OSLK12, XTCC15, BCC^{+}93, CHS^{+}99, EAS^{+}97, HCJ08, HMSM08, Pro07, RKR^{+}99, WSB11, WSP94]. supported [CMM^{+}02]. Supporting [ZA10, BBLR03]. survey [AB96, BCK07]. surveys [JEM07]. survival [AMN^{+}12]. Sweep [COB^{+}00, PDA^{+}08]. Switching [WS15]. SX [DTV00]. SX-4M [DTV00]. symbolic [SF03b, SF03a, VF95]. Symmetric [BDV03]. Symposium [SO11]. Synchronization [MM16]. synthetic [PTS^{+}13]. System [AlS16, FMA16, SZC16, Was95, YTD15, YZZ15, ZGS16, BB08, CO93, CJS^{+}02, CHS^{+}99, FP95, IMB^{+}13, KMR^{+}97, KZRR94, KABW11, KLN^{+}04, Lar93, MSSG11, MAG^{+}07, SCSJ09, SHHI01, TF04, WMR^{+}94, WSP94, XNQF04, Ros00]. System/6000 [Was95]. Systems [BD16, GBJ94, HXX16, Kuf15, MPA16, MJB15, PT09a, PT09b, TT16, YZZ15, ADLN08, Bai95, BHRT12, BTS^{+}14, BPK00, DSS^{+}05, DBVF01, DZKS13, DDMS14, LTB02, OS99, PVL^{+}04, SZ04, Sch94, STD^{+}14, SG96, WMR^{+}94, WGW08, Zhe10]. systolic [CIN^{+}96].

T [FP00a]. table [WSB11]. TACO [SJR14]. Task [CLYL16, LWW16, MJB15, WWLG16, OdSSP13, RR07, RKR^{+}99, RRV09]. task-level [RRV09]. tasks [BS01, DRR12, PKE^{+}10]. Teaching [DDMS15]. Teams [Ano07a]. Technique [HJ15, MOS16, VMRM16]. Techniques [AlS16, Abd02, BMY13, BPC^{+}09, KJG^{+}08, MS00, SKS01]. technologies [BB08, Mic97, vDKH01]. technology [BGH99, OSS94, VCT05]. Template [BMY13, COB^{+}00, PPS12a, CP03]. Template-based [PPS12a]. Templates [MWQ16]. temporal [PSU08]. ten [SN02]. Tensor [KHSJ95, Lan03]. TeraGyroid [BCCP05]. test [CIN^{+}96]. tests [Was95]. text [TFN11]. Their [GMCT15]. them [TSCT11]. theoretical [SH94]. Theory [PDROFRM16]. Things [WCZ^{+}16]. Thomas [Mic97]. thread [HPD09]. threads [BS01]. Three [VMRM16, KBRS95, Mar05]. three-dimensional [KBRS95]. tile [LKDB10]. Time [BD16, HXX16, SLBZ16, BCHL05, CGL08, DN04, EAS^{+}97, IRSD99, MAG^{+}07, MSH99, OdSSP13, ZGW08, ZA10, ARFS05]. time-course [ZGW08]. time-critical [ZA10]. time-series [BCHL05]. tolerance [ZA10]. Tolerant [HXX16]. tomography [DRL13, SCB02]. Tool [WSB11, AGG^{+}97, BKK^{+}11, CKS^{+}09, GRR^{+}03, SLJ^{+}00]. Toolkit [YZZ15, IJL^{+}01]. toolkits [BÇCD12]. Tools [KJG^{+}08, STD^{+}14, AB96, Met99a, Met99b, MOT97, RLL^{+}02]. top [SN02]. Topic [DPMB16]. Tpetra [BH12]. trace [CGL08]. tracking [CBCM93]. tradeoff [PMCF94]. tradeoffs [AJJF14]. Traffic [CGC16, BDM^{+}04]. Traffic-Aware [CGC16]. Transfers [THS^{+}15]. transformation [Lin04]. transformations [LKDB10, RR04]. Transient [Tým99]. transition [BBLR03, FP95, NDSG07]. translation [ARFS05, AJJF14, OPE^{+}95]. translator [OPE^{+}95]. transparent [KABW11, NPP^{+}00]. Transport [ZWCY16]. Tree [CBRRC^{+}16, War14]. Trees [PPJ^{+}15]. Triggered [HXX16]. Trilinos [HW12, HR12a, HR12b, OSLK12, Spo12]. Trios [OSLK12]. true [FO96]. Trust [DPMB16]. Trusted [HHD^{+}16]. TSP [Ano07a]. tuning [BFGH14, CMM^{+}02, DF14, FML^{+}14, HCJ08, MSCS14, STD^{+}14, SCB02]. turbulence [RX04]. tutorials [Met99a, Met99b]. Two [ACLS16, TC96, KMB09, Kok07, LKDB10, OB96]. two-dimensional [Kok07]. Two-Phase [ACLS16, TC96]. two-point [KMB09]. two-sided [LKDB10]. type [RMX05, Rou08b].

U [WOC99]. U-Net [WOC99]. U-Net/SLE [WOC99]. ubiquitous [ASCH^{+}07, ACB^{+}02]. UML [NCMF15]. Uncertain [YG16]. Uncertainties [ZJZ^{+}16]. Uncertainty [SZC16, GW11]. Understanding [Ano07b]. Unfavorable [Bai95]. UNICORE [Rom02]. unified [SZ09, TF04]. units [BB09]. Unix [Hil97]. Unstructured [SO15, SCSJ09]. UPC [Mar05, Zhe10]. Update [FO96]. Update-in-place [FO96]. Use [GW11, BH12, HS03, MPP^{+}04]. User [COB^{+}00, DPMB16, XST^{+}16, Hav00b, Hem00, WOC99]. user-customizable [WOC99]. User-Customized [XST^{+}16]. user-friendly [Hem00]. user-implemented [Hav00b]. User-level [COB^{+}00]. Users [ZG15]. Using [BD16, CTCG15, CS94, CSM15, CKKK15, CCBPGA15, DS97, HJ15, JBLB15, MWQ16, OPP11, RRM^{+}15, Sch94, SCD^{+}15, SJ15, SZC16, XTCC15, AMN^{+}12, ADS95, AGIS94, BAN02, CO93, DRL13, FSV14, HMCH07, IJL^{+}01, IMB^{+}13, JKR92, JJY^{+}03, JR10, JDVM10, KK11, LSB12, LLW^{+}14, LKDB10, MJLM07, PDA^{+}08, PKE^{+}10, SBM^{+}10, VvAC^{+}09, WHRH07, WCKD07, WC96]. utilized [KJG^{+}08].

Validation [BDV03]. Value [YZ16]. Variable [Sch03a, Tým99]. Variance [HY15, HB12]. Vector [BCHL05, CLTX16]. Vectorization [TSP^{+}15]. Vectorized [Kok07]. vectors [Hil97]. Verification [HHD^{+}16]. version [Ros00, Met99a, Met99b]. versus [OB96]. VFC [Ben99]. Video [ATZ^{+}15]. Vienna [Ben99, CMZ92]. View [Wu16, GKL^{+}96]. Violation [ZHL16]. Virtual [CGC16, GWZ16, KFFZ05, LWW16, XST^{+}16, ZHL16, ABB^{+}02, Bry96, WOC99, AGIS94]. Virtualization [NsP16]. Virtualized [HLPY16]. virus [LSB12]. Visualization [KGV97]. VLAM [ABB^{+}02, KVW^{+}07]. VLAM-G [ABB^{+}02, KVW^{+}07]. VLIW [YZZ15]. VM [JFPL16]. VoIP [CMTAC^{+}16]. Vol [Nag05b]. volatile [JIC^{+}14]. Volume [Ano99a, Ano00, Ano08a, Ano09, Ano11a, Ano12, Ano14, Ano97a]. volumes [Ano03]. VPP700 [DTV00]. VR [GWZ16]. VR-Cluster [GWZ16]. VRP [SLBZ16].

W [Ste97]. Walker [Nag05b]. water [KBRS95]. Watts [Ano07a]. wave [Mal01]. WCO [ZJZ^{+}16]. WCO-for-Biodiesel [ZJZ^{+}16]. Weather [O’K00, MGK^{+}13, Mal01, WCG95]. Web [BB08, FGROL^{+}16, GB15, IR02]. Weight [HY15]. Weighted [MWQ16]. Windows [SLBZ16, HLM^{+}11]. within [MRLF12]. work [OdSSP13, Sch94]. Workflow [ZWZ16, BCK07, KVW^{+}07, MAG^{+}07, Pro07, SVR^{+}07]. Workflows [MFB^{+}15, SPPH15, ASCH^{+}07, BBŁ08, BBBM12, DSS^{+}05, Gil09, JDVM10, KeKR^{+}11, SZ04, WD07, WHRH07]. Workload [DBVF01]. workspaces [KFFZ05]. workstation [Was95]. workstations [Hil97]. world [JR10]. Worldwide [VCT05]. Write [ZLL^{+}15]. Write-Combined [ZLL^{+}15]. Writing [Ano07c]. WSF2 [FGROL^{+}16].

XCAT [KBG^{+}02]. XCell [KGBB09]. Xeon [SCD^{+}15, SRO^{+}15, TSP^{+}15, ACCTA^{+}15, DGH^{+}15]. Xeon-Phi [ACCTA^{+}15]. xHPF [DS97]. XMatch [AMM05].

year [Met99a, Met99b]. Years [Szy07].

Zoltan [BÇCD12].

References

Appelbe:1996:STH

[AB96] Bill Appelbe and Donna Bergmark. Software tools for high-performance computing: survey and recommendations. Scientific Programming, 5(3):239–249, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Afsarmanesh:2002:VGG

[ABB^{+}02] H. Afsarmanesh, R. G. Belleman, A. S. Z. Belloum, et al. VLAM-G: a Grid-based virtual laboratory. Scientific Programming, 10(2):173–181, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                                      /iospress.metapress.com/
                                                      app/home/contribution.
                                                      asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                                      26referrer=parent%26backto=
                                                      issue%2C8%2C8%3Bjournal%
                                                      2C1%2C12%3Blinkingpublicationresults%
                                                      2C1%2C1.

Abdulla:2002:HBT

[Abd02] Waleed H. Abdulla. HMMbased techniques for speech segments extraction. Scientific Programming, 10(3): 221–239, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Aloisio:2002:GRB

[ACB^{+}02] Giovanni Aloisio, Massimo Cafaro, Euro Blasi, et al. The Grid Resource Broker, a ubiquitous grid computing framework. Scientific Programming, 10(2): 113–119, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C3%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Arroyo:2015:PSD

[ACCTA^{+}15] Maricela Arroyo, Carlos Couder-Castañeda, Alfredo Trujillo-Alcantara, Israel-Enrique Herrera-Diaz, and Nain Vera-Chavez. A performance study of a dual Xeon-Phi cluster for the forward modelling of gravitational fields. Scientific Programming, 2015(??):316012:1– 316012:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/316012/

.

Ancourt:1997:LAF

[ACIK97] Corinne Ancourt, Fabien Coelho, Francois Irigoin, and Ronan Keryell. Linear algebra framework for static High Performance Fortran code distribution. Scientific Programming, 6(1):3– 27, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ahn:2016:TPA

[ACLS16] Jun-Woo Ahn, Tai-Woo Chang, Sung-Hee Lee, and Yong Won Seo. Twophase algorithm for optimal camera placement. Scientific Programming, 2016 (??):4801784:1–4801784:16, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/4801784/.

Adalsteinsson:2008:CAC

[ADLN08] Helgi Adalsteinsson, Bert J. Debusschere, Kevin R. Long, and Habib N. Najm. Components for atomistic-to-continuum multiscale modeling of flow in micro- and nanofluidic systems. Scientific Programming, 16(4): 297–313, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anderson:1995:DMS

[ADS95] Charles W. Anderson, Saikumar V. Devulapalli, and Erik A. Stolz. Determining mental state from EEG signals using parallel implementations of neural networks. Scientific Programming, 4(3): 171–183, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Aslot:2003:QPA

[AE03] Vishal Aslot and Rudolf Eigenmann. Quantitative performance analysis of the SPEC OMPM2001 benchmarks. Scientific Programming, 11(2):105–124, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ayguade:1997:DRT

[AGG^{+}97] Eduard Ayguade, Jordi Garcia, Merce Girones, M. Luz Grande, and Jesus Labarta. DDT: a research tool for automatic data distribution in High Performance Fortran. Scientific Programming, 6(1): 73–94, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Averbuch:1994:PES

[AGIS94] A. Averbuch, E. Gabber, S. Itzikowitz, and B. Shoham. On the parallel elliptic single/multigrid solutions about aligned and nonaligned bodies using the Virtual Machine for Multiprocessors. Scientific Programming, 3(1):13– 32, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

AlHasan:2012:BKD

[AHCZ12] Mohammad Al Hasan, Jun Huan, Jake Chen, and Mohammed J. Zaki. Biological knowledge discovery and data mining. Scientific Programming, 20(1):1– 2, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Alaghband:1994:OFS

[AJ94] Gita Alaghband and Harry F. Jordan. Overview of the force scientific parallel language. Scientific Programming, 3(1): 33–47, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Arabas:2014:FTB

[AJJF14] Sylwester Arabas, Dorota Jarecka, Anna Jaruga, and Maciej Fijalkowski. Formula translation in Blitz++, NumPy and modern Fortran: A case study of the language choice tradeoffs. Scientific Programming, 22 (3):201–222, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Alsahli:2016:TAA

[AKA16] Abdulaziz Alsahli, Hameed Khan, and Sultan Alyahya. Toward an agile approach to managing the effect of requirements on software architecture during global software development. Scientific Programming, 2016 (??):8198039:1–8198039:16, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/8198039/.

Ai:2016:EMC

[ALL^{+}16] Wei Ai, Kenli Li, Shenglin Lan, Fan Zhang, Jing Mei, Keqin Li, and Rajkumar Buyya. On elasticity measurement in cloud computing. Scientific Programming, 2016 (??):7519507:1–7519507:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/7519507/.

AlShawi:2016:ADM

[AlS16] Amany AlShawi. Applying data mining techniques to improve information security in the cloud: A single cache system approach. Scientific Programming, 2016(??): 2385654:1–2385654:5, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/2385654/

.

Andreozzi:2005:XLS

[AMM05] Sergio Andreozzi, Danilo Montesi, and Rocco Moretti. XMatch: a language for satisfaction-based selection of Grid services. Scientific Programming, 13(4): 299–316, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Agrawal:2012:LCS

[AMN^{+}12] Ankit Agrawal, Sanchit Misra, Ramanathan Narayanan, Lalith Polepeddi, and Alok Choudhary. Lung cancer survival prediction using ensemble data mining on SEER data. Scientific Programming, 20(1):29– 42, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Allan:2008:MSS

[ANEA08] Benjamin A. Allan, Boyana Norris, Wael R. Elwasif, and Robert C. Armstrong. Managing scientific software complexity with Bocca and CCA. Scientific Programming, 16 (4):315–327, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Angus:1993:ADC

[Ang93] Ian G. Angus. Applications demand class-specific optimizations: the C++ compiler can do more. Scientific Programming, 2(4):123– 131, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1993:HPF

[Ano93] Anonymous. High Performance FORTRAN/journal of development. Scientific Programming, 2(1–2):1– 165, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1997:AIV

[Ano97a]  Anonymous.  Author in-

dex volume 6. Scientific Programming, 6(4):403– ??, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1997:CP

[Ano97b] Anonymous. Call for papers. Scientific Programming, 6(4):401, 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C7%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Anonymous:1999:AIV

[Ano99a] Anonymous. Author index volume 7 (1999). Scientific Programming, 7(3–4): 335–336, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=

                                        issue%2C12%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Anonymous:1999:HPJ

[Ano99b] Anonymous. High performance Java compilation and runtime issues. Scientific Programming, 7(2):85, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2000:AIV

[Ano00]  Anonymous.  Author in-

dex: Volume 7 (1999). Scientific Programming, 8(1): 335–336, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2001:I

[Ano01] Anonymous. Introduction. Scientific Programming, 9(2– 3):69–71, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2002:I

[Ano02] Anonymous. Introduction. Scientific Programming, 10 (3):183–184, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2003:RV

[Ano03] Anonymous. Reviewers for volumes 9–10. Scientific Programming, 11(1):77–78, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2004:I

[Ano04a] Anonymous. Introduction.

Scientific Programming, 12 (4):199, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2004:P

[Ano04b] Anonymous. Preface. Scientific Programming, 12(2): 63, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2005:F

[Ano05] Anonymous. Foreword. Scientific Programming, 13(2): 65–66, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRT

[Ano07a] Anonymous. Book review:

TSP^{SM} Coaching Development Teams (SEI Series in Software Engineering), by Watts S. Humphrey. Scientific Programming, 15(1): 67–69, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRU

[Ano07b] Anonymous. Book review:

Understanding and Implementing the Finite Element Method, by Mark Gockenbach. Scientific Programming, 15(2):117–119, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2007:BRW

[Ano07c] Anonymous. Book review: Writing Scientific Software: A Guide to Good Style, by Suely Oliveira and David Stewart. Scientific Programming, 15(3):189–190, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:AIV

[Ano08a]  Anonymous.  Author in-

dex volume 16 (2008). Scientific Programming, 16(4): 343–344, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:HPC

[Ano08b] Anonymous. High performance computing on Cell B.E. processors. Scientific Programming, 16(1): 99, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2008:RS

[Ano08c] Anonymous. Regular section. Scientific Programming, 16(1):79, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2009:AIV

[Ano09]  Anonymous.  Author in-

dex volume 17 (2009). Scientific Programming, 17(4): 347–348, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2011:AIV

[Ano11a]  Anonymous.  Author in-

dex volume 19 (2011). Scientific Programming, 19(4): 265–266, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2011:BR

[Ano11b] Anonymous. Book review.

Scientific Programming, 19 (2–3):179–184, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2012:AIV

[Ano12]  Anonymous.  Author in-

dex volume 20 (2012–2013). Scientific Programming, 20 (4):393–394, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:2014:AIV

[Ano14]  Anonymous.  Author in-

dex volume 22 (2014). Scientific Programming, 22(4): 331–332, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Alvarez-Rodriguez:2015:STH

[ÁRCPS15] Jose María Álvarez-Rodríguez, Ricardo Colomo-Palacios, and Vladimir Stantchev. Skillrank: Towards a hybrid method to assess quality and confidence of professional skills in social networks. Scientific Programming, 2015(??):451476:1– 451476:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/451476/

.

Agosta:2005:JJT

[ARFS05] Giovanni Agosta, Stefano Crespi Reghizzi, Gerlando Falauto, and Martino Sykora. JIST: Just-In-Time Scheduling translation for parallel processors. Scientific Programming, 13(3):239–253, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Addison:2003:OIA

[ARvW03] C. Addison, Y. Ren, and M. van Waveren. OpenMP issues arising in the development of parallel BLAS and LAPACK libraries. Scientific Programming, 11(2): 95–104, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Al-Shakarchi:2007:DWU

[ASCH^{+}07] Eddie Al-Shakarchi, Pasquale Cozza, Andrew Harrison, Carlo Mastroianni, Matthew Shields, Domenico Talia, and Ian Taylor. Distributing workflows over a ubiquitous P2P network. Scientific Programming, 15(4): 269–281, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Asif:2015:HSB

[ATZ^{+}15] Muhammad Asif, Imtiaz A.

Taj, S. M. Ziauddin, Maaz Bin Ahmad, and M. Tahir. A hybrid scheme based on pipelining and multitasking in mobile application processors for advanced video coding. Scientific Programming, 2015(??):197843:1– 197843:16, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/197843/

.

Bailey:1995:USC

[Bai95]  David H. Bailey.  Un-

favorable strides in cache memory systems. Scientific Programming, 4(2):53– 58, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). Also available as RNR technical report RNR-92-015.

Bailey:2008:BR

[Bai08]  David H. Bailey.  Book

review. Scientific Programming, 16(1):97–98, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bryan:2002:AER

[BAN02] Greg L. Bryan, Tom Abel, and Michael L. Norman. Achieving extreme resolution in numerical cosmology using adaptive mesh refinement: resolving primordial star formation. Scientific Programming, 10(4):291–302, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Beamer:2013:DOB

[BAP13] Scott Beamer, Krste Asanović, and David Patterson. Directionoptimizing breadth-first search. Scientific Programming, 21 (3–4):137–148, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baker:2008:SPA

[BB08] Mark Baker and Richard Boakes. Slogger: a profiling and analysis system based on Semantic Web technologies. Scientific Programming, 16(2–3):183–204, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bodin:2009:HMP

[BB09] Francois Bodin and Stephane Bihan. Heterogeneous multicore parallel programming for graphics processing units. Scientific Programming, 17 (4):325–336, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baranowski:2012:CWS

[BBBM12] Mikolaj Baranowski, Adam Belloum, Marian Bubak, and Maciej Malawski. Constructing workflows from script applications. Scientific Programming, 20(4): 359–377, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Budimlic:2010:CC

[BBC^{+}10] Zoran Budimlić, Michael Burke, Vincent Cavé, Kathleen Knobe, Geoff Lowney, Ryan Newton, Jens Palsberg, David Peixotto, Vivek Sarkar, Frank Schlimbach, and Sagnak Tasirlar. Concurrent collections. Scientific Programming, 18(3–4): 203–217, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Badin:2011:IAM

[BBDN11] Matthew Badin, Lubomir Bic, Michael Dillencourt, and Alexandru Nicolau. Improving accuracy for matrix multiplications on GPUs. Scientific Programming, 19(1): 3–11, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bali:2004:GEL

[BBF^{+}04] Bartosz Bali, Marian Bubak, Włodzimierz Funika, Roland Wismüller, Marcin Radecki, Tomasz Szepieniec, Tomasz Arod, and Marcin Kurdziel. Grid environment for online application monitoring and performance analysis. Scientific Programming, 12 (4):239–251, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bodin:1993:DPB

[BBG^{+}93] François Bodin, Peter Beckman, Dennis Gannon, Srinivas Narayana, and Shelby X. Yang. Distributed pC++: Basic ideas for an object parallel language. Scientific Programming, 2(3):7–22, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

ftp://cica.cica.indiana.
edu/pub/sage/pC++SciPro.
ps.gz.

Blazewicz:2011:CPA

[BBK^{+}11] Marek Blazewicz, Steven R.

Brandt, Michal Kierzynka, Krzysztof Kurowski, Bogdan Ludwiczak, Jian Tao, and Jan Weglarz. CaKernel — a parallel application programming framework for heterogeneous computing architectures. Scientific Programming, 19(4): 185–197, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Balis:2008:MGS

[BBŁ08] Bartosz Balis, Marian Bubak, and Bartłomiej Łabno. Monitoring of Grid scientific workflows. Scientific Programming, 16(2–3):205–216, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bischof:2003:IES

[BBLR03] Christian H. Bischof, H. Martin Bücker, Bruno Lang, and Arno Rasch. An interactive environment for supporting the transition from simulation to optimization. Scientific Programming, 11 (4):263–272, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blount:1999:EJN

[BC99] Brian Blount and Siddhartha Chatterjee. An evaluation of Java for numerical computing. Scientific Programming, 7(2):97–110, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C3%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Bischof:1992:AGD

[BCC^{+}92] Christian H. Bischof, Alan Carle, George F. Corliss, Andreas Griewank, and Paul Hovland. ADIFOR: Generating derivative code from Fortran programs. Scientific Programming, 1(1): 11–29, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bhatt:1993:OOS

[BCC^{+}93] Sandeep Bhatt, Marina Chen, James Cowie, Cheng-Yee Lin, and Pangfeng Liu. Object-oriented support for adaptive methods on parallel machines. Scientific Programming, 2(4):179–192, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bircsak:2000:EON

[BCC^{+}00] John Bircsak, Peter Craig, RaeLyn Crowell, et al. Extending OpenMP for NUMA machines. Scientific Programming, 8(3):163–181, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Boman:2012:ZIP

[BÇCD12] Erik G. Boman, Ümit V.

Çatalyürek, Cédric Chevalier, and Karen D. Devine. The Zoltan and Isorropia parallel toolkits for combinatorial scientific computing: Partitioning, ordering and coloring. Scientific Programming, 20(2): 129–150, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blake:2005:TES

[BCCP05] R. J. Blake, P. V. Coveney, P. Clarke, and S. M. Pickles. The TeraGyroid experiment — Supercomputing 2003. Scientific Programming, 13(1):1–17, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Banicescu:2005:VNT

[BCHL05] Ioana Banicescu, Ricolindo L. Cariño, Jane L. Harvill, and John Patrick Lestrade. Vector nonlinear time-series analysis of gamma-ray burst datasets on heterogeneous clusters. Scientific Programming, 13(2): 67–77, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bahsi:2007:CWM

[BCK07] Emir M. Bahsi, Emrah Ceyhan, and Tevfik Kosar. Conditional workflow management: a survey and analysis. Scientific Programming, 15(4):283–297, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bala:2001:APC

[BCS01] Piotr Bała, Terry Clark, and L. Ridgway Scott. Application of Pfortran and Co-Array Fortran in the parallelization of the GROMOS96 molecular dynamics module. Scientific Programming, 9(1):61–68, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C6%2C6%3Bjournal%
                                        2C2%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Beamonte:2016:DCP

[BD16] Raphaël Beamonte and Michel R. Dagenais. Detection of common problems in real-time and multicore systems using model-based constraints. Scientific Programming, 2016 (??):9792462:1–9792462:18,

????  2016.  CODEN

SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/9792462/.

Beguelin:1994:HHN

[BDG^{+}94] Adam Beguelin, Jack J.

Dongarra, George Al Geist, Robert Manchek, and Keith Moore. HeNCE: a heterogeneous network computing environment. Scientific Programming, 3(1): 49–60, Spring 1994. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://www.netlib.org/
utk/people/JackDongarra/
PAPERS/HeNCE-A-Heterogeneous-
Network-Computing-Environment.
pdf.

Brodtkorb:2010:SAH

[BDH^{+}10] Andre R. Brodtkorb, Christopher Dyken, Trond R. Hagen, Jon M. Hjelmervik, and Olaf O. Storaasli. Stateof-the-art in heterogeneous computing. Scientific Programming, 18(1):1–33, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Berry:1994:PPD

[BDLL94] Michael W. Berry, Jack J.

Dongarra, Brian H. LaRose, and Todd A. Letsche. PDS: a performance database server. Scientific Programming, 3 (2):147–156, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Barrett:2004:MTG

[BDM^{+}04] Chris Barrett, Martin Drozda, Madhav V. Marathe, S. S. Ravi, and James P. Smith. A mobility and traffic generation framework for modeling and simulating ad hoc communication networks. Scientific Programming, 12(1): 1–23, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Briguglio:2003:PPM

[BDV03] Sergio Briguglio, Beniamino Di Martino, and Gregorio Vlad. A performanceprediction model for PIC applications on clusters of symmetric multiprocessors: Validation with hierarchical HPF + OpenMP implementation. Scientific Programming, 11(2):159–176, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bochev:2012:SPI

[BEK^{+}12] P. Bochev, H. C. Edwards, R. C. Kirby, K. Peterson, and D. Ridzal. Solving PDEs with Intrepid. Scientific Programming, 20(2): 151–180, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Benkner:1999:VVF

[Ben99] Siegfried Benkner. VFC:

The Vienna Fortran Compiler. Scientific Programming, 7(1):67–81, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C5%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Benkner:2014:SIA

[BFGH14] Siegfried Benkner, Franz Franchetti, Hans Michael Gerndt, and Jeffrey K. Hollingsworth. Special issue on automatic application tuning for HPC architectures. Scientific Programming, 22(4):259–260, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Burkhart:1996:SPP

[BFH96] Helmar Burkhart, Robert Frank, and Guido Hachler. Structured parallel programming: how informatics can help overcome the software dilemma. Scientific Programming, 5(1):33–45, Spring 1996. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic).

Bik:1999:IIM

[BGH99] Aart J. C. Bik, Milind Girkar, and Mohammad R. Haghighat. Incorporating Intel(R) MMX^{T M} technology into a Java^{T M} JIT compiler. Scientific Programming, 7(1):167–184, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C7%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Ballance:1993:FBE

[BGLR93] Robert A. Ballance, Anthony J. Giancola, George F. Luger, and Timothy J. Ross. Framework-based environment for object-oriented scientific codes. Scientific Programming, 2(4):111–121, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bohm:1996:FIJ

[BH96] A. P. W. Bohm and R. E. Hiromoto. Functional implementations of the Jacobi Eigensolver. Scientific Programming, 5(2):111– 120, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Boys:2002:DOM

[BH02] R. J. Boys and D. A. Henderson. On determining the order of Markov dependence of an observed process governed by a hidden Markov model. Scientific Programming, 10(3):241–251, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Baker:2012:TUG

[BH12] C. G. Baker and M. A. Heroux. Tpetra, and the use of generic programming in scientific computing. Scientific Programming, 20(2): 115–128, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blazewicz:2013:PMR

[BHK^{+}13] Marek Blazewicz, Ian Hinder, David M. Koppelman, Steven R. Brandt, Milosz Ciznicki, Michal Kierzynka, Frank Löffler, Erik Schnetter, and Jian Tao. From physics model to results: an optimizing framework for crossarchitecture code generation. Scientific Programming, 21(1–2):1–16, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Barekas:2003:MAO

[BHP^{+}03] Vasileios K. Barekas, Panagiotis E. Hadjidoukas, Eleftherios D. Polychronopoulos, et al. A multiprogramming aware OpenMP implementation. Scientific Programming, 11(2):133–141, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bavier:2012:ABD

[BHRT12] Eric Bavier, Mark Hoemmen, Sivasankaran Rajamanickam, and Heidi Thornquist. Amesos2 and Belos: Direct and iterative solvers for large sparse linear systems. Scientific Programming, 20 (3):241–255, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bjorstad:2000:CFN

[Bjø00] Petter E. Bjørstad. Coordinate free numerics. Scientific Programming, 8(4):209, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
                                        26referrer=parent%26backto=
                                        issue%2C1%2C5%3Bjournal%
                                        2C3%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Bak:2011:GTD

[BKK^{+}11] Slawomir Bak, Marcin Krystek, Krzysztof Kurowski, Ariel Oleksiak, Wojciech Piatek, and Jan Waglarz. GSSIM — a tool for distributed computing experiments. Scientific Programming, 19(4):231–251, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bubak:2001:CJN

[BKLS01] Marian Bubak, Dawid Kurzyniec, Piotr Luszczek, and V. Sunderam. Creating Java to Native Code Interfaces with Janet. Scientific Programming, 9(1):39–50, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C4%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Budimlic:1999:CBO

[BKP99] Zoran Budimlić, Ken Kennedy, and Jeff Piper. The cost of being object-oriented: a preliminary study. Scientific Programming, 7(2): 87–95, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C2%2C8%3Bjournal%

                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Bozkus:1997:POH

[BMN^{+}97] Zeki Bozkus, Larry Meadows, Steven Nakamoto, Vincent Schuster, and Mark Young. PGHPF — an optimizing High Performance Fortran compiler for distributed memory machines. Scientific Programming, 6(1): 29–40, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bachelet:2013:TMT

[BMY13] Bruno Bachelet, Antoine Mahul, and Loïc Yon. Template metaprogramming techniques for concept-based specialization. Scientific Programming, 21(1–2):43– 61, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Borchers:2004:BRL

[Bor04] Brian Borchers. Book review: The Lanczos Method: Evolution and Application, by Louis Komzsik. Scientific Programming, 12(3): 197–198, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bellens:2009:CST

[BPC^{+}09] Pieter Bellens, Josep M. Perez, Felipe Cabarcas, Alex Ramirez, Rosa M. Badia, and Jesus Labarta. CellSs: Scheduling techniques to better exploit memory hierarchy. Scientific Programming, 17(1–2):77–95, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brunner:2000:SMD

[BPK00] Robert K. Brunner, James C.

Phillips, and Laxmikant V. Kalé. Scalable molecular dynamics for large biomolecular systems. Scientific Programming, 8(3):195–207, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brandic:2011:SIS

[BR11] Ivona Brandic and Ioan Raicu. Special issue on science-driven cloud computing. Scientific Programming, 19(2–3):71–73, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brownston:2009:BR

[Bro09] Lee S. Brownston. Book review. Scientific Programming, 17(4):339–341, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2011:BRa

[Bro11a] Adrian Brown. Book review.

Scientific Programming, 19 (1):63–65, ???? 2011. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2011:BRb

[Bro11b] Adrian Brown. Book review. Scientific Programming, 19 (4):259–264, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Brown:2012:BR

[Bro12] Adrian Brown. Book review. Scientific Programming, 20 (3):355–358, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bryson:1996:PCV

[Bry96] Steve Bryson. Pictorial communication in virtual and real environments, edited by Steve Ellis, Mary Kaiser, and Art Grunwald. Scientific Programming, 5(1):89– ??, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blikberg:2001:NPA

[BS01] Ragnhild Blikberg and Tor Sørevik. Nested parallelism: Allocation of threads to tasks and OpenMP implementation. Scientific Programming, 9(2–3):185– 194, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C11%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Bollman:1996:FDI

[BSF96] Dorothy Bollman, Jaime Seguel, and John Feo. Fast digit-index permutations. Scientific Programming, 5 (2):137–146, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Breslow:2014:EFP

[BTS^{+}14] Alex D. Breslow, Ananta Tiwari, Martin Schulz, Laura Carrington, Lingjia Tang, and Jason Mars. Enabling fair pricing on high performance computer systems with node sharing. Scientific Programming, 22(2):59– 74, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bushnell:2009:BR

[Bus09] David H. Bushnell. Book review. Scientific Programming, 17(3):275–277, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Bubak:2001:PLM

[BŻvA^{+}01] Marian Bubak, Dariusz Żbik, Dick van Albada, et al. Portable library of migratable sockets. Scientific Programming, 9(4): 211–222, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C2%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Chauveau:1999:MEH

[CB99] Stéphane Chauveau and François Bodin. Menhir: An environment for high performance Matlab. Scientific Programming, 7(3– 4):303–312, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C9%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Chang:1993:PIP

[CBCM93] L. Chang, G. Bourianoff, B. Cole, and S. Machida. Parallel implementation of particle tracking with space charge effects on an Intel iPSC/

860. Scientific Programming, 2(3):37–47, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Contreras-Bolton:2016:APA

[CBRRC^{+}16] Carlos Contreras-Bolton, Carlos Rey, Sergio Ramos-Cossio, Claudio Rodríguez, Felipe Gatica, and Victor Parada. Automatically produced algorithms for the generalized minimum spanning tree problem. Scientific Programming, 2016 (??):1682925:1–1682925:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/1682925/.

Couder-Castaneda:2015:PCM

[CCBPGA15] C. Couder-Castañeda, H. Barrios-Piña, I. Gitler, and M. Arroyo. Performance of a code migration for the simulation of supersonic ejector flow to SMP, MIC, and GPU using OpenMP, OpenMP+LEO, and OpenACC directives. Scientific Programming, 2015(??): 739107:1–739107:20, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/739107/

. Chandrasekaran:2015:PML

[CCTY15] Sunita Chandrasekaran, Barbara Chapman, Xinmin Tian, and Yonghong Yan. Programming models, languages, and compilers for manycore and heterogeneous architectures. Scientific Programming, 2015 (??):376317:1, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/376317/

.

Caron:2005:MDP

[CDD^{+}05] Eddy Caron, Bruno DelFabbro, Frédéric Desprez, Emmanuel Jeannot, and Jean-Marc Nicod. Managing data persistence in network enabled servers. Scientific Programming, 13(4):333–354, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Choi:1996:DIS

[CDO^{+}96] Jaeyoung Choi, Jack J. Dongarra, L. Susan Ostrouchov, Antoine P. Petitet, David W. Walker, and R. Clint Whaley. Design and implementation of the ScaLAPACK LU, QR, and Cholesky factorization routines. Scientific Programming, 5(3): 173–184, Fall 1996. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:

//www.netlib.org/netlib/
lapack/lawns/lawn80.ps;

 http://www.netlib.org/
netlib/lapack/lawnspdf/
lawn80.pdf;   http://www.
netlib.org/utk/papers/
factor/ftcover.html.

Cardellini:2014:DPS

[CFR14] Valeria Cardellini, Salvatore Filippone, and Damian W. I. Rouson. Design patterns for sparse-matrix computations on hybrid CPU/GPU platforms. Scientific Programming, 22(1):1–19, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chen:2016:OVM

[CGC16] Tao Chen, Xiaofeng Gao, and Guihai Chen. Optimized virtual machine placement with traffic-aware balancing in data center networks. Scientific Programming, 2016 (??):3101658:1–3101658:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/3101658/.

Chmiel:2005:EJA

[CGK^{+}05] Krzysztof Chmiel, Maciej Gawinecki, Pawel Kaczmarek, Michal Szymczak, and Marcin Paprzycki. Efficiency of JADE agent platform. Scientific Programming, 13(2):159–172, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chan:2008:EFN

[CGL08] Anthony Chan, William Gropp, and Ewing Lusk. An efficient format for nearly constant-time access to arbitrary time intervals in large trace files. Scientific Programming, 16(2–3):155– 165, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Carver:2015:SEC

[CHC15] Jeffrey C. Carver, Neil Chue Hong, and Selim Ciraci. Software engineering for CSE. Scientific Programming, 2015 (??):591562:1–591562:2, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2015/591562/

.

Cheng:1993:HCN

[Che93a]  Harry H. Cheng.  Han-

dling of complex numbers in the C^{H} programming language. Scientific Programming, 2(3):77–106, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cheng:1993:SCC

[Che93b] Harry H. Cheng. Scientific computing in the C^{H} programming language. Scientific Programming, 2(3): 49–75, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chapman:1997:OCL

[CHM^{+}97] Barbara Chapman, Matthew Haines, Piyush Mehrotra, Hans Zima, and John Van Rosendale. Opus: a coordination language for multidisciplinary applications. Scientific Programming, 6 (4):345–362, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C2%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Carter:1999:IMS

[CHS^{+}99] John B. Carter, Wilson C.

Hsieh, Leigh B. Stoller, et al. Impulse: Memory system support for scientific applications. Scientific Programming, 7(3–4):195–209, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C2%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Cornu:1996:DIT

[CIN^{+}96] Thierry Cornu, Paolo Ienne, Dagmar Niebur, Patrick Thiran, and Marc A. Viredaz. Design, implementation, and test of a multi-model systolic neural-network accelerator. Scientific Programming, 5(1): 47–61, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cao:2002:AAB

[CJS^{+}02] Junwei Cao, Stephen A. Jarvis, Subhash Saini, et al. ARMS: An agent-based resource management system for grid computing. Scientific Programming, 10(2): 135–148, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C8%3Bjournal%
                                        2C1%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Chapman:2008:F

[CK08] Barbara Chapman and Dieter Kranzlmüller. Foreword. Scientific Programming, 16 (2–3):101–103, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ciznicki:2015:MLB

[CKKK15] Milosz Ciznicki, Michal Kulczewski, Piotr Kopta, and Krzysztof Kurowski. Methods to load balance a GCR pressure solver using a stencil framework on multi- and many-core architectures. Scientific Programming, 2015(??):648752:1– 648752:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/648752/

.

Choi:2009:EMT

[CKS^{+}09] Wonil Choi, Hyunhee Kim, Wook Song, Jiseok Song, and Jihong Kim. ePRO-MP: a tool for profiling and optimizing energy and performance of mobile multiprocessor applications. Scientific Programming, 17(4): 285–294, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chow:2004:CCM

[CL04] P. Chow and C.-H. Lai.

Collaborating components in mesh-based electronic packaging. Scientific Programming, 12(2):65–70, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chapuis:2015:PSB

[CLA^{+}15] Guillaume Chapuis, Mathilde Le Boudic-Jamin, Rumen Andonov, Hristo Djidjev, and Dominique Lavenier. Parallel seed-based approach to multiple protein structure similarities detection. Scientific Programming, 2015(??): 279715:1–279715:12, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/279715/

.

Cai:2005:PPP

[CLM05] Xing Cai, Hans Petter Langtangen, and Halvard Moe. On the performance of the Python programming language for serial and parallel scientific computations. Scientific Programming, 13(1): 31–56, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chen:2016:PGA

[CLTX16] Zhi Chen, Tao Lin, Ningjiu Tang, and Xin Xia. A parallel genetic algorithm based feature selection and parameter optimization for support vector machine. Scientific Programming, 2016 (??):2739621:1–2739621:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/2739621/.

Choi:2016:TCB

[CLYL16] HeeSeok Choi, JongBeom Lim, Heonchang Yu, and EunYoung Lee. Task classification based energy-aware consolidation in clouds. Scientific Programming, 2016 (??):6208358:1–6208358:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/6208358/.

Cappello:2002:CSR

[CM02] Peter Cappello and Dimitrios Mourloukos. CX: a scalable, robust network for parallel computing. Scientific Programming, 10(2): 159–171, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C7%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Cesar:2002:DPT

[CMM^{+}02] Eduardo César, Anna Morajko, Tomàs Margalef, et al. Dynamic performance tuning supported by program specification. Scientific Programming, 10(1):35–44, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=

                                        issue%2C4%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Cortes-Mendoza:2016:BVS

[CMTAC^{+}16] Jorge M. Cortés-Mendoza, Andrei Tchernykh, Fermin A. Armenta-Cano, Pascal Bouvry, Alexander Yu. Drozdov, and Loic Didelot. Biobjective VoIP service management in cloud infrastructure. Scientific Programming, 2016(??):5706790:1– 5706790:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2016/5706790/

.

Chapman:1992:PVF

[CMZ92] Barbara M. Chapman, Piyush Mehrotra, and Hans P. Zima. Programming in Vienna Fortran. Scientific Programming, 1(1):31–50, Fall 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Calhoun:1993:SAD

[CO93] Donna Calhoun and Roy Overstreet. Sensitivity analysis of a dynamical system using C++. Scientific Programming, 2(4):157– 169, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Casanova:2000:APS

[COB^{+}00] Henri Casanova, Graziano Obertelli, Francine Berman, et al. The AppLeS parameter sweep template: Userlevel middleware for the Grid. Scientific Programming, 8(3):111–126, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Crooks:1995:LCD

[CP95] P. Crooks and R. H. Perrott.

Language constructs for data partitioning and distribution. Scientific Programming, 4(2):59–85, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Cole:2003:DCC

[CP03] Martin J. Cole and Steven G.

Parker. Dynamic compilation of C++ template code. Scientific Programming, 11 (4):321–327, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Clayton:2005:ASC

[CQF05] Brian C. Clayton, Thomas B.

Quillinan, and Simon N. Foley. Automating security configuration for the Grid. Scientific Programming, 13 (2):113–125, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Craus:2005:PFC

[CR05] Mitică Craus and Laurent.iu Rudeanu. Parallel framework for cooperative processes. Scientific Programming, 13 (3):205–217, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chapin:1994:SIS

[CS94] Steve J. Chapin and Eugene H. Spafford. Support for implementing scheduling algorithms using MESSIAHS. Scientific Programming, 3 (4):325–340, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Chen:2016:OCR

[CSC16] Zhengyu Chen, Jianhua Sun, and Hao Chen. Optimizing checkpoint restart with data deduplication. Scientific Programming, 2016 (??):9315493:1–9315493:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/9315493/.

Ciesielczyk:2015:ELR

[CSM15] Michał Ciesielczyk, Andrzej Szwabe, and Mikołaj Morzy. On efficient link recommendation in social networks using actor-fact matrices. Scientific Programming, 2015 (??):450215:1–450215:9, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/450215/

.

Chang:2015:ISI

[CTC^{+}15] Bao Rong Chang, Hsiu-Fen Tsai, Chia-Yen Chen, Chien-Feng Huang, and Hung-Ta Hsu. Implementation of secondary index on cloud computing NoSQL database in big data environment. Scientific Programming, 2015(??): 560714:1–560714:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/560714/

.

Chang:2015:EAH

[CTCG15] Bao Rong Chang, Hsiu-Fen Tsai, Chia-Yen Chen, and Cin-Long Guo. Empirical analysis of high efficient remote cloud data center backup using HBase and Cassandra. Scientific Programming, 2015(??):294614:1– 294614:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/294614/

.

Clark:1997:EDP

[CvHK97] Terry W. Clark, Reinhard v. Hanxleden, and Ken Kennedy. Experiences in data-parallel programming. Scientific Programming, 6 (1):153–158, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Crutchfield:1993:OOI

[CW93] William Y. Crutchfield and Michael L. Welcome. Objectoriented implementation of adaptive mesh refinement algorithms. Scientific Programming, 2(4):145–156, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

DiMartino:2001:WDS

[DBVF01] Beniamino Di Martino, Sergio Briguglio, Gregorio Vlad, and Giuliana Fogaccia. Workload decomposition strategies for shared memory parallel systems with OpenMP. Scientific Programming, 9(2– 3):109–122, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        http://iospress.metapress.
                                        com/app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Dolgopolovas:2014:PSC

[DDMS14] Vladimiras Dolgopolovas, Valentina Dagiene, Saulius Minkevicius, and Leonidas Sakalauskas. Python for scientific computing education: Modeling of queueing systems. Scientific Programming, 22(1):37–51, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dolgopolovas:2015:TSC

[DDMS15] Vladimiras Dolgopolovas, Valentina Dagienė, Saulius Minkevičius, and Leonidas Sakalauskas. Teaching scientific computing: A modelcentered approach to pipeline and parallel programming with C. Scientific Programming, 2015(??):820803:1– 820803:18, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/820803/

.

Doolin:1999:JCL

[DDS99] David M. Doolin, Jack Dongarra, and Keith Seymour. JLAPACK — compiling LA-PACK FORTRAN to Java. Scientific Programming, 7 (2):111–138, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C4%2C8%3Bjournal%

                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1;   http://www.
                                        netlib.org/utk/people/
                                        JackDongarra/PAPERS/f2jrep~

1.pdf. The software is available on the World-Wide Web

                                        at http://www.cs.utk.edu/
                                        f2j/.

Dongarra:2001:RAS

[DEŁ01] Jack Dongarra, Victor Eijkhout, and Piotr Łuszczek. Recursive approach in sparse matrix LU factorization. Scientific Programming, 9(1): 51–60, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C6%3Bjournal%
                                        2C2%2C9%3Blinkingpublicationresults%
                                        2C1%2C1;   http://www.
                                        netlib.org/netlib/utk/
                                        people/JackDongarra/PAPERS/
                                        recur-sparse-sciprog.pdf;
                                        http://www.netlib.org/
                                        utk/people/JackDongarra/
                                        PAPERS/rlu03.pdf.

Dennis:1996:SMF

[Den96] Jack B. Dennis. Static mapping of functional programs: an example in signal processing. Scientific Programming, 5(2):121–135, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dongarra:2001:ISB

[DEvdV01] Jack Dongarra, Victor Eijkhout, and Henk van der Vorst. An iterative solver benchmark. Scientific Programming, 9(4):223–231, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C3%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1;   http://www.
netlib.org/utk/people/
JackDongarra/PAPERS/sparse-
bench.pdf.

Durillo:2014:SMO

[DF14] Juan Durillo and Thomas Fahringer. From single- to multi-objective auto-tuning of programs: Advantages and implications. Scientific Programming, 22(4): 285–297, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Deboni:1996:HIF

[DFP96] Thomas Deboni, John Feo, and Doug Peters. Hybrid imperative and functional molecular mechanics application. Scientific Programming, 5(2):97–109, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dongarra:2015:HPI

[DGH^{+}15] Jack Dongarra, Mark Gates, Azzam Haidar, Yulu Jia, Khairul Kabir, Piotr Luszczek, and Stanimire Tomov. HPC programming on Intel Many-Integrated-Core hardware with MAGMA port to Xeon Phi. Scientific Programming, 2015(??):502593:1– 502593:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/502593/

.

Dinesh:2000:APS

[DHH00] T. B. Dinesh, Magne Haveraaen, and Jan Heering. An algebraic programming style for numerical software and its optimization. Scientific Programming, 8 (4):247–259, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C4%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Dinda:1999:SPH

[Din99] Peter A. Dinda. The statistical properties of host load. Scientific Programming, 7(3–4):211–229, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C3%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Deelman:2002:GC

[DK02] Ewa Deelman and Carl Kesselman. Grid computing. Scientific Programming, 10(2):101–102, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C1%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Decyk:2004:SMI

[DN04] Viktor K. Decyk and Charles D.

Norton. A simplified method for implementing run-time polymorphism in Fortran95. Scientific Programming, 12 (1):45–55, ???? 2004. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Decyk:1997:HEC

[DNS97] Viktor K. Decyk, Charles D.

Norton, and Bolesław K. Szymanski. How to express C++ concepts in Fortran90. Scientific Programming, 6 (4):363–390, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C3%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Dong:2016:TSO

[DPMB16] Wang Dong, Wang Peng, Li Meizi, and Zhang Bo. A topic space oriented user group discovering scheme in social network: A trust chain based interest measuring perspective. Scientific Programming, 2016 (??):4091323:1–4091323:25, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/4091323/.

Dichev:2013:ERN

[DRL13] Kiril Dichev, Fergal Reid, and Alexey Lastovetsky. Efficient and reliable network tomography in heterogeneous networks using Bit-Torrent broadcasts and clustering algorithms. Scientific Programming, 21(3–4): 79–92, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dummler:2012:CSM

[DRR12] Jörg Dümmler, Thomas Rauber, and Gudula Rünger. Combined scheduling and mapping for scalable computing with parallel tasks. Scientific Programming, 20(1): 45–67, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Dart:2014:SDN

[DRT^{+}14] Eli Dart, Lauren Rotman, Brian Tierney, Mary Hester, and Jason Zurawski. The Science DMZ: A network design pattern for dataintensive science. Scientific Programming, 22(2): 173–185, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

DeSturler:1997:SPH

[DS97] Eric De Sturler and Volker Strumpen. Scientific programming with High Performance Fortran: a case study using the xHPF compiler. Scientific Programming, 6 (1):127–152, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Deelman:2005:PFM

[DSS^{+}05] Ewa Deelman, Gurmeet Singh, Mei-Hui Su, James Blythe, Yolanda Gil, Carl Kesselman, Gaurang Mehta, Karan Vahi, G. Bruce Berriman, John Good, Anastasia Laity, Joseph C. Jacob, and Daniel S. Katz. Pegasus: a framework for mapping complex scientific workflows onto distributed systems. Scientific Programming, 13(3): 219–237, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Derby:1996:NLD

[DSZ96] Thomas Derby, Robert Schnabel, and Benjamin Zorn. New language design for prototyping numerical computation. Scientific Programming, 5(4):279–300, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Desgagne:2000:PME

[DTV00] Michel Desgagné, Stephen Thomas, and Michel Valin. Performance of MC2 and the ECMWF IFS forecast model on the Fujitsu VPP700 and NEC SX-4M. Scientific Programming, 8(1):23–30, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C3%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Ding:2013:CDF

[DZKS13] Wei Ding, Yuanrui Zhang, Mahmut Kandemir, and Seung Woo Son. Compilerdirected file layout optimization for hierarchical storage systems. Scientific Programming, 21(3–4):65– 78, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Edjlali:1997:RTC

[EAS^{+}97] Guy Edjlali, Gagan Agrawal, Alan Sussman, Jim Humphries, and Joel Saltz. Run-time and compiler support for programming in adaptive parallel environments. Scientific Programming, 6(2):215– 227, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Edwards:2012:MPP

[ESP^{+}12] H. Carter Edwards, Daniel Sunderland, Vicki Porter, Chris Amsler, and Sam Mish. Manycore performance-portability: Kokkos multidimensional array library. Scientific Programming, 20(2):89–114, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Eide:1999:FIC

[ESSL99] Eric Eide, James L. Simister, Tim Stack, and Jay Lepreau. Flexible IDL compilation for complex communication patterns. Scientific Programming, 7(3– 4):275–287, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C7%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Ferner:1999:ACS

[FB99] Clayton S. Ferner and Robert G. Babb II. Automatic choice of scheduling heuristics for parallel/ distributed computing. Scientific Programming, 7(1): 47–65, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C4%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Faure:2001:CGS

[FC01] Christèle Faure and Isabelle Charpentier. Comparing global strategies for coding adjoints. Scientific Programming, 9(1):1–10, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C1%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Fdez-Glez:2016:WNF

[FGROL^{+}16] J. Fdez-Glez, D. Ruano-Ordás, R. Laza, J. R. Méndez, R. Pavón, and F. Fdez-Riverola. WSF2: A novel framework for filtering Web spam. Scientific Programming, 2016 (??):6091385:1–6091385:18, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/6091385/.

Filippone:2014:BR

[Fil14] Salvatore Filippone. Book review. Scientific Programming, 22(1):53–55, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Faisal:2016:AEO

[FMA16] Mohammed Faisal, Hassan Mathkour, and Mansour Alsulaiman. AntStar: Enhancing optimization problems by integrating an ant system and algorithm. Scientific Programming, 2016 (??):5136327:1–5136327:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/5136327/.

Fursin:2014:CMT

[FML^{+}14] Grigori Fursin, Renato Miceli, Anton Lokhmotov, Michael Gerndt, Marc Baboulin, Allen D. Malony, Zbigniew Chamski, Diego Novillo, and Davide Del Vento. Collective mind: Towards practical and collaborative auto-tuning. Scientific Programming, 22 (4):309–329, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Fitzgerald:1996:UPA

[FO96] Steven M. Fitzgerald and Rodney R. Oldehoeft. Updatein-place analysis for true multidimensional arrays. Scientific Programming, 5(2):147– 160, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Foster:1992:PPP

[FOT92] I. Foster, R. Olson, and S. Tuecke. Productive parallel programming: The PCN approach. Scientific Programming, 1(1):51– 66, Fall 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ford:1995:PPA

[FP95] R. W. Ford and D. I. A.

Poll. Parallel processing approach to transition prediction for laminar flow control system design. Scientific Programming, 4(3):203– 217, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Fahringer:2000:PPE

[FP00a] T. Fahringer and A. Požgaj.

P^{3}T+: a performance estimator for distributed and parallel programs. Scientific Programming, 8(2): 73–93, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=253x52trrm4r87tkuw1h%
26referrer=parent%26backto=
issue%2C2%2C3%3Bjournal%
2C4%2C9%3Blinkingpublicationresults%
2C1%2C1.

Francioni:2000:DSH

[FP00b] Joan M. Francioni and Cherri M. Pancake. A debugging standard for highperformance computing. Scientific Programming, 8(2): 95–108, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=253x52trrm4r87tkuw1h%

                                        26referrer=parent%26backto=
                                        issue%2C3%2C3%3Bjournal%
                                        2C4%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Fatoohi:2001:DID

[FS01] Rod Fatoohi and Lance Smith. Development and implementation of a distributedobject job-execution environment. Scientific Programming, 9(1):27–37, ???? 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=f2779jvvqg63jq64qwtm%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C6%3Bjournal%
                                        2C2%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Fang:2014:API

[FSV14] Jianbin Fang, Henk Sips, and Ana Lucia Varbanescu. Aristotle: A performance impact indicator for the OpenCL kernels using local memory. Scientific Programming, 22 (3):239–257, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gustafson:1996:MPS

[GA96] John Gustafson and Srinivas Aluru. Massively parallel searching for better algorithms or how to do a cross product with five multiplications. Scientific Programming, 5(3):203–217, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gerber:1994:NCS

[GB94] Stefan Gerber and Helmar Burkhart. Number-crunching software and the input problem: guidelines and a case study. Scientific Programming, 3(1):1–11, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gundy-Burlet:2010:BR

[GB10] Karen Gundy-Burlet. Book review. Scientific Programming, 18(3–4):219–220, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ghayekhloo:2015:PSI

[GB15] Samira Ghayekhloo and Zeki Bayram. Prefiltering strategy to improve performance of Semantic Web service discovery. Scientific Programming, 2015(??):576463:1– 576463:15, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/576463/

.

Gerstenberger:2014:EHS

[GBH14] Robert Gerstenberger, Maciej Besta, and Torsten Hoefler. Enabling highlyscalable remote memory access programming with MPI-3 One Sided. Scientific Programming, 22(2):75–91, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

George:1994:PCS

[GBJ94] William George, Ralph G. Brickner, and S. Lennart Johnsson. Polyshift communications software for the Connection Machine systems CM-2 and CM-200. Scientific Programming, 3(1):83– 99, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gunzinger:1996:PEH

[GBK^{+}96] A. Gunzinger, B. Baumle, M. Klebl, M. Kocheisen, P. Kohler, R. Morel, U. Muller, and M. Rosenthal. Programming environment for a high-performance parallel supercomputer with intelligent communication. Scientific Programming, 5(1):25– 32, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gerndt:2002:GEP

[Ger02]  Michael Gerndt.  Guest-

editorial: PADDA 2001. Scientific Programming, 10 (1):1–2, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C1%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

GautierdeLahaut:1995:SAC

[GG95] Damien Gautier de Lahaut and Cecile Germain. Static approach for compiling communications in parallel scientific programs. Scientific Programming, 4(4):291– ??, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gravvanis:2005:PPC

[GG05] George A. Gravvanis and Konstantinos M. Giannoutakis. Parallel preconditioned conjugate gradient square method based on normalized approximate inverses. Scientific Programming, 13(2):79– 91, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Goldman:2004:IBS

[GGJ04] Jacki P. Goldman, William J.

Gullick, and Colin G. Johnson. Individual-based simulation of the clustering behaviour of epidermal growth factor receptors. Scientific Programming, 12(1):25– 43, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Getov:1999:MLP

[GGMS99] Vladimir Getov, Paul Gray, Sava Mintchev, and Vaidy Sunderam. Multi-language programming environments for high performance Java computing. Scientific Programming, 7(2):139–146, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C8%3Bjournal%
                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Gschwind:2009:HPC

[GGP09] Michael Gschwind, Fred Gustavson, and Jan F. Prins. High performance computing with the Cell Broadband Engine. Scientific Programming, 17(1–2):1–2, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Greer:2002:SCI

[GHH^{+}02] Bruce Greer, John Harrison, Greg Henry, et al. Scientific computing on the Itanium(R) processor. Scientific Programming, 10(4):

329–337, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gaidamour:2012:DCF

[GHST12] Jérémie Gaidamour, Jonathan Hu, Chris Siefert, and Ray Tuminaro. Design considerations for a flexible multigrid preconditioning library. Scientific Programming, 20 (3):223–239, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Grant:2000:CFP

[GHW00] Philip W. Grant, Magne Haveraaen, and Michael F. Webster. Coordinate free programming of computational fluid dynamics problems. Scientific Programming, 8(4):211–230, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://iospress.metapress.
com/app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C2%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Glikman:1995:PAM

[GIKP95] Eli Glikman, Ludmila Ioffe, Itzhak Kelson, and Shlomit S. Pinter. Parallel algorithms for molecular dynamics simulation of irradiation effects in crystals. Scientific Programming, 4(3):185–??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gil:2009:DKD

[Gil09] Yolanda Gil. From data to knowledge to discoveries: Artificial intelligence and scientific workflows. Scientific Programming, 17(3): 231–246, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gennart:1996:GVM

[GKL^{+}96] B. A. Gennart, B. Krummenacher, L. Landron, R. D. Hersch, B. Saugy, J.-C. Hadorn, and D. Muller. Giga view multiprocessor multidisk image server. Scientific Programming, 5(1):3– 13, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gerbessiotis:2004:RMA

[GL04] Alexandros V. Gerbessiotis and Seung-Yeop Lee. Remote memory access: a case for portable, efficient and library independent parallel programming. Scientific Programming, 12(3): 169–183, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gropp:2014:SIS

[GM14] William Gropp and Satoshi Matsuoka. Special issue: SC13 — The International Conference for High Performance Computing, Networking, Storage and Analysis. Scientific Programming, 22(2):57–58, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gao:2015:LPM

[GMCT15] Fei Gao, Katarzyna Musial, Colin Cooper, and Sophia Tsoka. Link prediction methods and their accuracy for different social networks and network metrics. Scientific Programming, 2015(??): 172879:1–172879:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/172879/

.

Golumbic:1993:ISA

[GR93] Martin Charles Golumbic and Vladimir Rainish. Instruction scheduling across control flow. Scientific Programming, 2(3):1–5, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gomes:2003:POG

[GRC03] Maria Cecília Gomes, Omer F.

Rana, and José C. Cunha. Pattern operators for grid environments. Scientific Programming, 11(3):237–261, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Germain-Renaud:1999:JBC

[GRN99] Cécile Germain-Renaud and Vincent Néri. Java-based coupling for parallel predictiveadaptive domain decomposition. Scientific Programming, 7(2):185–189, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C8%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Gonzalez:2003:TPM

[GRR^{+}03] J. A. González, C. Rodríguez, G. Rodríguez, et al. A tool for performance modeling of parallel programs. Scientific Programming, 11 (3):191–198, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gregor:2003:DPL

[GSM03] Douglas Gregor, Sibylle Schupp, and David R. Musser. Design patterns for library optimization. Scientific Programming, 11(4): 309–320, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gustafsson:1993:OOI

[Gus93] Kjell Gustafsson. Objectoriented implementation of software for solving ordinary differential equations. Scientific Programming, 2 (4):217–225, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gaurav:2011:UGC

[GW11] Gaurav and Steven F. Wojtkiewicz. Use of GPU computing for uncertainty quantification in computational mechanics: a case study. Scientific Programming, 19 (4):199–212, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Gao:2016:VCD

[GWZ16] Xianming Gao, Baosheng Wang, and Xiaozhe Zhang. VR-Cluster: Dynamic migration for resource fragmentation problem in virtual router platform. Scientific Programming, 2016 (??):3976965:1–3976965:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/3976965/.

Haveraaen:2000:CSA

[Hav00a]  Magne Haveraaen.  Case

study on algebraic software methodologies for scientific computing. Scientific Programming, 8(4):261– 273, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C5%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Haveraaen:2000:MCA

[Hav00b]  Magne Haveraaen.  Ma-

chine and collection abstractions for user-implemented data-parallel programming. Scientific Programming, 8 (4):231–246, ???? 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=2dyrcfdc1p5ytlerkc3u%
26referrer=parent%26backto=
issue%2C3%2C5%3Bjournal%
2C3%2C9%3Blinkingpublicationresults%
2C1%2C1.

Hockney:1994:PRP

[HB94] R. Hockney and M. Berry. Parkbench report: public international benchmarks for parallel computers. Scientific Programming, 3(2): iii, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:2012:MLV

[HB12] Zhen Hu and Raj Bhatnagar. Mining low-variance biclusters to discover coregulation modules in sequencing datasets. Scientific Programming, 20(1):15–27, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hind:1994:ESP

[HBCM94] Michael Hind, Michael Burke, Paul Carini, and Sam Midkiff. An empirical study of precise interprocedural array analysis. Scientific Programming, 3(3):255–271, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://www.mcs.newpaltz.
edu/~hind.

Hernandez:2008:PTM

[HCJ08] Oscar Hernandez, Barbara Chapman, and Haoqiang Jin. A performance tuning methodology with compiler support. Scientific Programming, 16(2–3):135– 153, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hemler:2000:KEU

[Hem00] Richard S. Hemler. Key elements of the user-friendly, GFDL SKYHI general circulation model. Scientific Programming, 8(1):39–47, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C5%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Huang:2016:RLT

[HHD^{+}16] Chenlin Huang, Chuanwang Hou, Huadong Dai, Yan Ding, Songling Fu, and Mengluo Ji. Research on Linux trusted boot method based on reverse integrity verification. Scientific Programming, 2016 (??):4516596:1–4516596:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/4516596/.

Herrera:2005:PSA

[HHML05] J. Herrera, E. Huedo, R. S. Montero, and I. M. Llorente. Porting of scientific applications to Grid Computing on GridWay. Scientific Programming, 13(4): 317–331, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Forum:1993:HPF

[Hig93] High Performance Fortran Forum. High Performance Fortran language specification. Scientific Programming, 2(1–2):1–170, Spring– Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hillman:1997:RSE

[Hil97] L. Hillman. Review: Scientist’s and engineer’s guide to workstations and supercomputers: coping with Unix, RISC, vectors, and programming, by Rubin H. Landau and Paul J. Fink, Jr. Scientific Programming, 6 (4):391–393, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hu:1996:IBA

[HJ96] Y. Charlie Hu and S. Lennart Johnsson. Implementing O(N) N-body algorithms efficiently in data-parallel languages. Scientific Programming, 5(4):337–364, 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ha:2015:EAF

[HJ15] Ok-Kyoon Ha and Yong-Kee Jun. An efficient algorithm for on-the-fly data race detection using an epochbased technique. Scientific Programming, 2015(??): 205827:1–205827:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/205827/

.

Huang:2010:ELA

[HJYC10] Lei Huang, Haoqiang Jin, Liqi Yi, and Barbara Chapman. Enabling locality-aware computations in OpenMP. Scientific Programming, 18 (3–4):169–181, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Howle:2012:PHP

[HKL^{+}12] Victoria E. Howle, Robert C. Kirby, Kevin Long, Brian Brennan, and Kimberly Kennedy. Playa: Highperformance programmable linear algebra. Scientific Programming, 20(3):257–273, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huedo:2005:DEI

[HLM^{+}05] E. Huedo, A. Lepinette, R. S. Montero, I. M. Llorente, and L. Vázquez. Development and execution of an impact cratering application on a computational Grid. Scientific Programming, 13(1):19– 30, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hill:2011:EOP

[HLM^{+}11] Zach Hill, Jie Li, Ming Mao, Arkaitz Ruiz-Alvarez, and Marty Humphrey. Early observations on the performance of Windows Azure. Scientific Programming, 19 (2–3):121–132, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hong:2016:AAQ

[HLPY16] Cheol-Ho Hong, Kyungwoon Lee, Hyunchan Park, and Chuck Yoo. ANCS: Achieving QoS through dynamic allocation of network resources in virtualized clouds. Scientific Programming, 2016 (??):4708195:1–4708195:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/4708195/.

Haines:1997:DPP

[HMC97] Matthew Haines, Piyush Mehrotra, and David Cronk. Data-parallel programming in a multithreaded environment. Scientific Programming, 6(2):187–200, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hill:2007:ISC

[HMCH07] Chris Hill, Dimitris Menemenlis, Bob Ciotti, and Chris Henze. Investigating solution convergence in a global ocean model using a 2048-processor cluster of distributed shared memory machines. Scientific Programming, 15(2): 107–115, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Haveraaen:2015:HPD

[HMR^{+}15] Magne Haveraaen, Karla Morris, Damian Rouson, Hari Radhakrishnan, and Clayton Carson. Highperformance design patterns for modern Fortran. Scientific Programming, 2015(??): 942059:1–942059:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2015/942059/

.

Huck:2008:KSA

[HMSM08] Kevin A. Huck, Allen D. Malony, Sameer Shende, and Alan Morris. Knowledge support and automation for performance analysis with PerfExplorer 2.0. Scientific Programming, 16(2–3):123– 134, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hansen:1992:OOH

[HMSW92] D. Hansen, D. Maier, J. Stanley, and J. Walpole. Objectoriented heterogeneous database for materials science. Scientific Programming, 1(2): 115–??, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hey:2002:PEP

[HP02] Tony Hey and Juri Papay. Performance engineering, PSEs and the GRID. Scientific Programming, 10 (1):3–17, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C2%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Hadjidoukas:2009:EFG

[HPD09] P. E. Hadjidoukas, G. Ch.

Philos, and V. V. Dimakopoulos. Exploiting finegrain thread parallelism on multicore architectures. Scientific Programming, 17(4): 309–323, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hilbrich:2013:MRE

[HPS^{+}13] Tobias Hilbrich, Joachim Protze, Martin Schulz, Bronis R. de Supinski, and Matthias S. Müller. MPI runtime error detection with MUST: Advances in deadlock detection. Scientific Programming, 21(3–4):109– 121, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

He:2016:DSS

[HPZ16] Hua He, Shanchen Pang, and Zenghua Zhao. Dynamic scalable stochastic Petri net: A novel model for designing and analysis of resource scheduling in cloud computing. Scientific Programming, 2016 (??):9259248:1–9259248:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/9259248/.

Heroux:2012:SITa

[HR12a] Michael A. Heroux and Damian Rouson. Special issue on the Trilinos project, Part 1 of 2. Scientific Programming, 20(2): 81, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Heroux:2012:SITb

[HR12b] Michael A. Heroux and Damian Rouson. Special issue on the Trilinos project, Part 2 of 2. Scientific Programming, 20(3): 221, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hwang:2003:IDU

[HS03] Yuan-Shin Hwang and Joel Saltz. Interprocedural definition-use chains of dynamic pointer-linked data structures. Scientific Programming, 11(1):3–37, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hamilton:1993:FOO

[HST^{+}93] Lisa Hamilton, Mark Stalzer, R. Steven Turley, John Visher, and Stephen Wandzura. Fastscat: an object-oriented program for fast scattering computation. Scientific Programming, 2(4):171– 178, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Hascoet:2008:CAR

[HUN08] L. Hascoët, J. Utke, and U. Naumann. Cheaper adjoints by reversing address computations. Scientific Programming, 16(1):81– 92, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Heroux:2012:NOT

[HW12] Michael A. Heroux and James M. Willenbring. A new overview of the Trilinos project. Scientific Programming, 20(2):83–88, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Huang:2016:ETF

[HXX16] Ling Huang, Xuhuan Xie, and Wenbo Xie. Eventtriggered faults tolerant control for stochastic systems with time delays. Scientific Programming, 2016 (??):5606234:1–5606234:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/5606234/.

Hongjiu:2015:EMC

[HY15] Liu Hongjiu and Hu Yanrong. An evaluating method with combined assigning-weight based on maximizing variance. Scientific Programming, 2015(??):290379:1– 290379:8, ???? 2015. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/290379/

.

Hempel:1999:AMP

[HZ99] Rolf Hempel and Falk Zimmermann. Automatic migration from PARMACS to MPI in parallel Fortran applications. Scientific Programming, 7(1):39–46, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C3%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Ibrahim:2009:ESD

[IB09] Khaled Z. Ibrahim and François Bodin. Efficient SIMDization and data management of the Lattice QCD computation on the Cell Broadband Engine. Scientific Programming, 17(1–2): 153–172, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ierotheou:2001:SAP

[IJL^{+}01] C. S. Ierotheou, S. P.

Johnson, P. F. Leggett, M. Cross, E. W. Evans, H. Jin, M. Frumkin, and J. Yan. The semi-automatic parallelisation of scientific application codes using a computer aided parallelisation toolkit. Scientific Programming, 9(2–3):163– 173, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C9%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Islam:2013:MSC

[IMB^{+}13] Tanzima Zerin Islam, Kathryn Mohror, Saurabh Bagchi, Adam Moody, Bronis R. de Supinski, and Rudolf Eigenmann. McrEngine: a scalable checkpointing system using data-aware aggregation and compression. Scientific Programming, 21 (3–4):149–163, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Iyengar:2002:AWS

[IR02] Arun Iyengar and Daniela Rosu. Architecting Web sites for high performance. Scientific Programming, 10 (1):75–89, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C8%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Ioannidis:1999:CCR

[IRSD99] Sotiris Ioannidis, Umit Rencuzogullari, Robert Stets, and Sandhya Dwarkadas. CRAUL: Compiler and runtime integration for adaptation under load. Scientific Programming, 7(3– 4):261–273, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C6%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Iosup:2008:IOG

[ITF^{+}08] Alexandru Iosup, Todd Tannenbaum, Matthew Farrellee, Dick Epema, and Miron Livny. Inter-operating grids through Delegated Match-Making. Scientific Programming, 16(2–3):233–253, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Janssens:2015:FEA

[JBLB15] Bart Janssens, Támas Bányai, Karim Limam, and Walter Bosschaerts. Finite element assembly using an embedded domain specific language. Scientific Programming, 2015(??):797325:1– 797325:22, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/

                                        journals/sp/2015/797325/

.

Juve:2010:ERP

[JDVM10] Gideon Juve, Ewa Deelman, Karan Vahi, and Gaurang Mehta. Experiences with resource provisioning for scientific workflows using Corral. Scientific Programming, 18(2):77–92, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jacob:2007:PMC

[JEM07] Joseph C. Jacob, Peter Eisenhardt, and David Makovoz. Parallel MOPEX: Computing mosaics of large-area Spitzer surveys on a cluster computer. Scientific Programming, 15(2):75–81, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jespersen:2010:ACC

[Jes10] Dennis C. Jespersen. Acceleration of a CFD code with a GPU. Scientific Programming, 18(3–4):193– 201, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jiang:2016:FAF

[JFPL16] Jianhua Jiang, Yunzhao Feng, Milan Parmar, and Keqin Li. FP-ABC: Fast and parallel ABC based energyefficiency live VM allocation policy in data centers. Scientific Programming, 2016(??): 9524379:1–9524379:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/9524379/

.

Jolivet:2014:SDD

[JHNP14] Pierre Jolivet, Frédéric Hecht, Frédéric Nataf, and Christophe Prud’homme. Scalable domain decomposition preconditioners for heterogeneous elliptic problems. Scientific Programming, 22(2): 157–171, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jung:2014:EFC

[JIC^{+}14] Myoungsoo Jung, Ellis H. Wilson III, Wonil Choi, John Shalf, Hasan Metin Aktulga, Chao Yang, Erik Saule, Umit V. Catalyurek, and Mahmut Kandemir. Exploring the future of out-of-core computing with computelocal non-volatile memory. Scientific Programming, 22 (2):125–139, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jin:2003:AMP

[JJY^{+}03] Haoqiang Jin, Gabriele Jost, Jerry Yan, et al. Automatic multilevel parallelization using OpenMP. Scientific Programming, 11(2): 177–190, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jost:2010:SIE

[JK10] Gabriele Jost and Alice Koniges. Special issue: Exploring languages for expressing medium to massive onchip parallelism. Scientific Programming, 18(3–4):125– 126, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jacobson:1992:ADD

[JKR92] P. Jacobson, B. Kågström, and M. Ränner. Algorithm development for distributed memory multicomputers using CONLAB. Scientific Programming, 1(??): 185–203, 1992. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jackson:2011:PCA

[JMR^{+}11] Keith R. Jackson, Krishna Muriki, Lavanya Ramakrishnan, Karl J. Runge, and Rollin C. Thomas. Performance and cost analysis of the Supernova factory on the Amazon AWS cloud. Scientific Programming, 19 (2–3):107–119, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Jost:2010:EUH

[JR10] Gabriele Jost and Bob Robins. Experiences using hybrid MPI/OpenMP in the real world: Parallelization of a 3D CFD solver for multi-core node clusters. Scientific Programming, 18 (3–4):127–138, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kosar:2011:PRE

[KABW11] Tevfik Kosar, Ismail Akturk, Mehmet Balman, and Xinqi Wang. PetaShare: a reliable, efficient and transparent distributed storage management system. Scientific Programming, 19(1): 27–43, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kazienko:2015:CAS

[KAS15] Przemysław Kazienko, Reda Alhajj, and Jaideep Srivastava. Computational aspects of social network analysis. Scientific Programming, 2015(??):961610:1–961610:2, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2015/961610/.

Kohn:1996:ICG

[KB96] Scott R. Kohn and Scott B. Baden. Irregular coarsegrain data parallelism under LPARX. Scientific Programming, 5(3):185–201, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kaiser:2001:OCC

[KB01] Timothy H. Kaiser and Scott B. Baden. Overlapping communication and computation with OpenMP and MPI. Scientific Programming, 9(2–3):73–81, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=7pab6qgbaf8vxg991rwy%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C11%3Bjournal%
                                        2C1%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Krishnan:2002:XSP

[KBG^{+}02] Sriram Krishnan, Randall Bramley, Dennis Gannon, et al. The XCAT Science Portal. Scientific Programming, 10(4):303–317, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Korn:1995:PTD

[KBRS95] C. Falco Korn, J. M. Bull, G. D. Riley, and P. K.

Stansby. Parallelization of a three-dimensional shallowwater estuary model on the KSR-1. Scientific Programming, 4(3):155–169, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kenny:2005:CIM

[KCO^{+}05] Stuart Kenny, Brian Coghlan, David O’Callaghan, John Ryan, Rob Byrom, Laurence Field, Steve Hicks, Manish Soni, Antony Wilson, Xiaomei Zhu, Roney Cordenonsi, Ari Datta, Linda Cornwall, Abdeslem Djaoui, and Norbert Podhorszki. The CanonicalProducer: an instrument monitoring component of the Relational Grid Monitoring Architecture (R-GMA). Scientific Programming, 13(2):151–158, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kuzmin:2015:PSS

[KCS15] Konstantin Kuzmin, Mingming Chen, and Bolesław K. Szymanski. Parallelizing SLPA for scalable overlapping community detection. Scientific Programming, 2015(??):461362:1– 461362:18, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/

                                        journals/sp/2015/461362/

.

Kurzak:2009:QFC

[KD09] Jakub Kurzak and Jack Dongarra. QR factorization for the Cell Broadband Engine. Scientific Programming, 17 (1–2):31–42, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kim:2011:AMA

[KeKR^{+}11] Hyunjoo Kim, Yaakoub el Khamra, Ivan Rodero, Shantenu Jha, and Manish Parashar. Autonomic management of application workflows on hybrid computing infrastructure. Scientific Programming, 19(2–3): 75–89, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kessler:1996:PDA

[Kes96] Christoph W. Kessler. Patterndriven automatic parallelization. Scientific Programming, 5(3):251–274, Fall 1996. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Keahey:2005:VWA

[KFFZ05] K. Keahey, I. Foster, T. Freeman, and X. Zhang. Virtual workspaces: Achieving quality of service and quality of life in the Grid. Scientific Programming, 13 (4):265–275, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

King:2008:SSL

[KG08] Brian R. King and Chittibabu Guda. Semi-supervised learning for classification of protein sequence data. Scientific Programming, 16(1): 5–29, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kistler:2009:PLB

[KGBB09] Michael Kistler, John Gunnels, Daniel Brokenshire, and Brad Benton. Programming the Linpack benchmark for the IBM Power XCell 8i processor. Scientific Programming, 17(1–2): 43–57, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kiran:2015:GSH

[KGMN15] D. C. Kiran, S. Gurunarayanan, Janardan Prasad Misra, and Abhijeet Nawal. Global scheduling heuristics for multicore architecture. Scientific Programming, 2015(??): 860891:1–860891:12, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/860891/

.

Koppler:1997:VDD

[KGV97] Rainer Koppler, Siegfried Grabner, and Jens Volkert. Visualization of distributed data structures for High Performance Fortranlike languages. Scientific Programming, 6(1):115–126, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kumar:1995:TPF

[KHSJ95] B. Kumar, C.-H. Huang, P. Sadayappan, and R. W. Johnson. Tensor product formulation of Strassen’s matrix multiplication algorithm with memory reduction. Scientific Programming, 4(4):275– ??, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kenny:2008:CAC

[KJG^{+}08] Joseph P. Kenny, Curtis L. Janssen, Mark S. Gordon, Masha Sosonkina, and Theresa L. Windus. A component approach to collaborative scientific software development: Tools and techniques utilized by the Quantum Chemistry Science Application Partnership. Scientific Programming, 16(4): 287–296, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Knepley:2009:MAP

[KK09] Matthew G. Knepley and Dmitry A. Karpeev. Mesh algorithms for PDE with Sieve I: Mesh distribution. Scientific Programming, 17 (3):215–230, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kunzman:2011:PHC

[KK11] David M. Kunzman and Laxmikant V. Kalé. Programming heterogeneous clusters with accelerators using object-based programming. Scientific Programming, 19 (1):47–62, ???? 2011. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kurowski:2004:DGS

[KLN^{+}04] K. Kurowski, B. Ludwiczak, J. Nabrzyski, A. Oleksiak, and J. Pukacki. Dynamic Grid scheduling with job migration and rescheduling in the GridLab resource management system. Scientific Programming, 12(4): 263–273, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kropf:1996:SPC

[KLS^{+}96] Peter G. Kropf, Edgar F. A. Lederer, Thomas Steffen, Karl Guggisberg, Jean-Guy Schneider, and Peter Schwab. SPINET: a parallel computing approach to spine simulations. Scientific Programming, 5(1):15– ??, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kindratenko:2009:ITP

[KMB09] Volodymyr V. Kindratenko, Adam D. Myers, and Robert J. Brunner. Implementation of the two-point angular correlation function on a high-performance reconfigurable computer. Scientific Programming, 17(3): 247–259, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kamachi:1997:KPH

[KMR^{+}97] T. Kamachi, A. Muller, R. Ruhl, Y. Seo, K. Suehiro, and M. Tamura. Kemari: a portable High Performance Fortran system for distributed memory parallel processors. Scientific Programming, 6(1):41– 58, Spring 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Koko:2007:VMC

[Kok07] Jonas Koko. Vectorized Matlab codes for linear twodimensional elasticity. Scientific Programming, 15(3): 157–172, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Knies:1994:HPF

[KOM94] Allan Knies, Matthew O’Keefe, and Tom MacDonald. High Performance Fortran: a practical analysis. Scientific Programming, 3(3):187– 199, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kholod:2016:CDD

[KPS16] Ivan Kholod, Ilya Petukhov, and Andrey Shorov. Cloud for distributed data analysis based on the actor model. Scientific Programming, 2016 (??):1050293:1–1050293:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/1050293/.

Kalaivani:2015:FRB

[KS15] P. Kalaivani and K. L. Shunmuganathan. Feature reduction based on genetic algorithm and hybrid model for opinion mining. Scientific Programming, 2015(??): 961454:1–961454:15, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/961454/

.

Katsaloulis:2005:SAL

[KTP05] P. Katsaloulis, T. Theoharis, and A. Provata. Statistical algorithms for long DNA sequences: Oligonucleotide distributions and homogeneity maps. Scientific Programming, 13(3): 177–188, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kufel:2015:NLS

[Kuf15] Lukasz Kufel. Network latency in systems event monitoring for multiple locations. Scientific Programming, 2015 (??):371620:1–371620:6, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2015/371620/

.

Korkhov:2007:VGI

[KVW^{+}07] Vladimir Korkhov, Dmitry Vasyunin, Adianto Wibisono, Adam S. Z. Belloum, Márcia A. Inda, Marco Roos, Timo M. Breit, and L. O. Hertzberger. VLAM-G: Interactive data driven workflow engine for Grid-enabled resources. Scientific Programming, 15(3): 173–188, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kohr:1994:POO

[KZRR94] David R. Kohr, Jr., Xingbin Zhang, Mustafizur Rahman, and Daniel A. Reed. Performance of an object-oriented, parallel operating system. Scientific Programming, 3 (4):301–324, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lancaster:2001:PSE

[Lan01] David Lancaster. A problem solving environment based on CORBA. Scientific Programming, 9(4):233–242, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C4%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Landry:2003:IHP

[Lan03] Walter Landry. Implementing a high performance tensor library. Scientific Programming, 11(4):273–290, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Larrabee:1993:PPP

[Lar93] Allan R. Larrabee. P4 parallel programming system, the Linda environment, and some experiences with parallel computation. Scientific Programming, 2(3):23– 35, Fall 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Li:2002:BAS

[LB02] Cen Li and Gautam Biswas. A Bayesian approach for structural learning with hidden Markov models. Scientific Programming, 10(3): 201–219, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Long:2012:SHL

[LBvBW12] Kevin Long, Paul T. Boggs, and Bart G. van Bloemen Waanders. Sundance: High-level software for PDEconstrained optimization. Scientific Programming, 20 (3):293–310, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lemeire:2007:CAP

[LDV07] Jan Lemeire, Erik Dirkx, and Frederik Verbist. Causal analysis for performance modeling of computer programs. Scientific Programming, 15(3):121–136, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lesk:2010:BR

[Les10] Michael Lesk. Book review. Scientific Programming, 18 (3–4):221–223, ???? 2010.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lu:2003:MLH

[LG03] Quanming Lu and Vladimir Getov. Mixed-language highperformance computing for plasma simulations. Scientific Programming, 11(1): 57–66, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lagos:2016:MAC

[LGC^{+}16] Carolina Lagos, Guillermo Guerrero, Enrique Cabrera, Stefanie Niklander, Franklin Johnson, Fernando Paredes, and Jorge Vega. A matheuristic approach combining local search and mathematical programming. Scientific Programming, 2016(??): 1506084:1–1506084:7, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/1506084/

.

Lickly:1993:CMP

[LH93] Daniel J. Lickly and Philip J.

Hatcher. C++ and massively parallel computers. Scientific Programming, 2 (4):193–202, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lin:2004:GTD

[Lin04] H. X. Lin. Graph transformation and designing parallel sparse matrix algorithms beyond data dependence analysis. Scientific Programming, 12(2):91–100, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ltaief:2010:STS

[LKDB10] Hatem Ltaief, Jakub Kurzak, Jack Dongarra, and Rosa M. Badia. Scheduling twosided transformations using tile algorithms on multicore architectures. Scientific Programming, 18(1):35– 50, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Laney:2014:AED

[LLW^{+}14] Daniel Laney, Steven Langer, Christopher Weber, Peter Lindstrom, and Al Wegener. Assessing the effects of data compression in simulations using physically motivated metrics. Scientific Programming, 22(2): 141–155, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lee:2015:OPE

[LNK^{+}15] Joo Hwan Lee, Nimit Nigania, Hyesoon Kim, Kaushik Patel, and Hyojong Kim.

OpenCL performance evaluation on modern multicore CPUs. Scientific Programming, 2015(??):859491:1– 859491:20, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/859491/

.

Labarta:2001:NOD

[LOHA01] J. Labarta, J. Oliver, D. S.

Henty, and Eduard Ayguadé. New OpenMP directives for irregular data access loops. Scientific Programming, 9(2–3):175–183, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C10%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Lin:1999:APS

[LP99] Yuan Lin and David Padua.

On the automatic parallelization of sparse and irregular Fortran programs. Scientific Programming, 7 (3–4):231–246, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%

                                        26referrer=parent%26backto=
                                        issue%2C4%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Liddell:2004:PCA

[LPHD04] Heather M. Liddell, D. Parkinson, G. S. Hodgson, and P. Dzwig. Parallel computing applications and financial modelling. Scientific Programming, 12(2):81– 90, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lastovetsky:2005:DPM

[LR05] Alexey Lastovetsky and Ravi Reddy. Data partitioning for multiprocessors with memory heterogeneity and memory constraints. Scientific Programming, 13(2):93– 112, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lam:2012:AIV

[LSB12] Ham Ching Lam, Srinand Sreevatsan, and Daniel Boley. Analyzing influenza virus sequences using binary encoding approach. Scientific Programming, 20(1):3– 13, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lan:2002:DLB

[LTB02] Zhiling Lan, Valerie E. Taylor, and Greg Bryan. Dynamic load balancing of SAMR applications on distributed systems. Scientific Programming, 10(4): 319–328, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Lin:2016:HTS

[LWW16] Weiwei Lin, Wentai Wu, and James Z. Wang. A heuristic task scheduling algorithm for heterogeneous virtual clusters. Scientific Programming, 2016 (??):7040276:1–7040276:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/7040276/.

Li:2016:NMO

[LWYS16] Yongfei Li, Shicheng Wang, Dongfang Yang, and Dawei Sun. A novel metric online monocular SLAM approach for indoor applications. Scientific Programming, 2016(??):5369780:1– 5369780:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/5369780/

.

Li:2016:ROP

[LZF16] Gang Li, Qingpu Zhang, and Zhengqian Feng. Research on optimal path of data migration among multisupercomputer centers. Scientific Programming, 2016(??): 5018213:1–5018213:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/5018213/

.

McGough:2007:GRT

[MAG^{+}07] A. Stephen McGough, Asif Akram, Li Guo, Marko Krznaric, Luke Dickens, David Colling, Janusz Martyniak, Roger Powell, Paul Kyberd, Chenxi Huang, Constantinos Kotsokalis, and Panayiotis Tsanakas. GRIDCC: a real-time Grid workflow system with QoS. Scientific Programming, 15(4): 213–234, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Malfetti:2001:AOW

[Mal01] Paolo Malfetti. Application of OpenMP to weather, wave and ocean codes. Scientific Programming, 9(2–3):99–107, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.

asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C4%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Mannarswamy:2008:BR

[Man08] Sandya S. Mannarswamy.

Book review. Scientific Programming, 16(4):341–342, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Marowka:2005:EMT

[Mar05]  Ami Marowka.  Execu-

tion model of three parallel languages: OpenMP, UPC and CAF. Scientific Programming, 13(2): 127–135, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mattson:1994:ELG

[Mat94] Timothy G. Mattson. Efficiency of Linda for general purpose scientific programming. Scientific Programming, 3(1):61–71, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mattson:2003:HGO

[Mat03] Timothy G. Mattson. How good is OpenMP. Scientific Programming, 11(2): 81–93, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

McCalpin:1996:CSS

[McC96] John D. McCalpin. A case study of some issues in the optimization of Fortran 90 array notation. Scientific Programming, 5(3):219–237, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://reality.sgi.com/
employees/mccalpin/papers/
f90.ps.

Markus:2010:CFF

[MCvM10] A. A. Markus, W. M. G. Courage, and M. C. L. M. van Mierlo. A computational framework for flood risk assessment in The Netherlands. Scientific Programming, 18 (2):93–105, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Metcalf:1999:IFCa

[Met99a]  Mike Metcalf.  Informa-

tion file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(1): 327–333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Metcalf:1999:IFCb

[Met99b]  Mike Metcalf.  Informa-

tion file on compilers, tools, books, courses, tutorials, and the standard for the Fortran language and its derivatives: Version of 20 May 1999 (the penultimate year of the millennium). Scientific Programming, 7(3– 4):327–333, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C11%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Malawski:2015:SMD

[MFB^{+}15] Maciej Malawski, Kamil Figiela, Marian Bubak, Ewa Deelman, and Jarek Nabrzyski. Scheduling multilevel deadline-constrained scientific workflows on clouds based on cost optimization. Scientific Programming, 2015(??):680271:1– 680271:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/680271/

.

Malakar:2013:DCS

[MGK^{+}13] Preeti Malakar, Thomas George, Sameer Kumar, Rashmi Mittal, Vijay Natarajan, Yogish Sabharwal, Vaibhav Saxena, and Sathish S. Vadhiyar. A divide and conquer strategy for scaling weather simulations with multiple regions of interest. Scientific Programming, 21 (3–4):93–107, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Merlin:1995:IHP

[MH95] John Merlin and Anthony Hey. An introduction to High Performance Fortran. Scientific Programming, 4 (2):87–113, Summer 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Michael:1997:RET

[Mic97]  G. A. Michael.  Re-

view: Enabling technologies for petaflops computing, by Thomas Sterling, Paul Messina, Paul H. Smith. Scientific Programming, 6 (4):395–397, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Michalakes:2000:SSP

[Mic00]  John Michalakes.  The

same-source parallel MM5. Scientific Programming, 8 (1):5–12, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C1%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Mathur:1995:AAC

[MJ95] Kapil K. Mathur and S. Lennart Johnsson. All-to-all communication on the Connection Machine CM-200. Scientific Programming, 4(4):251– 273, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Muddukrishna:2015:LAT

[MJB15] Ananya Muddukrishna, Peter A. Jonsson, and Mats Brorsson. Locality-aware task scheduling and data distribution for OpenMP programs on NUMA systems and manycore processors. Scientific Programming, 2015(??):981759:1– 981759:16, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/981759/

.

Madison:2007:PCR

[MJLM07] Richard Madison, Abhinandan Jain, Christopher Lim, and Mark Maimone. Performance characterization of a rover navigation algorithm using large-scale simulation.

Scientific Programming, 15 (2):95–105, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Muralidharan:2016:BBS

[MM16] D. Muralidharan and R. Muthaiah. Bus based synchronization method for CHIP-PER based NoC. Scientific Programming, 2016 (??):1907521:1–1907521:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/1907521/.

Moreira:2002:NJH

[MMG^{+}02] José E. Moreira, Samuel P. Midkiff, Manish Gupta, et al. NINJA: Java for high performance numerical computing. Scientific Programming, 10(1):19–33, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C3%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Morgenstern:1994:MPS

[Mor94] Craig Morgenstern. Methods for precise submesh allocation. Scientific Programming, 3(4):353–364, Winter 1994.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Morris:2015:EMI

[Mor15]  Karla Morris.  Emulat-

ing multiple inheritance in Fortran 2003/2008. Scientific Programming, 2015 (??):126069:1–126069:7, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/126069/

.

Mhadhbi:2016:ETH

[MOS16] Imene Mhadhbi, Slim Ben Othman, and Slim Ben Saoud. An efficient technique for hardware/software partitioning process in codesign. Scientific Programming, 2016 (??):6382765:1–6382765:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/6382765/.

Mossberg:1997:OOS

[MOT97] Eva Mossberg, Kurt Otto, and Michael Thune. Objectoriented software tools for the construction of preconditioners. Scientific Programming, 6(3):285–295, Fall 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Memishi:2016:FBR

[MPA16] Bunjamin Memishi, María S. Pérez, and Gabriel Antoniu. Feedback-based resource allocation in MapReducebased systems. Scientific Programming, 2016 (??):7241928:1–7241928:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/7241928/.

Miles:2004:TPA

[MPP^{+}04] Simon Miles, Juri Papay, Terry Payne, Michael Luck, and Luc Moreau. Towards a protocol for the attachment of metadata to grid service descriptions and its use in semantic discovery. Scientific Programming, 12 (4):201–211, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Moreau:2002:MOJ

[MR02] Luc Moreau and Daniel Ribbens. Mobile objects in Java. Scientific Programming, 10(1):91–100, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C9%2C9%3Bjournal%

2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Morris:2012:ECW

[MRLF12] Karla Morris, Damian W. I.

Rouson, M. Nicole Lemaster, and Salvatore Filippone. Exploring capabilities within ForTrilinos by solving the 3D Burgers equation. Scientific Programming, 20(3):275–292, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Maley:2000:CCS

[MS00] David Maley and Ivor Spence. Config: a case study in combining software engineering techniques. Scientific Programming, 8 (2):59–71, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=253x52trrm4r87tkuw1h%
26referrer=parent%26backto=
issue%2C1%2C3%3Bjournal%
2C4%2C9%3Blinkingpublicationresults%
2C1%2C1.

Martinez:2014:ELS

[MSCS14] Andrea Martínez, Anna Sikora, Eduardo César, and Joan Sorribes. ELASTIC: A large scale dynamic tuning environment. Scientific Programming, 22(4): 261–271, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Moon:1999:CCT

[MSH99] Sungdo Moon, Byoungro So, and Mary W. Hall. Combining compile-time and runtime parallelization. Scientific Programming, 7(3– 4):247–260, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=53f7mftrrm4r73yyrqau%
                                        26referrer=parent%26backto=
                                        issue%2C5%2C12%3Bjournal%
                                        2C6%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Mubarak:2013:PGA

[MSLS13] Misbah Mubarak, Seegyoung Seol, Qiukai Lu, and Mark S. Shephard. A parallel ghosting algorithm for the flexible distributed mesh database. Scientific Programming, 21 (1–2):17–42, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Manguoglu:2011:PMS

[MSSG11] Murat Manguoglu, Faisal Saied, Ahmed Sameh, and Ananth Grama. Performance models for the Spike banded linear system solver. Scientific Programming, 19(1): 13–25, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Muller:2003:OCB

[Mül03] Matthias S. Müller. An OpenMP compiler benchmark. Scientific Programming, 11(2):125–131, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Mi:2016:MCF

[MWQ16] Aizhong Mi, Lei Wang, and Junyan Qi. A multiple classifier fusion algorithm using weighted decision templates. Scientific Programming, 2016 (??):3943859:1–3943859:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/3943859/.

Nikolopoulos:2003:SNR

[NAA^{+}03] Dimitrios S. Nikolopoulos, Ernest Artiaga, Eduard Ayguadé, et al. Scaling non-regular shared-memory codes by reusing custom loop schedules. Scientific Programming, 11(2): 143–158, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagel:2004:BRI

[Nag04] Dan Nagel. Book review:

Industrial Strength Parallel Computing: Programming Massively Parallel Processors, by Alice E. Koniges. Scientific Programming, 12 (1):57–62, ???? 2004. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2005:BRH

[Nag05a]  Dan Nagle.  Book re-

view: High Performance Linux Clusters, by A. Joseph and D. Sloan. Scientific Programming, 13(2):173–175, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2005:BRM

[Nag05b] Dan Nagle. Book review: MPI — The Complete Reference, Vol. 1, The MPI Core, 2nd ed., Scientific and Engineering Computation Series, by Marc Snir, Steve Otto, Steven Huss– Lederman, David Walker and Jack Dongarra. Scientific Programming, 13(1):57– 63, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2009:BRa

[Nag09a] Dan Nagle. Book review. Scientific Programming, 17 (3):279–282, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2009:BRb

[Nag09b] Dan Nagle. Book review.

Scientific Programming, 17 (4):343–345, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2011:BRa

[Nag11a] Dan Nagle. Book review. Scientific Programming, 19(1): 67–70, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2011:BRb

[Nag11b] Dan Nagle. Book review.

Scientific Programming, 19 (4):253–258, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nagle:2012:BR

[Nag12] Dan Nagle. Book review.

Scientific Programming, 20 (3):349–353, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nanthaamornphong:2015:EUC

[NCMF15] Aziz Nanthaamornphong, Jeffrey Carver, Karla Morris, and Salvatore Filippone. Extracting UML class diagrams from object-oriented Fortran: ForUML. Scientific Programming, 2015(??): 421816:1–421816:15, ???? 2015. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/421816/

.

Norton:2007:TAM

[NDSG07] Charles D. Norton, Viktor K. Decyk, Bolesław K. Szymanski, and Henry Gardner. The transition and adoption to modern programming concepts for scientific computing in Fortran. Scientific Programming, 15(1):27– 44, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nielsen:2008:MCB

[NJ08] Ida M. B. Nielsen and Curtis L. Janssen. Multicore challenges and benefits for high performance scientific computing. Scientific Programming, 16(4): 277–285, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nakano:2002:SAS

[NKV^{+}02] Aiichiro Nakano, Rajiv K. Kalia, Priya Vashishta, et al. Scalable atomistic simulation algorithms for materials research. Scientific Programming, 10(4):263–270, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Narayanamoorthy:2016:IOF

[NM16] S. Narayanamoorthy and S. Maheswari. The intelligence of octagonal fuzzy number to determine the fuzzy critical path: A new ranking method. Scientific Programming, 2016(??): 6158208:1–6158208:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/6158208/

.

Norton:2007:HPC

[Nor07] Charles D. Norton. High performance computing for mission-enabling space applications. Scientific Programming, 15(2):71–73, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nikolopoulos:2000:TRD

[NPP^{+}00] Dimitrios S. Nikolopoulos, Theodore S. Papatheodorou, Constantine D. Polychronopoulos, et al. A transparent runtime data distribution engine for OpenMP. Scientific Programming, 8 (3):143–162, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

No:2016:MMC

[NsP16] Jaechun No and Sung soon Park. MultiCache: Multilayered cache implementation for I/O virtualization. Scientific Programming, 2016 (??):3780163:1–3780163:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/3780163/.

OBoyle:1996:EPV

[OB96] M. F. P. O’Boyle and J. M. Bull. Expert programmer versus parallelizing compiler: a comparative study of two approaches for distributed shared memory. Scientific Programming, 5(1):63– 88, Spring 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Orellana:2016:FAS

[OCC^{+}16] Julio Proaño Orellana, Blanca Caminero, Carmen Carrión, Luis Tomas, Selome Kostentinos Tesfatsion, and Johan Tordsson. FPGAaware scheduling strategies at hypervisor level in cloud environments. Scientific Programming, 2016 (??):4670271:1–4670271:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/4670271/.

Olivier:2013:CMW

[OdSSP13] Stephen L. Olivier, Bronis R. de Supinski, Martin Schulz, and Jan F. Prins. Characterizing and mitigating work time inflation in task parallel programs. Scientific Programming, 21(3–4):123– 136, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Oehmke:2000:SAA

[OHS00] Robert Oehmke, Janis Hardwick, and Quentin F. Stout. Scalable algorithms for adaptive statistical designs. Scientific Programming, 8(3): 183–193, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

OKeefe:2000:GEP

[O’K00] Matthew O’Keefe. Guesteditorial parallel software design for weather simulation codes. Scientific Programming, 8(1):1–3, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Oldehoeft:2000:BPS

[Old00] Rod Oldehoeft. Best papers from SC2000. Scientific Programming, 8(3): 109–110, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

OKeefe:1995:FPT

[OPE^{+}95] Matthew O’Keefe, Terence Parr, B. Kevin Edgar, Steve Anderson, Paul Woodward, and Hank Dietz. Fortran-P translator: towards automatic translation of Fortran 77 programs for massively parallel processors. Scientific Programming, 4(1):1– 21, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ostermann:2011:UNE

[OPP11] Simon Ostermann, Kassian Plankensteiner, and Radu Prodan. Using a new eventbased simulation framework for investigating resource provisioning in Clouds. Scientific Programming, 19(2– 3):161–178, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

OHallaron:1999:SSS

[OS99] David R. O’Hallaron and Bolesław K. Szymanski. Software systems for scalable computers. Scientific Programming, 7(3–4): 191–193, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C1%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Oldfield:2012:TST

[OSLK12] Ron A. Oldfield, Gregory D.

Sjaardema, Gerald F. Lofstead II, and Todd Kordenbrock. Trilinos I/O Support (Trios). Scientific Programming, 20(2): 181–196, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ozturan:1994:CTP

[OSS94] Can Ozturan, Balaram Sinharoy, and Bolesław K. Szymanski. Compiler technology for parallel scientific computation. Scientific Programming, 3(3):201– 225, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Otto:1993:PAC

[Ott93] Steve W. Otto. Parallel array classes and lightweight sharing mechanisms. Scientific Programming, 2(4):203– 216, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ozturan:2004:RBD

[Özt04] Can Özturan. Resource bartering in data grids. Scientific Programming, 12(3): 155–168, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

PARKBENCH:1994:PRP

[PAR94] PARKBENCH Committee/ Assembled by R. Hockney (Chairman) and M. Berry (Secretary). PARK-BENCH report: Public international benchmarks for parallel computers. Scientific Programming, 3(2):101– 146, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Plazek:2001:CMP

[PBK01] Joanna Płażek, Krzysztof Banaś, and Jacek Kitowski. Comparison of messagepassing and shared memory implementations of the GM-RES method on MIMD computers. Scientific Programming, 9(4):195–209, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=
issue%2C1%2C4%3Bjournal%
2C3%2C12%3Blinkingpublicationresults%
2C1%2C1.

Perez-Carro:2014:CJM

[PCGLO14] Pascual Pérez-Carro, Francisco Grimaldo, Miguel Lozano, and Juan M. Orduña. Characterization of the Jason multiagent platform on multicore processors. Scientific Programming, 22(1): 21–35, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prins:1999:ICF

[PCS99] Jan F. Prins, Siddhartha Chatterjee, and Martin Simons. Irregular computations in Fortran — expression and implementation strategies. Scientific Programming, 7(3–4):313–326, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C10%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Peachey:2008:FFD

[PDA^{+}08] T. C. Peachey, N. T. Diamond, D. A. Abramson, W. Sudholt, A. Michailova, and S. Amirriazi. Fractional factorial design for parameter sweep experiments using Nimrod/E. Scientific Programming, 16(2–3):217– 230, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pike:2005:IDP

[PDGQ05] Rob Pike, Sean Dorward, Robert Griesemer, and Sean Quinlan. Interpreting the data: Parallel analysis with Sawzall. Scientific Programming, 13(4):277–298, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Perez-Diaz:2016:BAC

[PDROFRM16] N. Pérez-Díaz, D. Ruano-Ordás, F. Fdez-Riverola, and J. R. Méndez. Boosting accuracy of classical machine learning antispam classifiers in real scenarios by applying rough set theory. Scientific Programming, 2016 (??):5945192:1–5945192:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/5945192/.

Perrott:2008:MMR

[Per08] Ron Perrott. Memorial: In memory of Robert Gordon Babb II. Scientific Programming, 16(1):1, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papiani:1995:GBI

[PHH95] Mark Papiani, Anthony J. G. Hey, and Roger W. Hockney. Graphical benchmark information service. Scientific Programming, 4(4):219– 227, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Preissl:2010:OCC

[PKE^{+}10] Robert Preissl, Alice Koniges, Stephan Ethier, Weixing Wang, and Nathan Wichmann. Overlapping communication with computation using OpenMP tasks on the GTS magnetic fusion code. Scientific Programming, 18 (3–4):139–151, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Plale:2004:FBD

[Pla04] Beth Plale. Framework for bringing data streams to the grid. Scientific Programming, 12(4):213–223, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Ponnusamy:1994:GCM

[PMCF94] R. Ponnusamy, N. Mansour, A. Choudhary, and G. C. Fox. Graph contraction for mapping data on parallel computers: a quality-cost tradeoff. Scientific Programming, 3(1):73– 82, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pase:1994:CFP

[PMM94] Douglas M. Pase, Tom MacDonald, and Andrew Meltzer. CRAFT Fortran programming model. Scientific Programming, 3(3):

227–253, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papadimitriou:2012:SAP

[PMTL12] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. The software architecture for performing scientific computation with the JLAPACK libraries in ScalaLab. Scientific Programming, 20(4): 379–391, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papadimitriou:2014:MLS

[PMTL14] Stergios Papadimitriou, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. MATLABlike scripting of Java scientific libraries in ScalaLab. Scientific Programming, 22 (3):187–199, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Petcu:2005:DIG

[PPD05] Dana Petcu, Marcin Paprzycki, and Diana Dubu. Design and implementation of a Grid extension for Maple. Scientific Programming, 13 (2):137–149, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Paszynska:2015:QOE

[PPJ^{+}15] A. Paszyńska, M. Paszyński, K. Jopek, M. Woźniak, D. Goik, P. Gurgul, H. AbouEisha, M. Moshkov, V. M. Calo, A. Lenharth, D. Nguyen, and K. Pingali. Quasi-optimal elimination trees for 2D grids with singularities. Scientific Programming, 2015(??): 303024:1–303024:18, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/303024/

.

Pawlowski:2012:AEAa

[PPS12a] Roger P. Pawlowski, Eric T.

Phipps, and Andrew G. Salinger. Automating embedded analysis capabilities and managing software complexity in multiphysics simulation, Part I: Templatebased generic programming. Scientific Programming, 20 (2):197–219, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pawlowski:2012:AEAb

[PPS^{+}12b] Roger P. Pawlowski, Eric T.

Phipps, Andrew G. Salinger, Steven J. Owen, Christopher M. Siefert, and Matthew L. Staten. Automating embedded analysis capabilities and managing software complexity in multiphysics simulation, Part II: Application to partial differential equations. Scientific Programming, 20 (3):327–345, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prentice:1999:LE

[Pre99] J. K. Prentice. Letter to the editors. Scientific Programming, 7(1):83–84, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Paul:2014:CEM

[PRM^{+}14] Indrani Paul, Vignesh Ravi, Srilatha Manne, Manish Arora, and Sudhakar Yalamanchili. Coordinated energy management in heterogeneous processors. Scientific Programming, 22(2):93– 108, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Prodan:2007:SRW

[Pro07] Radu Prodan. Specification and runtime workflow support in the ASKALON Grid environment. Scientific Programming, 15(4): 193–211, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Papadimitriou:2015:SGC

[PSM^{+}15] Stergios Papadimitriou, Kirsten Schwark, Seferina Mavroudi, Kostas Theofilatos, and Spiridon Likothanasis. ScalaLab and GroovyLab: Comparing Scala and Groovy for scientific computing. Scientific Programming, 2015(??): 498618:1–498618:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/498618/

.

Patnaik:2008:INN

[PSU08] Debprakash Patnaik, P. S.

Sastry, and K. P. Unnikrishnan. Inferring neuronal network connectivity from spike data: a temporal data mining approach. Scientific Programming, 16(1):49– 77, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pllana:2009:ISP

[PT09a] Sabri Pllana and Jesper Larsson Träff. Introduction to the scientific programming special issue: Software development for multi-core computing systems. Scientific Programming, 17(4): 283–284, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Pllana:2009:RSP

[PT09b] Sabri Pllana and Jesper Larsson Träff. Reviewers for scientific programming special issue on software development for multi-core computing systems. Scientific Programming, 17(4): 337, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Park:2013:EBB

[PTS^{+}13] Jongsoo Park, Ping Tak Peter Tang, Mikhail Smelyanskiy, Daehyun Kim, and Thomas Benson. Efficient backprojection-based synthetic aperture radar computation with many-core processors. Scientific Programming, 21(3–4):165–179, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Park:2001:PPE

[PVKE01] Insung Park, Michael J. Voss, Seon Wook Kim, and Rudolf Eigenmann. Parallel programming environment for OpenMP. Scientific Programming, 9(2–3):143– 161, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C8%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Parent:2004:ALB

[PVL^{+}04] Johan Parent, Katja Verbeeck, Jan Lemeire, Ann Nowe, Kris Steenhaut, and Erik Dirkx. Adaptive load balancing of parallel applications with multi-agent reinforcement learning on heterogeneous systems. Scientific Programming, 12(2):71– 79, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rehg:1999:ITD

[RKR^{+}99] James M. Rehg, Kathleen Knobe, Umakishore Ramachandran, Rishiyur S. Nikhil, and Arun Chauhan. Integrated task and data parallel support for dynamic applications. Scientific Programming, 7(3– 4):289–302, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=53f7mftrrm4r73yyrqau%
26referrer=parent%26backto=
issue%2C8%2C12%3Bjournal%
2C6%2C9%3Blinkingpublicationresults%
2C1%2C1.

Rommelse:2004:EAC

[RLC04] J. R. Rommelse, H. X. Lin, and T. F. Chan. Efficient active contour and K-means algorithms in image segmentation. Scientific Programming, 12(2):101–120, ???? 2004.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rackl:2002:EBL

[RLL^{+}02] Günther Rackl, Thomas Ludwig, Markus Lindermeier, et al. Efficiently building on-line tools for distributed heterogeneous environments. Scientific Programming, 10(1):67–74, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=9ejnuvwuvby9737jte27%
                                        26referrer=parent%26backto=
                                        issue%2C7%2C9%3Bjournal%
                                        2C2%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Rouson:2005:DMA

[RMX05] Damian W. I. Rouson, Karla Morris, and Xiaofeng Xu. Dynamic memory deallocation in Fortran 95/2003 derived type calculus. Scientific Programming, 13(3): 189–203, ???? 2005. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Reid:2007:CAN

[RN07] John Reid and Robert W. Numrich. Co-arrays in the next Fortran Standard. Scientific Programming, 15(1): 9–26, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Romberg:2002:UGI

[Rom02] Mathilde Romberg. The UNICORE Grid infrastructure. Scientific Programming, 10(2):149–157, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
26referrer=parent%26backto=
issue%2C6%2C8%3Bjournal%
2C1%2C12%3Blinkingpublicationresults%
2C1%2C1.

Rosmond:2000:SVN

[Ros00] Thomas E. Rosmond. A scalable version of the Navy Operational Global Atmospheric Prediction System spectral forecast model. Scientific Programming, 8(1): 31–38, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C4%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Rouson:2008:CSC

[Rou08a] Damian W. I. Rouson. Complexity in scalable computing. Scientific Programming, 16(4):275–276, ???? 2008.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rouson:2008:TAD

[Rou08b] Damian W. I. Rouson. Towards analysis-driven scientific software architecture: The case for abstract data type calculus. Scientific Programming, 16(4):329–339, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rauber:2004:ILO

[RR04] Thomas Rauber and Gudula Rünger. Improving locality for ODE solvers by program transformations. Scientific Programming, 12(3): 133–154, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rauber:2007:MTD

[RR07] Thomas Rauber and Gudula Rünger. Mixed task and data parallel executions in general linear methods. Scientific Programming, 15(3): 137–155, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Radhakrishnan:2015:UCP

[RRM^{+}15] Hari Radhakrishnan, Damian W. I. Rouson, Karla Morris, Sameer Shende, and Stavros C. Kassinos. Using coarrays to parallelize legacy Fortran applications: Strategy and case study. Scientific Programming, 2015(??): 904983:1–904983:12, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/904983/

.

Rico:2009:ATL

[RRV09] Alejandro Rico, Alex Ramirez, and Mateo Valero. Available task-level parallelism on the Cell BE. Scientific Programming, 17(1–2): 59–76, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rosing:1994:FLC

[RS94] Matt Rosing and Robert Schnabel. Flexible language constructs for large parallel programs. Scientific Programming, 3(3):169– 186, Fall 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Rosing:1995:LLM

[RS95] Matt Rosing and Joel Saltz.

Low latency messages on distributed memory multiprocessors. Scientific Programming, 4(1):35–43, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Reussner:2002:SCB

[RST02] Ralf Reussner, Peter Sanders, and Jesper Larsson Träff. SKaMPI: a comprehensive benchmark for public benchmarking of MPI. Scientific Programming, 10(1): 55–65, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C6%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Rouson:2004:DMQ

[RX04] Damian W. I. Rouson and Yi Xiong. Design metrics in quantum turbulence simulations: How physics influences software architecture. Scientific Programming, 12 (3):185–196, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Smith:2001:DMM

[SB01] Lorna Smith and Mark Bull. Development of mixed mode MPI/OpenMP applications. Scientific Programming, 9(2–3):83–98, Spring– Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.

asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C3%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Srirama:2011:SPS

[SBJV11] Satish Narayana Srirama, Oleg Batrashev, Pelle Jakovits, and Eero Vainikko. Scalability of parallel scientific applications on the cloud. Scientific Programming, 19 (2–3):91–105, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shan:2010:PMP

[SBM^{+}10] Hongzhang Shan, Filip Blagojević, Seung-Jai Min, Paul Hargrove, Haoqiang Jin, Karl Fuerlinger, Alice Koniges, and Nicholas J. Wright. A programming model performance study using the NAS parallel benchmarks. Scientific Programming, 18(3–4): 153–167, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shen:2016:POS

[SC16] Zhe-Ping Shen and Walter W. Chen. Profile orientation and slope stability analysis. Scientific Programming, 2016 (??):7029786:1–7029786:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/7029786/.

Soto:2016:EPS

[SCAP16] Ricardo Soto, Broderick Crawford, Boris Almonacid, and Fernando Paredes. Efficient parallel sorting for migrating birds optimization when solving machine-part cell formation problems. Scientific Programming, 2016 (??):9402503:1–9402503:39, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/9402503/.

Smallen:2002:AST

[SCB02] Shava Smallen, Henri Casanova, and Francine Berman. Applying scheduling and tuning to on-line parallel tomography. Scientific Programming, 10(4):271–289, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schweitzer:2015:PEM

[SCD^{+}15] P. Schweitzer, S. Cipière, A. Dufaure, H. Payno, Y. Perrot, D. R. C. Hill, and L. Maigne. Performance evaluation of multithreaded Geant4 simulations using an Intel Xeon Phi cluster. Scientific Programming, 2015(??): 980752:1–980752:10, ???? 2015. CODEN SCIPEV.

ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/980752/

.

Schmidt:1994:UNS

[Sch94] Henning Schmidt. Using naming strategies to make massively parallel systems work. Scientific Programming, 3(4):289–300, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schonfelder:2003:VPA

[Sch03a] J. L. Schonfelder. Variable precision arithmetic: a Fortran 95 module. Scientific Programming, 11(1):67– 76, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

http://pcwww.liv.ac.uk/
~jls/vpa20.f95;   http:
//pcwww.liv.ac.uk/~jls/
vpa20.htm.

Schupp:2003:LBC

[Sch03b] Sibylle Schupp. Lifting a butterfly — a componentbased FFT. Scientific Programming, 11(4):291–307, ???? 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sahni:2009:SSA

[SCSJ09] Onkar Sahni, Christopher D.

Carothers, Mark S. Shephard, and Kenneth E.

Jansen. Strong scaling analysis of a parallel, unstructured, implicit solver and the influence of the operating system interference. Scientific Programming, 17 (3):261–274, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schattler:2000:RPP

[SDS00] Ulrich Schättler, Günther Doms, and Jürgen Steppeler. Requirements and problems in parallel model development at DWD. Scientific Programming, 8(1): 13–22, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C2%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Skinner:1995:PPC

[SE95] Gregg Skinner and Rudolf Eigenmann. Parallel performance of a combustion chemistry simulation. Scientific Programming, 4(3): 127–139, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sheshadri:2003:GSPb

[SF03a] K. Sheshadri and Peter Fritzson. A general symbolic PDE solver generator: Beyond explicit schemes. Scientific Programming, 11(3): 225–235, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sheshadri:2003:GSPa

[SF03b] K. Sheshadri and Peter Fritzson. A general symbolic PDE solver generator: Explicit schemes. Scientific Programming, 11(1):39–55, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Smith:1996:DDS

[SG96] Barry F. Smith and William D.

Gropp. Design of datastructure-neutral libraries for the iterative solution of sparse linear systems. Scientific Programming, 5(4):329– 336, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schulz:2008:OSO

[SGM^{+}08] Martin Schulz, Jim Galarowicz, Don Maghrak, William Hachfeld, David Montoya, and Scott Cranford. Open — SpeedShop: an open source infrastructure for parallel performance analysis. Scientific Programming, 16 (2–3):105–121, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schonauer:1994:EGB

[SH94] W. Schonauer and H. Hafner. Explaining the gap between theoretical peak performance and real performance for supercomputer architectures. Scientific Programming, 3 (2):157–168, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sato:2001:CEO

[SHHI01] Mitsuhisa Sato, Hiroshi Harada, Atsushi Hasegawa, and Yutaka Ishikawa. Clusterenabled OpenMP: An OpenMP compiler for the SCASH software distributed shared memory system. Scientific Programming, 9(2–3):123– 130, Spring–Summer 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C6%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Skillicorn:1997:QAA

[SHM97] David B. Skillicorn, Jonathan M. D. Hill, and W. F. Mc-Coll. Questions and answers about BSP. Scientific Programming, 6(3): 249–274, Fall 1997. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL ftp://

ftp.comlab.ox.ac.uk/pub/
Documents/techpapers/Jonathan.
Hill/SkillHillMcColl_QA.
ps.gz.

Shu:1994:ADP

[Shu94]  Wei Shu.  Adaptive dy-

namic process scheduling on distributed memory parallel computers. Scientific Programming, 3(4):341– 352, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sinharoy:1999:COI

[Sin99] Balaram Sinharoy. Compiler optimization to improve data locality for processor multithreading. Scientific Programming, 7(1):21– 37, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C2%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Shahrivari:2015:FPA

[SJ15] Saeed Shahrivari and Saeed Jalili. Fast parallel allsubgraph enumeration using multicore machines. Scientific Programming, 2015(??):

901321:1–901321:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/901321/

.

Schonherr:2014:TSS

[SJR14] Jan H. Schönherr, Ben Juurlink, and Jan Richling. TACO: A scheduling scheme for parallel applications on multicore architectures. Scientific Programming, 22(3): 223–237, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Satoh:2001:COT

[SKS01] Shigehisa Satoh, Kazuhiro Kusano, and Mitsuhisa Sato. Compiler optimization techniques for OpenMP programs. Scientific Programming, 9(2–3):131–142, Spring–Summer 2001. CO-DEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=7pab6qgbaf8vxg991rwy%
26referrer=parent%26backto=
issue%2C7%2C11%3Bjournal%
2C1%2C9%3Blinkingpublicationresults%
2C1%2C1.

Shi:2009:ISC

[SKU^{+}09] Guochun Shi, Volodymyr V. Kindratenko, Ivan S. Ufimtsev, Todd J. Martinez, James C. Phillips, and Steven A. Gottlieb. Implementation of scientific computing applications on the Cell Broadband Engine. Scientific Programming, 17(1– 2):135–151, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Shi:2016:HBP

[SLBZ16] Chenghua Shi, Tonglei Li, Yu Bai, and Fei Zhao. A heuristics-based parthenogenetic algorithm for the VRP with potential demands and time windows. Scientific Programming, 2016 (??):8461857:1–8461857:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/8461857/.

Sun:2016:MSR

[SLD^{+}16] Xiaobing Sun, Bin Li, Yucong Duan, Wei Shi, and Xiangyue Liu. Mining software repositories for automatic interface recommendation. Scientific Programming, 2016 (??):5475964:1–5475964:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/5475964/.

Song:2000:MST

[SLJ^{+}00] H. J. Song, X. Liu, D. Jakobsen, et al. The MicroGrid: a scientific tool for modeling Computational Grids. Scientific Programming, 8(3): 127–141, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Slota:2012:SQP

[Slo12]  Renata Slota.  Storage

QoS provisioning for execution programming of dataintensive applications. Scientific Programming, 20(1): 69–80, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Schopf:2002:GTT

[SN02] Jennifer M. Schopf and Bill Nitzberg. Grids: The top ten questions. Scientific Programming, 10(2): 103–111, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                        26referrer=parent%26backto=
                                        issue%2C2%2C8%3Bjournal%
                                        2C1%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Snelling:1995:AAG

[Sne95] David F. Snelling. Applications analysis: Guest Editor’s introduction. Scientific Programming, 4(3):123– ??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Snyder:2007:SPF

[Sny07]  W. Van Snyder.  Scien-

tific programming in Fortran. Scientific Programming, 15(1):3–8, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sharp:1993:CSO

[SO93] Michael D. Sharp and Steve W. Otto. Class-specific optimizing compiler. Scientific Programming, 2(4):235– 238, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sen:2011:SII

[SO11] Alper Sen and Can Ozturan.

Special issue on the 9th International Symposium on Parallel and Distributed Computing. Scientific Programming, 19(1):1, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Soner:2015:GME

[SO15] Seren Soner and Can Ozturan. Generating multibillion element unstructured meshes on distributed memory parallel machines. Scientific Programming, 2015(??): 437480:1–437480:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/437480/

.

Spotz:2012:PRA

[Spo12] William F. Spotz. PyTrilinos: Recent advances in the Python interface to Trilinos. Scientific Programming, 20(3):311–325, ???? 2012. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Santana-Perez:2015:TRS

[SPPH15] Idafen Santana-Perez and María S. Pérez-Hernández. Towards reproducibility in scientific workflows: An infrastructure-based approach. Scientific Programming, 2015 (??):243180:1–243180:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/243180/

.

Szustak:2015:AMH

[SRO^{+}15] Lukasz Szustak, Krzysztof Rojek, Tomasz Olas, Lukasz Kuczynski, Kamil Halbiniak, and Pawel Gepner. Adaptation of MPDATA heterogeneous stencil computation to Intel Xeon Phi coprocessor. Scientific Programming, 2015(??):642705:1– 642705:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/642705/

.

Soto:2016:SPC

[SRTCL16] Ricardo Soto, Eduardo Rodriguez-Tello, Stéphane Caro, and Frédéric Lardeux. Scientific programming in computational intelligence. Scientific Programming, 2016(??): 5039793:1–5039793:2, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/5039793/

.

Schaffer:2000:DPA

[SS00] Daniel S. Schaffer and Max J.

Suárez. Design and performance analysis of a massively parallel atmospheric general circulation model. Scientific Programming, 8 (1):49–57, 2000. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=h82chcapth0xynh5tw5w%
26referrer=parent%26backto=
issue%2C6%2C6%3Bjournal%
2C5%2C9%3Blinkingpublicationresults%
2C1%2C1.

Silva:1997:IPD

[SSC97] Luis M. Silva, Joao Gabriel Silva, and Simon Chapple. Implementation and performance of DSMPI. Scientific Programming, 6(2):201– 214, Summer 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Stockinger:2002:GDM

[SSM^{+}02] Heinz Stockinger, Asad Samar, Shahzad Muzaffar, et al. Grid Data Mirroring Package (GDMP). Scientific Programming, 10(2): 121–133, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=1f99bpyvlg7t46lx8ue3%
                                        26referrer=parent%26backto=
                                        issue%2C4%2C8%3Bjournal%
                                        2C1%2C12%3Blinkingpublicationresults%
                                        2C1%2C1.

Shahriary:2015:CBA

[SSN15] Saeed Reza Shahriary, Mohsen Shahriari, and Rafidah MD Noor. A community-based approach for link prediction in signed social networks. Scientific Programming, 2015(??):602690:1– 602690:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2015/602690/

.

Schone:2014:TMM

[STD^{+}14] Robert Schöne, Jan Treibig, Manuel F. Dolz, Carla Guillen, Carmen Navarrete, Michael Knobloch, and Barry Rountree. Tools and methods for measuring and tuning the energy efficiency of HPC systems. Scientific Programming, 22(4): 273–283, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Stevenson:1997:RGI

[Ste97] D. E. Stevenson. Review:

Great ideas in computer science: a gentle introduction, by Alan W. Biermann. Scientific Programming, 6 (4):399–400, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Spafford:2015:ADS

[SV15] Kyle L. Spafford and Jeffrey S. Vetter. Automated design space exploration with Aspen. Scientific Programming, 2015(??):157305:1– 157305:10, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/157305/

.

Singh:2007:OWD

[SVR^{+}07] Gurmeet Singh, Karan Vahi, Arun Ramakrishnan, Gaurang Mehta, Ewa Deelman, Henan Zhao, Rizos Sakellariou, Kent Blackburn, Duncan Brown, Stephen Fairhurst, David Meyers, G. Bruce Berriman, John Good, and Daniel S. Katz. Optimizing workflow data footprint. Scientific Programming, 15 (4):249–268, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Sitek:2016:HPF

[SW16] Paweł Sitek and Jarosław Wikarek. A hybrid programming framework for modeling and solving constraint satisfaction and optimization problems. Scientific Programming, 2016 (??):5102616:1–5102616:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/5102616/.

Sakellariou:2004:LCR

[SZ04] Rizos Sakellariou and Henan Zhao. A low-cost rescheduling policy for efficient mapping of workflows on Grid systems. Scientific Programming, 12(4):253–262, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Savage:2009:EMA

[SZ09] John E. Savage and Mohammad Zubair. Evaluating multicore algorithms on the unified memory model.

Scientific Programming, 17 (4):295–308, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Samudrala:2015:PFD

[SZAG15] Sai Kiranmayee Samudrala, Jarosław Zola, Srinivas Aluru, and Baskar Ganapathysubramanian. Parallel framework for dimensionality reduction of large-scale datasets. Scientific Programming, 2015(??): 180214:1–180214:12, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/180214/

.

Song:2016:SOU

[SZC16] Jiekun Song, Kaixin Zhang, and Zijian Cao. 3Es system optimization under uncertainty using hybrid intelligent algorithm: A fuzzy chance-constrained programming model. Scientific Programming, 2016 (??):2675759:1–2675759:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/2675759/.

Song:2016:MMB

[SZXL16] Aibo Song, Maoxian Zhao, Yingying Xue, and Junzhou Luo. MHDFS: A memorybased Hadoop framework for large data storage. Scientific Programming, 2016 (??):1808396:1–1808396:12, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/1808396/.

Szymanski:2007:FPL

[Szy07] Bolesław K. Szymanski. Fortran programming language and scientific programming: 50 years of mutual growth. Scientific Programming, 15 (1):1–2, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Thakur:1996:ETP

[TC96] Rajeev Thakur and Alok Choudhary. An Extended Two-Phase Method for Accessing Sections of Outof-Core Arrays. Scientific Programming, 5(4):301– 317, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

                                        www.mcs.anl.gov/~thakur/
                                        papers/ext2ph.ps.

Truong:2004:SGU

[TF04] Hong-Linh Truong and Thomas Fahringer. SCALEA-G: a unified monitoring and performance analysis system for the Grid. Scientific Programming, 12(4): 225–237, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Turcu:2011:RTD

[TFN11] Gabriela Turcu, Ian Foster, and Svetlozar Nestorov. Reshaping text data for efficient processing on Amazon EC2. Scientific Programming, 19 (2–3):133–145, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Takizawa:2015:ODT

[THS^{+}15] Hiroyuki Takizawa, Shoichi Hirasawa, Makoto Sugawara, Isaac Gelado, Hiroaki Kobayashi, and Wen mei W. Hwu. Optimized data transfers based on the OpenCL event management mechanism. Scientific Programming, 2015(??): 576498:1–576498:16, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/576498/

.

Tao:2002:MAB

[TKS02] Jie Tao, Wolfgang Karl, and Martin Schulz. Memory access behavior analysis of NUMA-based shared memory programs. Scientific Programming, 10(1):45–53, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=9ejnuvwuvby9737jte27%
26referrer=parent%26backto=
issue%2C5%2C9%3Bjournal%
2C2%2C12%3Blinkingpublicationresults%
2C1%2C1.

Tran:2015:CLC

[TLC15] Nhat-Phuong Tran, Myungho Lee, and Dong Hoon Choi. Cache locality-centric parallel string matching on manycore accelerator chips. Scientific Programming, 2015(??): 937694:1–937694:20, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/937694/

.

Tanaka:2014:ISB

[TOF^{+}14] Teruo Tanaka, Ryo Otsuka, Akihiro Fujii, Takahiro Katagiri, and Toshiyuki Imamura. Implementation of dspline-based incremental performance parameter estimation method with ppOpen-AT. Scientific Programming, 22(4):299–307, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tang:2013:FLC

[TPKP13] Ping Tak Peter Tang, Jongsoo Park, Daehyun Kim, and Vladimir Petrov. A framework for low-communication 1-D FFT. Scientific Programming, 21(3–4):181–195,

????  2013.  CODEN

SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Thakar:2011:LSD

[TSCT11] Ani Thakar, Alex Szalay, Ken Church, and Andreas Terzis. Large science databases — are cloud services ready for them? Scientific Programming, 19(2–3):147–159, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Tian:2015:ESV

[TSP^{+}15] Xinmin Tian, Hideki Saito, Serguei V. Preis, Eric N. Garcia, Sergey S. Kozhukhov, Matt Masten, Aleksei G. Cherkasov, and Nikolay Panchenko. Effective SIMD vectorization for Intel Xeon Phi coprocessors. Scientific Programming, 2015(??): 269764:1–269764:14, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/269764/

.

Tang:2016:EER

[TT16] Xiaoyong Tang and Weizhen Tan. Energy-efficient reliabilityaware scheduling algorithm on heterogeneous systems. Scientific Programming, 2016 (??):9823213:1–9823213:13,

????  2016.  CODEN

SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/9823213/.

Tyma:1999:TVC

[Tým99] Paul Týma. Transient variable caching in Java’s stackbased intermediate representation. Scientific Programming, 7(2):157–166, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

                                        /iospress.metapress.com/
                                        app/home/contribution.
                                        asp%3Fwasp=f277qlrwwjr5m4vxjywv%
                                        26referrer=parent%26backto=
                                        issue%2C7%2C8%3Bjournal%
                                        2C7%2C9%3Blinkingpublicationresults%
                                        2C1%2C1.

Vermeulen:1993:OSI

[VC93] Al Vermeulen and Margaret Chapman. OON-SKI: an introduction. Scientific Programming, 2(4):109– 110, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Varela:2005:WCA

[VCT05] Carlos A. Varela, Paolo Ciancarini, and Kenjiro Taura. Worldwide computing: Adaptive middleware and programming technology for dynamic Grid environments. Scientific Programming, 13(4):255–263,

????  2005.  CODEN

SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

vanDam-Kleese:2001:ETI

[vDKH01] Kerstin van Dam-Kleese and Michael Hopewell. Enabling technologies for improved data management: Hardware. Scientific Programming, 9(1):11–25, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f2779jvvqg63jq64qwtm%
26referrer=parent%26backto=
issue%2C2%2C6%3Bjournal%
2C2%2C9%3Blinkingpublicationresults%
2C1%2C1.

Viklund:1995:OOO

[VF95] Lars Viklund and Peter Fritzson. ObjectMath — an object-oriented language and environment for symbolic and numerical processing in scientific computing. Scientific Programming, 4(4):229– 250, Winter 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Vishwas:2009:IPM

[VGC09] B. C. Vishwas, Abhishek Gadia, and Mainak Chaudhuri. Implementing a parallel matrix factorization library on the Cell Broadband Engine. Scientific Programming, 17(1–2):3–29, ???? 2009.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Verner:1993:DGF

[VHBR93] Diane A. Verner, Gregory L. Heileman, Kent G. Budge, and Allen C. Robinson. Development of generic field classes for finite element and finite difference problems. Scientific Programming, 2 (4):227–234, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Velu:2016:ETB

[VMRM16] Vengadeshwaran Velu, Norman Mariun, Mohd Amran Mohd Radzi, and Nashiren Farzilah Mailah. Equalization technique for balancing the modulation ratio characteristics of the single-phase-to-three-phase matrix converter. Scientific Programming, 2016 (??):6187926:1–6187926:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/6187926/.

Vollebregt:1997:ALP

[Vol97]  Edwin Vollebregt.  Ab-

stract level parallelization of finite difference methods. Scientific Programming, 6 (4):331–344, Winter 1997. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http:/

/iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=a5tkxhqqy9eefb7hwkak%
26referrer=parent%26backto=
issue%2C1%2C7%3Bjournal%
2C9%2C9%3Blinkingpublicationresults%
2C1%2C1.

Vetter:2013:SIS

[VR13] Jeffrey S. Vetter and Padma Raghavan. Special issue: Selected papers from Super Computing 2012. Scientific Programming, 21(3–4): 63–64, ???? 2013. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Visser:2002:FHM

[VRM02] Ingmar Visser, Maartje E. J.

Raijmakers, and Peter C. M. Molenaar. Fitting hidden Markov models to psychological data. Scientific Programming, 10(3):185–199, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Verstak:2003:BBS

[VRW^{+}03] Alex Verstak, Naren Ramakrishnan, Layne T. Watson, et al. BSML: a binding schema markup language for data interchange in problem solving environments. Scientific Programming, 11(3):199–224, 2003. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Varbanescu:2009:BHR

[VvAC^{+}09] Ana Lucia Varbanescu, Alexander S. van Amesfoort, Tim Cornwell, Ger van Diepen, Rob van Nieuwpoort, Bruce G. Elmegreen, and Henk Sips. Building high-resolution sky images using the Cell/B.E. Scientific Programming, 17 (1–2):113–134, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2002:OPG

[Wan02] Ping Wang. OpenMP programming for a global inverse model. Scientific Programming, 10(3):253–261, 2002. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Warren:1996:PDP

[War96] Karen H. Warren. PDDP, a data parallel programming model. Scientific Programming, 5(4):319–327, Winter 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Warren:2014:IPH

[War14] Michael S. Warren. 2HOT: an improved parallel hashed oct-tree N-body algorithm for cosmological simulation. Scientific Programming, 22 (2):109–124, ???? 2014. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wasserman:1995:BTN

[Was95] Harvey J. Wasserman. Benchmark tests on the new IBM RISC System/6000 590 workstation. Scientific Programming, 4(1):23– 24, Spring 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Williams:1995:SCF

[WB95] Dan Williams and Luc Bauwens. Simulation of compressible flow on a massively parallel architecture. Scientific Programming, 4(3):193– 201, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2016:GOS

[WB16] Chun-Feng Wang and Yan-Qin Bai. Global optimization for solving linear multiplicative programming based on a new linearization method. Scientific Programming, 2016(??):3204368:1– 3204368:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/3204368/

.

Wolski:1996:CEC

[WC96] Rich Wolski and David Cann. Compiler-enforced cache coherence using a functional language. Scientific Programming, 5(2):161–171, Summer 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wolters:1995:DPN

[WCG95] Lex Wolters, Gerard Cats, and Nils Gustafsson. Dataparallel numerical weather forecasting. Scientific Programming, 4(3):141–??, Fall 1995. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2007:EPP

[WCKD07] Joseph Wang, Yong Cao, Raed Kafafy, and Viktor Decyk. Electric propulsion plume simulations using parallel computer. Scientific Programming, 15(2): 83–94, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wang:2016:IEH

[WCZ^{+}16] Wenfeng Wang, Xi Chen, Hongwei Zheng, Zhihan Lv, Zhengjia Liu, Jing Qian, and Ping Hu. Intelligence in ecology: How Internet of Things expands insights into the missing CO_{2} sink. Scientific Programming, 2016(??): 4589723:1–4589723:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/

journals/sp/2016/4589723/

.

Walker:2007:DCW

[WD07] David Walker and Ewa Deelman. Dynamic computational workflows: Discovery, optimisation and scheduling. Scientific Programming, 15 (4):191–192, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wester:2008:SR

[Wes08] Michael Wester. Software review. Scientific Programming, 16(1):93–96, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Weissman:1993:POO

[WGF93] Jon B. Weissman, Andrew S.

Grimshaw, and R. D. Ferraro. Parallel object-oriented computation applied to a finite element problem. Scientific Programming, 2(4):133– 144, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wylie:2008:PMA

[WGW08] Brian J. N. Wylie, Markus Geimer, and Felix Wolf. Performance measurement and analysis of large-scale parallel applications on leadership computing systems. Scientific Programming, 16 (2–3):167–181, ???? 2008.

CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Willis:1993:ADA

[WHG93] A. G. Willis, M. P. Healey, and B. E. Glendenning. AIPS++ n-dimensional array classes. Scientific Programming, 2(4):239–246, Winter 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Walker:2007:DSS

[WHRH07] David W. Walker, Lican Huang, Omer F. Rana, and Yan Huang. Dynamic service selection in workflows using performance data. Scientific Programming, 15(4): 235–247, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wheat:1994:POS

[WMR^{+}94] Stephen R. Wheat, Arthur B. Maccabe, Rolf Riesen, David W. van Dresser, and T. Mack Stallcup. PUMA: an operating system for massively parallel systems. Scientific Programming, 3(4):275– 288, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Welsh:1999:UNS

[WOC99] Matt Welsh, David Oppenheimer, and David Culler. U-Net/SLE: a Java-based usercustomizable virtual network interface. Scientific Programming, 7(2):147–156, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=f277qlrwwjr5m4vxjywv%
26referrer=parent%26backto=
issue%2C6%2C8%3Bjournal%
2C7%2C9%3Blinkingpublicationresults%
2C1%2C1.

Wang:2015:HMS

[WS15] Jie-Sheng Wang and Na-Na Shen. Hybrid multiple softsensor models of grinding granularity based on cuckoo searching algorithm and hysteresis switching strategy. Scientific Programming, 2015(??):146410:1– 146410:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/146410/

.

Wilcox:2011:TSS

[WSB11] Chris Wilcox, Michelle Mills Strout, and James M. Bieman. Tool support for software lookup table optimization. Scientific Programming, 19(4):213–229, ???? 2011. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wu:1994:OSS

[WSP94] Min-You Wu and Wolfgang Schroder-Preikschat. Operating system support for massively parallel computer architectures: an introduction. Scientific Programming, 3(4):273–??, Winter 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wirawan:2009:HPP

[WSZK09] Adrianto Wirawan, Bertil Schmidt, Huiliang Zhang, and Chee Keong Kwoh. High performance protein sequence database scanning on the Cell Broadband Engine. Scientific Programming, 17 (1–2):97–111, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Wu:2016:DSR

[Wu16]  Xiuguo Wu.  Data sets

replicas placements strategy from cost-effective view in the cloud. Scientific Programming, 2016 (??):1496714:1–1496714:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/1496714/.

Wang:2016:HNT

[WWLG16] Guan Wang, Yuxin Wang, Hui Liu, and He Guo.

HSIP: A novel task scheduling algorithm for heterogeneous computing. Scientific Programming, 2016 (??):3676149:1–3676149:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/3676149/.

Xu:2016:IBH

[XML^{+}16] Honglong Xu, Rui Mao, Hao Liao, He Zhang, Minhua Lu, and Guoliang Chen. Index based hidden outlier detection in metric space. Scientific Programming, 2016 (??):8048246:1–8048246:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/8048246/.

Xue:2004:PPD

[XNQF04] Caijun Xue, Hong Nie, Qingying Qiu, and Peien Feng. A peer-to-peer distributed collaborative optimization system. Scientific Programming, 12(2): 121–131, ???? 2004. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Xiao:2016:UCV

[XST^{+}16] Lei Xiao, Yu Sheng, Guanlan Tan, Jianxin Wang, and Yi Pan. A user-customized virtual network platform for NaaS cloud. Scientific Programming, 2016(??): 9315672:1–9315672:6, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/9315672/

.

Xu:2015:MGS

[XTCC15] Rengan Xu, Xiaonan Tian, Sunita Chandrasekaran, and Barbara Chapman. Multi-GPU support on single node using directive-based programming model. Scientific Programming, 2015(??): 621730:1–621730:15, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/621730/

.

Yang:2016:NPM

[YDC^{+}16] Quan Yang, Zhihui Du, Zhoujian Cao, Jian Tao, and David A. Bader. A new parallel method for binary black hole simulations. Scientific Programming, 2016 (??):2360492:1–2360492:14, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/2360492/.

Yang:2016:URA

[YG16] Guibin Yang and Hongyu Gao. Uncertain risk assessment of knowledge management: Based on set pair analysis. Scientific Programming, 2016(??):2025892:1– 2025892:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/2025892/

.

Yamazaki:2015:CLR

[YTD15] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Computing low-rank approximation of a dense matrix on multicore CPUs with a GPU and its application to solving a hierarchically semiseparable linear system of equations. Scientific Programming, 2015(??):246019:1– 246019:17, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/246019/

.

Yang:2016:RSB

[YZ16] Kai Yang and Zhuhong Zhang. Racing sampling based microimmune optimization approach solving constrained expected value programming. Scientific Programming, 2016(??): 2148362:1–2148362:9, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2016/2148362/

.

Yang:2015:IOE

[YZZ15] Xu Yang, Mingbin Zeng, and Yanjun Zhang. Implementing and optimizing of entire system toolkit of VLIW DSP processors for embedded sensor-based systems. Scientific Programming, 2015 (??):507896:1–507896:7, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

                                        https://www.hindawi.com/
                                        journals/sp/2015/507896/

.

Zhu:2010:SFT

[ZA10] Qian Zhu and Gagan Agrawal. Supporting fault-tolerance for time-critical events in distributed environments. Scientific Programming, 18(1): 51–76, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2016:IAS

[ZCXQ16] Zhiying Zhang, Xiaozhen Chen, Wenwen Xiao, and Guijie Qi. Identifying and analyzing strong components of an industrial network based on cycle degree. Scientific Programming, 2016 (??):7340859:1–7340859:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.

hindawi.com/journals/sp/
2016/7340859/.

Zygmunt:2015:CUA

[ZG15] Anna Zygmunt and Bogdan Gliwa. The comparison of users activity on the example of Polish and American blogosphere. Scientific Programming, 2015(??): 907547:1–907547:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/907547/

.

Zhou:2016:API

[ZGS16] Yancong Zhou, Xudong Guo, and Xiaochen Sun. Acquisition pricing and inventory decisions on dualsource spare-part system with final production and remanufacturing. Scientific Programming, 2016 (??):8038045:1–8038045:10, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/8038045/.

Zhang:2008:DGS

[ZGW08] Ji Zhang, Qigang Gao, and Hai Wang. Discover gene specific local co-regulations from time-course gene expression data. Scientific Programming, 16(1):31–47, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zheng:2010:OUP

[Zhe10]  Yili Zheng.  Optimizing

UPC programs for multicore systems. Scientific Programming, 18(3–4):183– 191, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhou:2016:VMP

[ZHL16] Zhou Zhou, Zhigang Hu, and Keqin Li. Virtual machine placement algorithm for both energy-awareness and SLA violation reduction in cloud data centers. Scientific Programming, 2016 (??):5612039:1–5612039:11, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.
                                        hindawi.com/journals/sp/
                                        2016/5612039/.

Zima:2007:FLA

[Zim07]  Hans P. Zima.  From

FORTRAN 77 to localityaware high productivity languages for peta-scale computing. Scientific Programming, 15(1):45–65, ???? 2007. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:2016:ROR

[ZJZ^{+}16] Yong Zhang, Yunjian Jiang, Miner Zhong, Nana Geng, and Dandan Chen. Robust optimization on regional WCO-for-biodiesel supply chain under supply and demand uncertainties. Scientific Programming, 2016 (??):1087845:1–1087845:15, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/1087845/.

Zhang:2015:WCL

[ZLL^{+}15] Wenzhe Zhang, Kai Lu, Mikel Luján, Xiaoping Wang, and Xu Zhou. Writecombined logging: An optimized logging for consistency in NVRAM. Scientific Programming, 2015(??): 398369:1–398369:13, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2015/398369/

.

Zaki:2008:BDM

[ZRP08] Mohammed J. Zaki, Naren Ramakrishnan, and Srinivasan Parthasarathy. Biological data mining. Scientific Programming, 16(1): 3, ???? 2008. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhang:1999:CEC

[ZS99] Xiaodong Zhang and Lin Sun. Comparative evaluation and case studies of sharedmemory and data-parallel execution patterns. Scientific Programming, 7(1):1– 19, ???? 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/
app/home/contribution.
asp%3Fwasp=64cr5a4mg33tuhcbdr02%
26referrer=parent%26backto=
issue%2C1%2C7%3Bjournal%
2C8%2C9%3Blinkingpublicationresults%
2C1%2C1.

Zhou:2010:ABD

[ZSS^{+}10] Min Zhou, Onkar Sahni, Mark S. Shephard, Christopher D. Carothers, and Kenneth E. Jansen. Adjacencybased data reordering algorithm for acceleration of finite element computations. Scientific Programming, 18 (2):107–123, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Zhao:2016:BRS

[ZWCY16] Yang Zhao, Lei Wang, Xue Chen, and Futao Yang. Bulk restoration for SDN-based transport network. Scientific Programming, 2016(??): 1938312:1–1938312:8, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/1938312/

.

Zhang:2016:ODR

[ZWW16] Jianhua Zhang, Shuliang Wang, and Yixing Wang. Optimal design on robustness of scale-free networks based on degree distribution. Scientific Programming, 2016(??): 5659687:1–5659687:7, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/
journals/sp/2016/5659687/

.

Zheng:2016:RAS

[ZWZ16] Wei Zheng, Chen Wang, and Dongzhan Zhang. A randomization approach for stochastic workflow scheduling in clouds. Scientific Programming, 2016 (??):9136107:1–9136107:13, ???? 2016. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.
hindawi.com/journals/sp/
2016/9136107/.

Zeng:2015:RIF

[ZYLZ15] Yi Zeng, Shiqun Yin, Jiangyue Liu, and Miao Zhang. Research of improved FPgrowth algorithm in association rules mining. Scientific Programming, 2015 (??):910281:1–910281:6, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL

https://www.hindawi.com/

                                        journals/sp/2015/910281/

.